A Complete Bibliography of Publications in the Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences

halieutiques et aquatiques (2020–2029)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

21 August 2024 Version 1.23

Title word cross-reference

13 [326]. ^{14} [169, 576]. ^{15} [326]. ^{86} [267, 636]. ^{87} [267, 636]. _{2} [648]. _{Y Y} [687]. Δ [576, 169].

1 [514]. 127-year [525]. 19 [168]. 1980s [14].

3 [116]. 3-D [116]. 3-dimensional [527]. 3D [368].

aBACI [229]. ability [461, 86, 184, 403]. Abiotic [120, 581]. abundance [413, 641, 704, 98, 12, 369, 484, 131, 400, 410, 237, 376, 62, 429, 522, 457, 551, 207, 3, 611, 652, 373, 496, 412, 537, 312, 381]. abundant [437]. access [546, 530]. accessibility [132]. acclimation [277]. account [62]. Accounting [509, 151, 340, 675]. accounts [537]. accrual [34]. accumulation [73]. accuracy [561]. Accurate [110]. acid [548, 102, 494, 316]. acidic [567]. acidification [702]. acidity [494]. Acipenser [704, 354, 164, 58, 222, 386]. acoustic [679, 200, 185, 85, 89, 15, 475, 252, 338, 498, 293, 95, 1, 193]. acoustically [678]. acoustics [393]. acquisition [506]. across [469, 152, 365, 167, 81, 114, 620, 233, 451, 667, 35, 276, 654, 316, 366, 551, 357, 632, 627, 668, 517, 497]. Act [273]. Active [623, 252]. activities [558, 190, 222]. activity [148, 360]. Acute [178]. adaptation [395]. adaptations [719, 639]. Adaptive [105, 624, 352]. added [659]. addition [78]. additions [133, 61, 312]. Adirondack [287]. adjusting [682]. adjustment [16]. Adult [515, 146, 594, 201, 621, 226, 703, 553, 29, 171, 539]. adults [394]. advance [716, 600]. advancing [472, 286]. advantage [376]. advantages [172]. advice [511]. aestivalis [64]. affect [597, 170, 173, 235]. affected [189]. affecting [264, 642, 596]. affects [86, 208, 16, 311]. after [406, 391, 168, 229, 126]. Age [55, 559, 144, 605, 548, 468, 501, 151, 6, 114, 580, 211, 516, 651, 592, 354, 505, 222, 464, 169, 576, 38, 380, 403, 42, 481, 333, 694]. age-0 [211, 222]. Age-at-size [144]. age-based [505]. age-dependent [592]. age-specific [468]. ageing [315, 511]. agent [594, 488, 150]. agent-based [488, 150]. agents [42]. aggregation [560, 66]. aggregations [536, 667]. aging [144, 598]. Agricultural [627, 632]. al [533, 532]. alarm [512]. alarming [363]. Alaska [134, 56, 307, 320, 175]. albacares [55, 204]. Alberta [328]. alewife [64, 429, 543, 373]. alewives [51]. Algonquin [657]. algorithm [588]. Alien [65]. alligator [530]. allis [366]. allometric [412]. allometries [26]. allowing [22]. along [7, 139, 34, 117, 639]. Alosa [246, 690, 64, 543, 699]. Alpine [482]. alpinus [422, 451, 210]. alternative [470, 598, 715]. alterniflora [202]. alters [440, 215, 108, 524, 320]. ambient [58, 222]. America [666, 686, 699]. American [390, 690, 556, 7, 243, 141, 493, 458, 144, 326, 232, 270, 699, 140, 177]. americanus [141, 458, 144, 232, 100]. amino [548]. Ammocrypta [283]. Ammodytes [466]. among [445, 44, 212, 198, 80, 91, 718, 105, 285, 503]. amphidromous [249, 686]. amplicon [175]. Ana [447]. anadromous [581, 422, 212, 418, 686, 623, 395, 656, 646, 176, 210, 64, 543, 154, 76, 14, 612]. anadromy [95]. analyses [255, 432, 689, 92, 563, 557, 516, 542, 336, 320, 115]. analysis [405, 485, 702, 348, 131, 74, 703, 447, 340, 316, 5, 149, 323, 78, 386, 59, 688, 97, 591]. Analytical [601, 130]. analyzing [719, 397]. ancestral [699]. anchovy [719, 711]. ancient [670]. angled [16, 663]. angler [36, 336, 45]. anglers [304, 306]. angling [31, 610, 662, 295, 126]. Anguilla [15, 316]. animals [671]. Anisakis [606]. annual [121, 89]. anogenus [277, 280]. Anoplopoma [573]. Antarctic [508, 714]. antenna [98, 369]. antennas [260]. anthropogenic [46, 364, 257]. Aphanopus [11]. Apostichopus [531]. app [336]. Appalachian [289]. Apparent [670]. apple [707, 706]. apple-to-orange [707, 706]. Application [83, 204, 576, 614, 541, 711, 470, 36, 621, 516, 3, 57, 398]. Applications [444, 585]. applied [125, 536, 426]. Applying [665]. appreciation [156, 319, 645]. approach [371, 719, 393, 351, 152, 9, 685, 143, 83, 578, 629, 93, 716, 598, 600, 587, 574, 79, 150, 611, 567, 130, 383, 537, 372]. Approaches [286, 255, 221]. aquaculture [570, 458]. aquatic [73, 77, 273, 190, 671, 278, 426]. aragonitic [25]. archival [104, 171, 57]. archived [14]. Arctic [134, 419, 365, 179, 422, 390, 360, 628, 451, 430, 340, 163, 210, 235, 76, 498, 380, 364, 345]. Area [231, 569, 714, 371, 205]. areas [560, 338, 463, 509, 675, 431]. arrival [40]. arsenic [565, 547]. artedi [374]. Artificial [332, 640, 213, 355, 551, 409]. ash [524]. assays [2]. assemblage [108, 376]. assemblages [437, 195, 640, 294]. assembly [523]. assess [484, 483]. assessed [104, 448]. Assessing [437, 558, 202, 710, 399, 290, 59, 403, 45, 659, 125, 660, 655, 184]. assessment [461, 607, 683, 9, 143, 387, 456, 74, 618, 708, 136, 514, 352, 88, 132, 431, 150, 100, 511, 531, 573, 383, 126]. assessments [107, 141, 505, 161, 269]. assignment [579, 359, 342]. associated [486, 179, 6, 395, 164, 197, 436]. Association [251]. associations [274]. assumptions [161]. asymmetrical [229]. at-risk [392, 669, 643]. at-sea [118]. at-vessel [448]. Athabasca [442, 633]. Atlantic [271, 55, 225, 122, 405, 423, 138, 435, 377, 90, 397, 579, 476, 406, 614, 689, 473, 459, 550, 486, 563, 506, 8, 387, 213, 678, 212, 620, 211, 407, 239, 198, 180, 518, 80, 395, 296, 67, 186, 400, 494, 10, 104, 205, 253, 89, 379, 488, 112, 15, 376, 60, 649, 154, 181, 512, 502, 416, 500, 552, 356, 662, 544, 40, 342, 492, 699, 697, 238, 528, 21, 166, 515, 715, 446, 521, 467, 228, 192, 126, 616, 54, 385]. atlanticus [370]. Atractosteus [530]. attempts [12, 206]. augments [559]. automated [368]. Automatic [588]. autonomous [536, 260]. Autumn [234]. autumnalis [34]. availability [134, 456, 35, 718, 149]. avoid [707, 706]. avoidance [118]. Ayres [704].

Ba [664]. Ba/Ca [664]. back [472]. back-calculation [472]. background [399]. bacterial [524, 429]. bait [371, 638, 583]. baits [583]. Balaenoptera [513]. balance [420]. Balancing [456, 651, 283]. Balistes [3]. Baltic [713]. band [144]. Bank [227, 140, 177]. Banks [683]. bar [248]. bar-built [248]. barotrauma [663]. barriers [581]. Barrow [134]. bartramii [63]. base [542, 494]. based [271, 461, 125, 263, 519, 9, 687, 242, 536, 578, 518, 483, 201, 382, 236, 592, 106, 136, 488, 475, 505, 457, 588, 555, 38, 88, 150, 321, 221, 372, 717]. baseline [57]. Basin [327, 597, 273, 401, 201, 516, 330, 357, 279, 334, 61, 27, 115]. basins [92]. basis [550, 540]. bass [32, 314, 214, 27, 715, 172, 45, 306, 636]. Bay [719, 406, 711, 94, 199, 191, 498]. Bayesian [351, 711, 143, 290, 498, 40, 69, 52]. be [90, 548]. beach [466]. beam [251]. Bear [102]. beaver [310]. beds [443]. before [4, 229, 662]. beget [436]. behavior [371, 628, 638, 171, 150]. Behavioral [206, 540, 638]. behaviors [164]. behaviour [476, 256, 8, 146, 341]. Behavioural [31, 200, 16, 170, 512]. behaviours [219, 492]. being [265]. below [515]. benchmarks [601]. Benefits [535, 33, 43, 426]. benthic [538, 231, 640, 539]. Bering [464, 481]. Bet [525]. Bet-hedging [525]. better [260]. between [691, 393, 579, 549, 473, 387, 408, 483, 621, 120, 458, 571, 250, 699, 380, 42, 562, 192]. Bias [12, 187, 472, 519, 685, 341, 150, 129, 25]. biased [224]. Big [249]. bigeye [55]. bigmouth [297, 525, 652]. bioaccumulation [350, 565, 547]. biochemical [583]. biochronologies [129]. biocomplexity [676, 46]. biodiversity [363, 443, 430, 661]. Bioeconomic [340]. bioenergetic [103]. biofilms [539]. Biologgers [663]. biological [423, 563, 418, 72, 625, 514, 161, 320]. biology [44]. biomass [94, 199, 83, 582, 131, 493, 324]. biomass/recruitment [493]. biomonitor [364]. Biophysical [8, 380]. biopsy [157]. biota [73, 442, 633]. biotic [209, 289]. birds [445]. Biscay [719, 711]. bitumen [68]. Bivalvia [282]. black [484, 334]. blooms [101]. blue [191, 448, 203]. blueback [64]. bluefin [271, 400, 205, 342, 385]. bluefish [583]. boards [682]. bodied [239, 615]. body [469, 396, 700, 689, 244, 224, 185, 344, 76, 394, 684]. bomb [576]. boreal [565, 547, 650]. borealis [332, 361, 112]. borer [524]. Borrowing [36]. both [664, 170, 443]. Bottom [215, 337]. Bottom-up [337]. bound [23]. boundaries [258]. breeding [597, 86, 284]. British [216, 577, 672, 149, 491, 110, 125, 291, 165, 118, 578, 288, 631, 553, 590, 220, 677, 630, 514, 517]. Brittany [376]. broadband [393, 536]. brood [125]. brook [687, 264, 7, 147, 85, 695]. brown [31, 148, 292, 452, 198, 180, 546, 147, 596, 500, 310, 305, 599]. browning [440, 254]. Bryde [513]. brydei [513]. bubble [354, 48, 302]. Buccinum [521]. buffalo [297, 525, 652]. buffer [87]. build [615, 523]. built [248]. bull [258, 352]. bully [139]. burden [157]. Bycatch [361, 90, 449, 4, 528, 13, 345, 409].

C [508, 326, 169, 576]. Ca [664]. calcification [648]. calcium [222]. calculation [472]. calibrated [119]. California [418, 223, 676, 447, 658, 259, 175]. californicus [531]. Caligus [635]. Callinectes [191]. Cambridge [498]. cameras [355]. campechanus [371, 457]. Can [133, 548, 564, 566, 718, 150, 193]. Canada [294, 567, 392, 272, 138, 110, 125, 291, 585, 480, 165, 453, 286, 218, 657, 363, 318, 273, 578, 701, 518, 494, 466, 379, 281, 220, 643, 75, 280, 95, 431, 632, 627]. Canadian [188, 265, 123, 142, 419, 584, 28, 458, 602, 190, 60, 301, 279, 364, 345]. capability [343]. capacity [624, 418, 106, 308]. capriscus [3]. captive [512, 284, 394, 282]. captive-propagated [282]. captive-reared [512]. captivity [86, 80]. capture [98, 146, 173, 204, 4, 498]. capture-related [204]. captured [104]. carbo [11]. Carcass [61]. Carcharhinus [448]. carcharias [138, 476]. Carcharodon [138, 476]. Carnation [149]. carp [116, 652, 334]. cascade [412]. Case [106, 453, 418, 613, 141, 266, 220, 376, 60, 269, 383, 63, 345, 140, 177, 414]. Castor [310]. catastrophic [442, 633]. Catch [718, 271, 560, 66, 94, 199, 582, 382, 621, 516, 355, 62, 356, 662, 88, 511, 126, 65]. catch-and-release [662]. catch-at-age [516]. catch-based [88]. catch-curve [621]. catchability [613, 16, 638]. catches [205]. catching [682]. catchment [358]. catfish [482]. Catostomus [206, 275]. caught [333]. cautionary [689, 438]. Celtic [8, 716]. censored [62]. censoring [537]. Central [491, 676, 513, 57]. century [580, 619]. century-long [580]. cerebralis [388]. cerebrovascular [178]. cf. [275]. chalcogrammus [481]. challenges [585, 437, 256, 602, 168]. change [188, 265, 584, 309, 209, 190, 656, 646, 543, 119, 372]. Changes [594, 262, 244, 83, 145, 233, 534, 238, 380, 294, 511, 467, 381]. changing [549, 300, 296, 671, 526, 417, 126]. channel [123, 149, 331]. channels [360, 182]. char [422, 540, 623, 451, 340, 210, 235, 76, 498, 417, 207]. Characteristics [635, 263, 250, 358, 153, 668, 562]. Characterization [583]. characterize [93]. Characterizing [78]. charr [655]. chemical [299, 217]. chemistry [593, 218, 657, 308]. Chesapeake [191]. chi [335]. China [343]. Chinook [174, 291, 549, 103, 216, 248, 170, 477, 201, 669, 553, 415, 99, 347, 429, 39, 46, 654, 49, 527, 529, 109, 20, 57, 409, 503, 159]. choice [626]. Choosing [285]. chronologies [217]. chronometer [576]. chub [260]. chum [303, 491, 365, 390, 349]. cisco [374]. citizen [304]. Cladocera [294]. clam [364]. clarity [300]. clarkii [209, 276]. class [621, 516]. classes [114, 354, 464, 42]. classification [28, 99, 588, 130]. classifiers [59]. classy [130]. Clear [335]. clemensi [635]. Climate [624, 81, 695, 413, 469, 634, 265, 568, 218, 575, 378, 296, 209, 656, 646, 586, 671, 543, 129, 45, 126, 420, 712, 372, 385, 167]. climate-driven [712]. climate-related [378, 45]. climate-stressed [413]. climatic [595]. cline [253]. clock [694]. closely [39]. closures [550, 566, 718]. Clupea [423, 614, 8, 578, 713, 307, 324, 635]. clusters [650]. CMIP6 [575]. co [344, 648]. co-occurring [344]. coal [442, 633]. coast [473, 266, 699, 491, 626, 718]. coastal [575, 304, 443, 451, 330, 253, 677, 465, 526, 82, 595, 51, 718, 492, 504, 1, 539, 11]. cobia [470]. coccolithophores [648]. cod [413, 90, 473, 211, 36, 359, 10, 253, 56, 471, 21, 380, 166, 431, 54, 381]. coded [382]. codend [474, 361]. codends [142, 402]. coexisting [114]. Coho [110, 125, 520, 178, 245, 187, 553, 330, 35, 399, 275]. Cohort [344, 481, 54]. coincide [451]. Cold [586]. Cold-water [586]. collaboratively [449]. collapse [491, 719, 391, 471, 353]. collecting [304]. Colombia [327]. colonisation [298]. colonization [123]. Colorado [260]. Columbia [491, 110, 125, 291, 165, 578, 220, 216, 478, 577, 118, 642, 288, 631, 553, 61, 590, 677, 630, 672, 514, 149, 27, 517]. Combining [410, 560, 3]. come [359, 523]. Comment [631, 630, 672, 673, 532, 533]. commercial [626, 518, 355, 441]. commercially [298]. commersonii [206]. Commission [714]. common [541, 139, 161, 652, 599]. common-garden [599]. communication [132]. communities [538, 255, 437, 584, 218, 313, 647, 289, 189, 78, 128]. Community [523, 440, 188, 134, 660, 179, 83, 443, 475, 595, 71]. community-based [475]. Comparative [336]. compared [660, 612, 306]. Comparing [541, 88, 13, 22, 221]. Comparison [641, 236, 229, 23, 497, 83, 470, 282, 461]. comparisons [707, 621, 706]. compensate [48, 302]. compensation [26, 25]. competing [506]. Competition [570, 198, 167, 81, 655, 407, 527, 537]. Competitive [571, 86]. compilation [666]. complete [176]. complex [473, 459, 94, 199, 180, 667]. complexity [572]. complicated [207]. component [480]. composed [24]. composition [299, 440, 660, 179, 151, 470, 108, 651, 208, 194, 23, 528, 69]. compositions [14, 386]. compound [557]. compound-specific [557]. comprehensive [118, 598, 608]. compressed [691]. concentrations [632, 627]. concept [598, 155]. conceptual [19]. concern [263, 226, 311]. concordant [668]. condition [413, 256, 244, 211, 185, 61, 51, 197, 247, 639]. conditions [469, 134, 170, 478, 309, 80, 67, 490, 120, 205, 168, 380]. confirms [390]. Conflicts [445]. confluentus [258]. Congruence [431]. Connecting [522]. Connectivity [593, 693, 425, 44, 208, 112]. Consequences [300, 423, 12, 145, 460, 147, 270, 109, 161]. Conservation [225, 476, 110, 291, 585, 263, 570, 559, 212, 337, 239, 226, 163, 692, 661, 311, 75, 284, 680, 285, 321]. conservation-oriented [284]. conserving [586]. Considerations [195, 265, 375]. considering [426]. consistent [341]. conspecifics [174]. Constance [244]. constraints [418, 372]. construct [339]. constructed [123]. consumer [603]. consumers [215]. consumption [456, 421]. contacting [564]. contemporary [57]. content [421]. contents [710]. contingents [579]. continuous [519]. Continuum [155, 292]. contrast [696]. Contrasting [660, 82, 622, 144]. contrasts [377]. contribute [336]. contribution [110, 291, 182, 544, 599]. contributions [518, 267]. control [322, 227, 619, 389, 229, 441, 712, 608]. conventional [336, 1, 3]. convergence [114]. cooperative [613]. coordinated [113]. coral [449]. core [576]. Coregonus [244, 374, 482]. Correction [167, 199, 708, 404, 656, 565, 633, 302, 675, 632, 177, 472]. correlates [303, 76]. Coryphaena [171]. cotidianus [139]. Could [87, 90, 509, 675]. coulterii [398]. count [410]. counts [641, 561, 144, 40, 57]. coupled [605]. Coupling [1]. Courter [532]. covariates [370, 40]. covary [477]. cover [296]. COVID [168]. COVID-19 [168]. CPUE [62]. crab [160, 191, 203]. crabs [403]. crassispina [219]. Crassostrea [436]. crayfish [453]. creates [499]. Creek [529, 149]. creel [554]. crisis [634]. criterion [180]. critical [438, 471]. Cross [667]. Cross-scale [667]. crosses [212]. Crowsnest [328]. crustacean [188, 587]. cryptic [484]. cucumber [531]. cue [512]. cues [85]. Current [495, 255, 585, 27]. currents [492]. curve [382, 621]. cutthroat [261, 561, 572, 120, 209, 411, 276, 489, 539]. cyanobacteria [378]. cyanobacterium [34]. cycle [484, 669, 257]. cylindraceum [287]. cyprinellus [297, 525]. cyprinid [334].

D [673, 116]. dace [285]. daily [711, 62, 385]. Dakota [297]. dam [546, 270, 502, 499]. dam-induced [502]. damage [566]. Dams [649, 622, 180, 642, 515]. Daphnia [312]. dark [681]. darter [283]. Data [505, 278, 560, 666, 98, 107, 143, 94, 199, 339, 260, 370, 348, 532, 470, 613, 74, 304, 516, 651, 410, 340, 424, 376, 29, 62, 498, 637, 293, 88, 50, 652, 383, 119, 22, 345, 70, 414]. Data-limited [278, 370, 74, 505, 88, 414]. data-poor [340]. Data-rich [505]. data-weighting [22]. datasets [647]. David [240]. day [403]. days [542]. ddRAD [210]. dead [116]. deal [685]. dear [217]. decadal [307]. decade [73, 348]. decades [339]. decision [518]. decisions [17]. declarations [560]. decline [550, 494, 416]. declined [196]. declines [580]. declining [297]. Decreased [648]. decreasing [396]. deep [605, 360, 361]. deep-water [360, 361]. deepwater [576, 11, 694]. degree [542, 403]. degree-day [403]. degree-days [542]. degrees [294]. delayed [515]. delays [502]. delineate [677]. delta [538, 69, 214]. deluge [367]. demand [263]. demersal [682, 465, 463, 402]. demo [488]. demo-genetic [488]. Demographic [563, 614, 477, 241, 109, 320]. demographics [297]. demography [397]. demonstrate [106, 637]. demonstrates [426]. demonstration [66]. dendritic [108]. Dense [39]. densities [536, 317, 539]. Density [145, 469, 485, 121, 428, 229]. Density-dependent [145, 485, 428]. dentatus [79]. dependent [485, 339, 145, 592, 29, 428, 64, 609, 647]. deplete [515]. depleted [545]. depletion [234, 70]. deploying [355]. deposited [235]. depredation [582]. depth [606, 540, 139, 76, 500, 48, 302]. derived [406, 689, 169]. descending [663]. desert [624]. design [519, 84, 343]. design-based [519]. designs [229]. detect [184]. detectable [612]. detected [130]. Detecting [607, 424, 79, 258]. Detection [514, 98, 236, 281, 220, 338, 193]. detections [369, 236]. detects [536]. determinants [661]. determination [600]. determined [14]. Determining [678]. Detroit [338]. develop [389]. developed [449]. Developing [493, 716, 521]. development [657, 36, 662, 43, 694]. developmental [512]. devices [402, 681]. diadromous [33, 124]. diatom [183]. diatoms [9, 455]. did [478]. Didymosphenia [183]. Diel [352, 148, 76]. diet [606, 508, 267, 205, 181, 311]. Dietary [102]. differ [58, 154, 42]. Differences [506, 699, 197, 423, 146, 341, 34, 612]. different [461, 233, 552, 316, 149, 305, 119]. Differential [165, 563, 135]. differentially [597, 170]. differentiation [579, 352, 105]. Digital [604]. diluted [68]. dimensional [527]. diminishing [676]. dimorphic [700]. dimorphism [450]. Diploid [58]. Diplosoma [526]. Direct [551, 531, 110, 291, 158, 144, 254, 521, 175]. direct-developing [521]. directions [255]. directly [383, 434]. disappearance [219]. discard [554]. discards [421]. discernible [11]. discharge [597, 149]. Discovery [644]. discriminate [59]. Discrimination [393, 24, 99, 326]. Disease [256, 328, 429, 383]. Disentangling [257]. disparate [637]. dispersal [213, 314, 488, 29, 308, 492, 446, 639]. displacement [447]. displacements [154]. dissima [699]. dissolved [354, 358, 70]. distances [371]. Distinct [451, 718]. distributed [484]. distribution [597, 263, 690, 541, 215, 375, 209, 163, 658, 665, 307, 637, 280, 481, 13, 517, 128]. distributional [258]. distributions [271, 477, 26, 27]. disturbances [149]. diverge [313]. divergent [545]. Diverse [417, 437, 376, 269]. diversion [123, 288, 631, 630]. Diversity [462, 660, 92, 556, 374, 330, 46, 697, 688, 496, 282]. DNA [110, 291, 585, 207, 446, 659]. Do [250, 174, 359, 48, 302, 42]. Does [173, 428, 564]. doing [368]. dolomieu [314, 27, 172]. DOM [358]. downstream [622, 679, 534, 270, 499]. downstream-migrating [270]. Dreissena [101]. dreissenid [325]. drift [359, 609, 527, 259, 323]. drift-feeding [609, 527]. Driftless [569]. drive [134, 262, 606, 135, 718, 668]. driven [712, 381]. driver [374, 603]. Drivers [550, 162, 448, 392, 439, 365, 19, 654, 496, 404]. Drought [534, 367, 153]. Drought-related [534]. due [187, 193, 372]. Duration [86, 403]. duration-to-moult [403]. during [560, 594, 10, 56, 206, 168, 649, 502, 14, 317, 510, 228, 11]. Dutch [251]. dwelling [428]. Dworshak [312]. dynamic [493, 667, 527, 566, 119]. dynamics [371, 397, 391, 690, 167, 81, 241, 296, 325, 214, 714, 574, 358, 5, 509, 675, 50, 507].

Early [67, 355, 396, 485, 216, 452, 8, 678, 629, 236, 359, 235, 512, 362, 43]. eastern [283, 4, 342, 481]. eat [102]. eco [309]. eco-evolutionary [309]. Ecological [693, 375, 147, 357, 246, 255, 77, 460, 374, 618, 708, 108, 325, 139, 499, 87]. ecologically [276, 298]. Ecology [225, 256, 655, 337, 316, 659]. ecomorphs [540]. Economic [556, 555]. economic-based [555]. Ecosystem [653, 303, 440, 614, 183, 702, 94, 199, 239, 603, 716, 106, 191, 714, 269, 331, 717, 65, 495]. ecosystem-based [106, 717]. ecosystems [480, 313, 137, 49, 504, 668]. ecotype [442, 633]. ecotypes [102]. edge [163]. Editorial [487]. eDNA [585, 660, 443, 2]. Edward [95]. eel [541, 15, 270, 384, 316, 504, 418]. eelgrass [443]. Effect [326, 644, 425, 581, 185, 354, 306, 25]. effective [371]. effectiveness [229, 284]. Effects [73, 327, 700, 501, 183, 569, 638, 713, 543, 277, 181, 502, 662, 454, 51, 274, 117, 275, 261, 31, 178, 262, 445, 485, 568, 6, 570, 682, 532, 223, 198, 141, 137, 325, 135, 534, 340, 2, 254, 222, 658, 82, 184, 399, 171, 258, 512, 356, 492, 567, 539]. Efficacy [674, 4]. Efficiency [368, 84, 651, 615, 69, 681]. effort [66, 408, 651, 511, 434]. efforts [87]. egg [711, 116]. eggs [713]. eight [590, 343]. electrofishing [660, 264, 652]. Electronic [671, 697]. elemental [701, 217]. Elevated [235, 58, 277, 357, 510, 648]. elevation [597, 120]. embayment [492]. Emerald [524]. emergence [691]. emergency [634]. emerging [602, 328]. emigration [1]. emission [73]. empirical [666, 470, 191]. empirically [321]. enables [574]. enclosure [35]. encounter [317]. endangered [33, 406, 570, 339, 367, 442, 633, 274, 257, 321]. endemic [239]. endocrine [238]. endothermic [678]. Energetic [292, 172]. energetics [105]. energy [506, 515]. engineer [183]. England [51, 528]. enhances [625]. Enhancing [269]. enigmatic [459]. Enriched [625]. entrainment [53, 69]. environment [549, 570, 241, 625, 324, 650, 648, 119]. Environmental [585, 439, 179, 245, 223, 222, 661, 446, 188, 134, 584, 9, 370, 141, 667, 565, 547, 340, 526, 654, 40, 279, 207, 334]. environments [541, 395, 526, 394]. Epigenetic [694, 333]. Epinephelus [269]. episodically [525]. epizootic [635]. equal [691]. era [48, 302, 368]. eradication [264]. Erie [233, 47, 96]. ESA [201]. ESA-listed [201]. Escape [554, 564, 681]. escapement [201]. Esox [396, 6, 638]. establishment [190, 283, 526, 388, 334]. estimate [12, 36, 493, 410, 89, 1, 611, 57]. estimated [72, 511]. estimates [689, 468, 6, 369, 249, 679, 201, 187, 169, 551, 203, 69, 320]. Estimating [299, 704, 98, 94, 382, 703, 475, 349, 498, 555, 3, 698, 582, 220, 62, 293, 199]. Estimation [291, 5, 38, 271, 641, 110, 605, 711, 365, 151, 685, 333, 694, 479]. estimators [415]. estuarine [666, 351]. estuarine-opportunist [351]. estuary [329, 216, 248, 690]. eulachon [4, 175]. euphausiids [375]. Eurasian [310]. European [557, 15, 384, 316, 504, 497, 482]. eutrophication [372]. evaluate [351, 519]. Evaluating [33, 273, 231, 415, 692, 161, 293, 111, 589, 712, 409, 90, 15, 284]. Evaluation [188, 264, 315, 449, 304, 679, 441, 715, 97, 591, 157, 103, 687, 615, 505, 637, 717]. events [671]. eventually [510]. every [564]. Evidence [118, 64, 463, 166, 159, 550, 378, 578, 518, 400, 289, 530, 284, 386, 385]. evidence-based [578, 518]. Evolution [225, 309, 471]. evolutionary [460, 309]. evolving [420]. ex [582]. ex-vessel [582]. examining [437]. example [220, 111]. Exceptionally [226]. exercise [58]. exhaustive [58]. exhibit [216]. exilicauda [335]. exists [284]. exogenous [267]. exotic [188]. Expanded [691]. expansion [419, 526, 48, 302, 45]. experience [180]. experiment [16, 35, 384, 599, 653]. Experimental [66, 157, 234, 546, 534, 229, 713, 531, 275]. experiments [448]. expert [40]. explain [469, 632, 627, 403]. explained [153]. explicit [711, 127, 493]. Explicitly [717]. exploitation [249, 203]. exploited [601, 11]. exploration [148]. exposed [178]. exposure [512, 364, 381]. exposures [624]. expression [594, 418, 42]. expressions [601]. extended [384]. extension [219]. extent [476]. external [36]. externally [11]. Extreme [384, 671]. extremes [520]. Eye [169, 335, 576].

facilitate [133, 101, 649]. facilities [15]. facing [625]. Factors [642, 189, 614, 581, 264, 135, 596, 569, 565, 547, 326, 334]. failed [505]. fall [365, 546]. fall-timed [546]. family [592]. farm [506, 523]. farmed [5, 616]. farms [435, 44, 514, 644, 446]. fasciola [281, 282]. fasciolaris [282]. faster [356]. fatty [102, 316]. fauna [421]. favourable [656, 646]. features [657]. fecundity [423, 356]. federally [642]. feedback [428]. Feeding [192, 606, 623, 609, 527, 705, 247, 43, 513]. female [86, 226, 356]. females [703, 109]. fence [641]. fewer [535]. fiber [310]. fidelity [401]. field [68, 511, 193]. field-sampling [511]. Fifteen [273]. fimbria [573]. fin [6]. Fine [371, 280]. Fine-scale [371, 280]. fingerprints [701]. First [390, 249, 10, 228]. Fish [472, 606, 443, 299, 560, 392, 664, 134, 624, 255, 605, 480, 437, 660, 607, 444, 300, 614, 548, 256, 102, 581, 264, 545, 83, 318, 84, 369, 249, 158, 322, 350, 408, 484, 535, 558, 701, 710, 211, 195, 233, 367, 483, 603, 224, 236, 173, 542, 108, 19, 182, 135, 93, 615, 208, 565, 547, 16, 237, 600, 29, 254, 465, 640, 23, 184, 643, 609, 124, 601, 574, 298, 499, 495, 705, 289, 48, 302, 551, 530, 113, 709, 105, 604, 343, 611, 257, 59, 129, 96, 130, 650, 71, 668, 383, 698, 497, 496]. fish [192, 128, 639, 684, 193, 52, 335, 331, 653]. fisher [150]. Fisheries [168, 110, 125, 291, 585, 265, 300, 152, 550, 559, 146, 30, 223, 304, 518, 602, 415, 438, 336, 17, 716, 667, 355, 106, 340, 505, 421, 587, 124, 647, 555, 88, 509, 675, 528, 50, 589, 402, 97, 531, 13, 119, 717, 65]. Fishery [626, 327, 719, 425, 142, 575, 66, 556, 143, 160, 249, 582, 118, 449, 337, 231, 141, 173, 651, 361, 204, 297, 475, 60, 4, 554, 665, 463, 427, 566, 718, 434, 126, 608, 345]. fishery-independent [141, 651]. fishery-induced [160]. fishes [693, 33, 445, 286, 501, 183, 622, 629, 121, 447, 586, 592, 625, 615, 569, 326, 500, 169, 576, 595, 279, 284, 333]. Fishing [412, 305, 450, 371, 327, 432, 263, 501, 223, 190, 355, 16, 647, 454, 566, 111, 434, 414]. Fishing-induced [305, 566]. fishway [180, 206, 343]. fitness [245, 145, 625]. fitness-related [245, 625]. Fitting [346]. five [91]. Fixed [474, 410]. fixed-location [410]. fjord [360]. flatfish [1]. flavescens [338]. fleet [251, 718]. flexibility [266]. flexible [84]. Flight [37]. floodplain [322, 367, 530]. floods [546]. flounder [683, 424, 79, 100]. flow [397, 206, 502, 124, 352, 149, 259, 323, 172, 321, 275]. flows [534, 362]. Floy [104]. fluctuating [241]. fluctuation [640]. fluctuations [21]. flumes [198]. fluxes [366]. focus [717]. following [702, 146, 494, 610, 529]. fontinalis [687]. Food [406, 211, 68, 183, 137, 677]. for-hire [434]. forage [480]. Foraging [259, 67, 703, 254, 352]. forecasting [574]. forecasts [433]. forest [538]. Forestry [577, 673, 672]. forests [524]. form [246, 608]. formation [10]. formulations [470]. foundation [705]. four [563, 233, 288, 631, 630]. Fourier [605, 59]. Fragmentation [108, 32]. fragmented [331]. framework [614, 141, 493, 410, 184, 441]. frameworks [541, 587, 13]. France [384]. Fraser [435, 216, 220]. free [582, 665]. French [487]. fresh [170, 240]. Freshwater [664, 469, 392, 272, 327, 585, 700, 351, 286, 68, 541, 28, 581, 363, 318, 603, 451, 602, 565, 547, 430, 254, 543, 643, 504, 284, 589, 128, 684]. freshwaters [301]. Fringe [258, 205]. frontiers [709]. fry [422, 213, 224]. full [376]. fully [449]. fulvescens [164, 222, 386]. function [339, 301]. functional [608]. Fundy [406, 94, 199]. fungal [356]. future [255, 605, 155, 190, 168, 27].

Gadus [10, 380, 481, 54]. gar [530]. garden [599]. gas [354, 457, 48, 302]. gastric [144]. gates [246]. Gaussian [191]. gear [327, 432, 266]. gears [84]. geminata [183]. gene [594, 323, 42]. generalized [621]. generate [691]. generated [679]. generating [436]. generation [389, 599]. generic [601]. genes [253]. Genetic [163, 411, 282, 125, 579, 570, 143, 245, 32, 540, 374, 359, 253, 488, 194, 210, 399, 528, 386, 521, 513, 599]. genetics [298]. Genome [323]. Genome-wide [323]. Genomic [395, 139, 46, 197]. genotyping [210]. genus [82]. geochemistry [452]. geographic [430]. Georges [227]. ghost [130]. giant [531, 539]. gill [157]. gillnet [244, 665]. gillnets [497]. glanis [482]. glauca [448]. global [634, 313, 385]. go [359]. goal [283]. goals [436]. Gobiomorphus [139]. goby [571, 661, 357, 41]. gonadal [408]. gorbuscha [591]. government [619, 644]. gradient [7, 35, 639]. gradients [218, 575, 108, 139, 34, 193]. Grand [683, 140, 177]. graphic [49]. Graphical [191, 388]. grass [116]. gray [3]. grayling [198, 163]. Great [300, 348, 618, 708, 190, 617, 279, 334, 102, 273, 629, 147, 139, 357]. greater [580]. green [704]. Greenland [593, 360, 205]. gross [603]. ground [400]. groundfish [517]. groundfishes [4]. grounds [647, 513]. group [441]. grouper [269]. groups [527]. GrowChinook [49]. growing [403]. Growth [247, 228, 666, 405, 624, 396, 549, 152, 216, 486, 439, 151, 6, 570, 346, 202, 572, 603, 296, 400, 542, 703, 185, 10, 61, 534, 35, 82, 250, 184, 47, 399, 49, 79, 384, 705, 38, 51, 504, 197, 394, 385, 462, 275, 653]. Guidance [542]. guide [345]. guideline [252]. guilds [255]. Gulf [593, 614, 223, 231, 104, 56, 457, 169, 595, 269, 507, 333]. Gut [421].

Habitat [32, 686, 504, 444, 103, 690, 519, 287, 478, 577, 66, 127, 318, 422, 118, 374, 572, 186, 546, 53, 451, 466, 85, 569, 56, 465, 692, 522, 232, 609, 672, 124, 310, 566, 149, 695, 274, 27, 117, 562, 63, 684, 52, 335, 426]. habitat-use [451]. habitats [224, 586, 176, 316, 530]. haddock [227]. hake [470, 658]. halibut [593, 118, 104, 697, 119]. Hamilton [193]. hand [376]. haplotypes [616]. Harbour [193]. harengus [713, 324, 423, 614, 8]. harm [283]. Harvest [17, 538, 92, 233, 415, 441, 380, 715, 203, 45, 712, 450, 608]. hatcheries [416]. Hatchery [478, 148, 174, 125, 24, 486, 245, 532, 395, 67, 621, 676, 187, 553, 347, 552, 529, 20]. hatchery-origin [174, 187]. hatchery-stocked [621]. Hawai’i [582]. haystack [636]. headwater [572, 108, 524]. health [634, 558, 651]. heat [159]. heating [653]. heatwave [56]. heavily [11]. hedging [525]. Heliocidaris [219]. hematopoietic [220]. hepatic [408]. herring [246, 423, 614, 550, 439, 8, 578, 64, 713, 307, 324, 635, 528, 247]. heterogeneity [386]. hidden [453, 46]. Hierarchical [652, 521, 143, 210, 386, 69, 52]. High [365, 401, 234, 239, 120, 226, 236, 669, 502, 311, 368, 697, 1, 172]. high-elevation [120]. high-flow [502]. High-resolution [365]. high-survival [502]. high-throughput [236, 368]. higher [536, 317]. highlight [115]. highly [665, 65]. hindcasting [574]. Hippoglossoides [140, 177]. Hippoglossus [697]. hippurus [171]. hire [434]. historic [693, 660]. historical [12, 143, 431, 364, 57]. historically [427]. histories [233]. history [377, 549, 689, 452, 30, 145, 418, 395, 330, 176, 235, 46, 26, 362, 544, 699, 688, 639, 450, 462]. history-associated [395]. hitch [335]. holography [604]. Homarus [141, 458, 144, 232]. homing [401, 670]. homogenization [289]. Honshu [219]. hook [537]. hooknose [170]. Hoplostethus [370]. Horizon [75]. hormone [468, 399]. host [594, 23]. Hudson [690, 89]. human [190]. Humboldt [495]. humpback [260]. hunting [638]. Huron [382, 621, 516]. hybrid [148, 506]. hybridization [387, 30, 212, 120]. hydraulic [339]. hydroacoustic [536, 611]. hydroacoustics [437, 94, 199, 497]. hydrodynamic [116, 325]. hydroelectric [642, 53, 288, 631, 630]. hydrographic [205]. hydrologic [624]. Hydrological [322]. hydrology [121]. hydropower [122, 622, 180, 15, 48, 302, 343]. hyperstability [66]. hypolimnetic [70]. hypothesis [227, 562]. hypoxia [215, 465, 277, 274, 275, 96]. hypoxic [331].

ice [460, 296, 610]. ice-angling [610]. Iceland [111]. iconic [431]. Ictiobus [297, 525]. Idaho [103, 57, 312]. ideal [665]. identification [125, 221]. identify [365, 7, 628, 636]. Identifying [560, 629, 46, 496, 614]. If [615]. II [714]. illumination [4]. immigration [546]. Impact [131, 72, 396, 231, 237, 229, 511, 96]. Impacts [219, 157, 620, 558, 667, 191, 311, 364, 294]. impaired [452]. impairment [610]. Impeding [546]. imperiled [46, 335, 115]. imperilled [286, 284, 285]. implantation [185]. Implications [488, 149, 476, 183, 30, 212, 535, 407, 443, 268, 438, 667, 430, 2, 692, 543, 661, 352, 384, 166, 274, 100, 11]. Importance [220, 614, 296, 421, 429]. important [480, 298]. improve [142, 80, 18, 566, 509, 675, 50]. Improved [151, 529, 128, 669, 389]. improvement [494]. improves [683, 201, 332, 517, 573]. Improving [74, 433, 132, 681, 479, 386]. in-season [143]. in-stream [131]. including [90, 442, 633]. Inclusion [414]. incompatible [619]. Inconspicuous [138]. incorporate [141]. incorporated [265]. Incorporating [92, 369, 688, 320, 338, 383, 717]. incorporation [406, 326]. increase [546, 61]. increased [362]. increases [396, 572, 539]. Increasing [266, 361, 229, 342]. increasingly [671]. increments [346, 385]. incubation [713]. independent [141, 651]. index [674, 74, 495, 573]. indicator [587]. indicators [584, 83, 71]. indices [683, 230, 537]. Indigenous [263, 526]. Indirect [254]. Individual [342, 296, 341, 564, 639]. individuals [535]. indivisible [634]. induce [113]. Induced [350, 160, 145, 502, 566, 305, 639, 331]. induces [442, 633, 25]. industrial [202]. industry [613]. industry-cooperative [613]. infection [435, 256, 644]. Infectious [42, 594, 220]. infer [127]. inferences [519]. inferred [593, 359, 316]. Inferring [532, 26]. infestation [165]. Influence [209, 538, 397, 167, 81, 581, 211, 190, 553, 47, 46, 714, 654, 20, 203, 481]. influences [577, 245, 222, 672, 394]. influencing [189]. inform [337, 136, 688, 321]. information [619, 74, 36, 3, 320]. Informative [483]. informed [141]. informs [163]. infrared [605]. inhabiting [182, 650]. inhibitor [2]. initial [211]. INLA [680]. inland [152, 75]. inner [406]. innovative [130]. inputs [511]. inquiry [619]. insights [664, 702, 325, 105, 150, 694, 173]. Instream [572, 352, 172]. intakes [604]. integrated [370, 325, 688]. Integrating [89, 637, 50, 611, 517, 624, 143, 84, 613, 69]. integrative [124]. intelligence [355]. interacting [441]. interaction [425]. Interactions [549, 458, 198, 209, 268, 625, 571, 237, 192, 691]. interactive [261, 132]. Interannual [330, 77, 508]. interdependencies [124]. interface [49]. interfemale [160]. intergenerational [314]. interior [61]. intermediate [215, 30]. intermedius [11]. international [454, 714]. Interplay [465, 380]. Interpopulation [503]. Interpretation [77]. interrogation [130]. Interspecific [407, 198, 237]. Intertidal [466]. intra [160]. intra- [160]. intracoelomic [185]. introduced [268, 99]. Introducing [432]. Introduction [225, 680]. Introgression [616, 253]. Inuvialuit [419]. invaded [65]. invader [188, 27]. Invaders [113, 453]. invasion [524, 661]. Invasive [325, 693, 351, 158, 101, 121, 120, 236, 190, 571, 389, 308, 298, 357, 604, 43, 334, 482]. Inverse [536]. Invertebrate [123, 259]. Investigating [714]. investigation [341]. invisible [391]. involving [613]. ion [632, 627]. ionoregulatory [612]. Island [95]. Islands [644]. isotope [114, 702, 508, 557, 703, 316]. isotopes [7, 677, 636]. isotopic [23, 442, 633, 113]. issues [663, 75]. Iteroparity [544].

jack [170]. Jamestown [297]. Japan [219]. Japanese [574]. Joaquin [214]. John [379]. Joint [447, 128]. Joint-species [447]. jointly [235]. junctions [117]. Juvenile [510, 691, 31, 666, 157, 178, 397, 165, 103, 390, 211, 655, 679, 198, 67, 53, 553, 61, 354, 465, 229, 35, 338, 47, 49, 527, 574, 51, 105, 644, 197, 247, 539, 54, 275]. juvenile-to-adult [553].

Kamchatka [462]. keeled [583]. Keeping [130]. keta [303, 491, 365, 349]. key [555]. keystone [453]. kidney [429]. Killarney [567]. killer [317]. kinetics [58, 222]. king [160]. kisutch [520, 553, 35, 275]. knowledge [419, 40]. known [641]. Kokanee [30, 267, 354, 353, 312]. Kootenay [78, 353]. krill [714]. krill-predator [714]. krill-predators [714].

Lab [687]. Lab-based [687]. Labrador [340, 54]. lack [467]. lagging [265]. Lake [657, 244, 610, 598, 153, 668, 538, 485, 102, 702, 378, 460, 401, 655, 540, 239, 101, 382, 621, 516, 53, 494, 268, 410, 135, 670, 208, 164, 194, 176, 222, 250, 500, 695, 386, 43, 567, 87, 482, 102, 456, 655, 233, 382, 621, 516, 325, 534, 91, 194, 429, 47, 181, 416, 290, 78, 386, 96, 373, 353, 335, 398]. lake-rich [153]. lake-specific [250]. Lakes [273, 188, 538, 234, 179, 218, 460, 243, 162, 404, 590, 522, 358, 189, 632, 627, 294, 497, 455, 420, 300, 348, 629, 618, 708, 147, 190, 139, 617, 357, 279, 334]. Lampmussel [281]. lamprey [7, 679, 200, 341, 389, 206, 612, 510]. lampricides [389]. Lampsilis [281, 282]. lance [466]. land [108, 366, 632, 627]. landing [554]. landings [425]. Landlocked [239, 612, 212, 198, 395, 67, 186, 238]. landscape [263, 657, 451, 153, 274]. landscape-scale [274]. landscapes [152, 430]. Large [125, 182, 495, 597, 122, 216, 92, 127, 580, 313, 239, 483, 101, 226, 451, 162, 404, 268, 615, 670, 109, 595, 334, 482, 636, 639]. large-bodied [239, 615]. Large-scale [125, 226]. largemouth [214, 45, 306, 636]. larger [396, 535]. largest [640, 697]. larvae [7, 713, 639]. Larval [112, 44, 8, 160, 315, 400, 236, 254, 47, 705, 604, 496]. LASSO [575]. lastridge [402]. latent [574]. later [174, 619]. lateral [25]. latipes [145]. latitudinal [623]. Laurentian [300, 348, 618, 708, 190, 617, 334]. lavaretus [482]. Lavinia [335]. Lawrence [273, 701, 308, 280, 593, 614, 104]. lead [380]. Leading [265]. leads [30, 464]. lean [655, 540]. leaping [583]. learning [605, 355, 588, 426]. legacy [149]. Lemhi [103]. length [472, 473, 439, 606, 242, 233, 651, 38, 88, 589, 69, 479]. length-based [88]. length-stratified [38]. lens [169, 576]. lenses [335]. lentic [49, 426]. less [158]. lessons [155, 355, 488, 285]. leucomaenis [623]. leveed [248]. level [640, 153, 573]. levels [606, 149]. lewisi [209, 276]. lice [165, 5, 644]. Life [377, 68, 669, 461, 549, 485, 689, 452, 30, 145, 484, 418, 629, 395, 359, 330, 176, 235, 26, 654, 362, 544, 699, 257, 688, 639, 450, 462]. Life-cycle [669]. life-history [689, 145, 418, 26, 544]. life-history-induced [639]. life-stage [330]. Lifetime [468]. light [332, 522, 409]. like [103]. likelihood [470]. likely [435]. limit [578]. limited [370, 74, 603, 505, 88, 278, 510, 414]. limits [62, 589, 434]. limnological [294]. linear [83]. linked [490]. Linking [241, 603, 647, 294, 256]. lipid [234, 462]. listed [272, 201]. listerianum [526]. little [249, 306, 260]. littoral [179, 322]. live [436, 130]. lived [468, 185, 592, 297, 507, 694]. lobster [90, 556, 231, 141, 493, 458, 144, 232, 414]. lobsters [403]. local [262, 65]. location [626, 410, 20]. loci [395]. Long [244, 297, 380, 77, 216, 468, 580, 195, 80, 185, 592, 694]. Long-lived [297, 468, 185, 592, 694]. Long-term [244, 380, 77, 216, 195, 80, 185]. longfin [323]. longimanus [448]. Longitudinal [516, 562]. longline [582, 448, 517, 537]. Look [518]. Lookingglass [529]. loops [428]. Loss [56, 32, 695, 69]. losses [257]. Lost [545, 582]. lotic [609]. louse [635]. low [29, 181, 149]. low-flow [149]. low-thiamine [181]. Lower [434, 210]. lucius [396, 638]. lunch [582]. Lutjanus [371, 457].

M [687]. Machine [426, 605, 355, 588]. mackerel [579, 528]. macroinvertebrate [179, 189]. macrophyte [101]. macrophytes [609]. macroscale [184]. Madeira [11]. Magdalena [327]. magnetic [154]. magnitude [58]. mahi [171]. mahi-mahi [171]. main [516]. Maine [502, 556, 231, 270]. mainstem [117]. maintained [282]. maintaining [287]. major [400]. making [518, 253]. Male [164, 86, 170, 264, 508, 347, 356]. mammal [111]. manage [505]. Management [60, 714, 138, 585, 265, 444, 287, 92, 28, 264, 318, 339, 30, 449, 618, 708, 341, 438, 696, 716, 667, 106, 692, 617, 124, 384, 463, 5, 111, 509, 675, 709, 166, 274, 100, 436, 591, 11, 321, 221, 608, 717]. management-scale [341]. Managing [602, 459, 239]. mandated [402]. manipulations [275]. Mapping [232, 647]. marginal [324]. marina [202]. Marine [76, 495, 440, 405, 480, 437, 406, 44, 216, 501, 678, 477, 313, 211, 224, 451, 421, 232, 111, 509, 675, 434]. marine-derived [406]. marinus [679, 200, 341, 206, 612, 510]. Maritimes [458]. maritimum [435]. mark [93, 29, 457, 384, 203]. markets [718]. masquinongy [6]. mass [461]. masses [701]. mat [34]. mat-scale [34]. Matching [701]. materials [514]. maternal [267, 713]. matter [7, 524, 564, 358]. Maturation [309, 568, 347, 228, 394, 140, 177]. mature [560]. maturity [685, 233, 10, 380]. maximum [436]. may [101, 380]. measurement [158]. Measures [71, 449]. measuring [702]. medaka [145]. mediate [120, 250, 684]. mediated [526]. mediates [561]. medirostris [704]. Mediterranean [554, 463]. meets [85, 355]. melanostomus [357]. membras [713, 324]. Memoriam [37]. Memoriam/In [37]. menhaden [715]. mentella [142]. mercury [73, 350]. meromictic [378]. mesh [474, 361]. Mesocentrotus [219]. mesocosms [68]. meta [131]. meta-analysis [131]. metabarcoding [629, 443, 421, 659]. Metabolic [276, 700, 486]. metabolism [331]. metal [294]. metapopulation [33, 262, 120, 29]. method [536]. methodology [472]. methods [641, 505, 457, 293, 88, 368, 97, 22, 345]. métier [555]. metric [403]. metrics [256, 26]. Mexico [457, 169, 595, 269, 507, 333]. Michigan [91, 194, 524, 429, 386, 373]. microcephalus [360]. microchemistry [359, 14, 95]. Microcoleus [34]. Microcystis [101]. microecological [25]. microhabitat [352]. Microlitter [379]. microplastic [618, 708, 707, 706, 617]. Microplastics [590]. Micropterus [314, 27, 172]. microstructures [315]. mid [478]. mid-Columbia [478]. migrant [642, 669]. migrants [342]. migrating [157, 679, 15, 270, 500]. Migration [122, 104, 303, 299, 260, 367, 596, 670, 649, 552, 595, 197, 42, 599, 11]. migration-associated [197]. migrations [246, 581, 594, 484, 359]. migratory [435, 292, 127, 367, 451, 410, 164, 308, 499, 492, 697, 530, 462]. mill [144]. millimetres [252]. mine [442, 633]. miniature [679]. mining [558]. Minnesota [162, 404]. minnow [670]. Minor [80, 714]. mismatches [408]. mitigate [45]. mitigation [409]. mitochondrial [616]. Mixed [289, 386, 110, 125, 291, 143, 94, 199, 205, 415, 220, 97, 539]. mixed-species [199]. mixed-stock [110, 125, 291, 205, 415, 220, 97]. mixture [704, 98]. mixtures [386]. mobile [127, 665, 130, 115]. model [461, 626, 683, 83, 484, 231, 36, 116, 466, 592, 204, 136, 488, 505, 637, 40, 50, 652, 69, 257, 688, 13, 591, 573, 412, 321]. model-resistant [505]. modeled [691]. Modeling [520, 287, 592, 574, 70, 152, 325, 669, 29, 150, 63, 345, 372]. Modelling [116, 492, 541, 370, 375, 280, 680, 119]. models [704, 461, 9, 127, 8, 339, 346, 692, 290, 498, 601, 637, 433, 278, 128, 52, 479, 331, 191]. modern [368]. modes [259]. modification [310]. modulates [640, 228]. modulation [324]. monitored [449, 448]. Monitoring [420, 119, 84, 118, 195, 443, 628, 336, 355, 163, 617, 604, 71, 70, 398]. montane [258]. Moore [673]. morhua [10, 380, 54]. morio [269]. morphological [9, 25]. morphology [30, 67, 236]. morphology-based [236]. morphometrics [368]. morphometry [632, 627]. Mortality [126, 396, 234, 90, 568, 501, 160, 382, 621, 226, 187, 592, 204, 448, 362, 705, 293, 1, 257, 306, 54]. motility [212]. moult [403]. mountain [358]. Mountains [489]. movement [213, 483, 628, 410, 93, 665, 290, 203, 41]. Movements [53, 351, 260, 186, 259]. MSE [441]. mullet [583]. Multi [307, 587, 301, 327, 690, 555, 567]. Multi-decadal [307]. multi-gear [327]. Multi-indicator [587]. multi-métier [555]. multi-proxy [567]. Multi-scale [301]. multi-species [327, 555]. multi-system [301]. multidecadal [29]. Multidisciplinary [28]. Multimethod [600]. multimodel [49]. Multiple [7, 654, 93, 625, 111, 567, 13, 608, 426]. multiscale [569]. multispecies [425, 456, 449, 441, 434, 52]. multistate [351, 365, 93, 498]. municipal [364]. Murray [36]. muskellunge [6]. mussel [101, 325]. mussels [286, 379, 284]. my [217]. Mya [364]. Myctophidae [659]. mykiss [86, 676, 354, 95, 259, 197, 394, 462]. Myxobolus [388].

N [704, 508, 23, 326]. N-mixture [704]. namaycush [401, 655, 540, 516, 87]. narrowband [393]. Nass [580]. Natal [314, 593, 656, 646, 670]. National [87]. native [261, 392, 693, 445, 183, 114, 535, 622, 121, 120, 214, 447, 268, 569, 571, 99, 276, 661, 699, 113, 95, 87, 616]. natural [174, 478, 657, 80, 592, 35, 298, 238, 257, 306, 497, 305]. natural-origin [174]. naturalized [529]. naturally [567]. Nature [540, 634]. navigable [41]. navigation [693]. near [605, 536, 166]. near-infrared [605]. near-surface [536]. nearshore [134, 82]. necrosis [220]. need [707, 706, 637]. needle [636]. needs [286]. negative [428]. Neogobius [357]. Nephrops [554, 414]. nerka [146, 594, 267, 354, 220, 433, 42, 353]. net [536]. network [344, 551, 412]. networks [425, 83, 251]. neural [83]. Neuromorphological [512]. Newfoundland [473, 166, 54, 140, 177]. newly [123, 186]. Next [389]. Next-generation [389]. niche [114, 374, 91, 237, 442, 633, 113, 373]. niches [640]. Nitellopsis [101]. nitrate [34]. No [582, 185]. NOAA [168]. noisy [424]. non [261, 392, 114, 214, 447, 569, 526, 276, 699, 95, 604, 41, 306, 87, 616]. non-indigenous [526]. non-invasive [604]. non-native [261, 392, 114, 214, 447, 569, 276, 699, 95, 87, 616]. non-navigable [41]. non-tournament [306]. Nonanadromous [225, 154]. nondietary [102]. nonindigenous [453]. nonlinear [507]. nonparametric [433]. Nonstationary [477]. nonwadeable [84]. norm [309]. North [297, 460, 666, 577, 390, 686, 243, 4, 326, 672, 699, 521, 513]. north-temperate [460]. northeast [595, 167, 81, 375, 253, 696, 380, 192]. northeastern [323]. northern [396, 476, 219, 401, 703, 18, 332, 340, 112, 638, 457, 317, 558, 431, 175]. northward [138]. Northwest [210, 303, 579, 476, 459, 112, 63]. northwestern [55, 342]. norvegicus [554, 414]. Norway [504, 414]. Norwegian [500]. Note [156, 319, 487, 645, 689]. Notropis [277, 280]. Nova [423]. Novel [333, 468, 685, 702, 674, 395, 383]. nucleotide [110, 291, 342, 175]. nudus [219]. Numbers [461]. Nunavik [422]. Nunavut [210]. nursery [29]. nursery-dependent [29]. nurture [540]. nutrient [133, 366, 632, 627, 78, 312]. nutrients [406]. nutritional [440, 377, 7, 372].

O [648]. obesus [55]. objectives [608]. observations [168]. observed [271]. observer [613]. obstacles [246]. obtained [497]. obtusa [101]. Occupancy [281, 262, 365, 280, 52]. occurring [344]. ocean [313, 677, 3, 167, 81, 375, 4, 342, 659, 63]. oceanic [253, 448, 14]. Oceano [55]. oceanographic [575]. odontocetes [582]. off [423, 473, 283, 166, 54, 140, 177]. offer [172]. offs [92, 105]. offshore [251]. offspring [212, 662]. Oga [219]. ogive [685]. oil [68, 457]. old [142, 15, 464, 611]. Ommastrephes [63]. Oncorhynchus [303, 491, 520, 86, 365, 146, 594, 209, 676, 553, 267, 99, 354, 347, 220, 35, 276, 349, 433, 20, 95, 259, 197, 42, 591, 57, 394, 353, 409, 462, 275]. one [634, 640, 25]. one-side-only [25]. only [25]. Ontario [657, 280, 294, 567, 73, 456, 325, 281, 181, 416, 455, 193]. ontogenetic [620, 335]. ontogeny [598]. open [432, 44, 243, 3]. open-marine [44]. open-water [243]. openness [361]. operated [642]. operations [622]. opercula [468]. opportunist [351]. opportunities [585, 256, 602]. Opposing [427, 356]. opposite [464]. Optimal [566, 63, 608]. optimized [49]. optimizing [84]. options [287]. orange [370, 707, 706]. Oregon [330, 529, 539]. organic [7, 524, 23, 358]. orientation [154]. orientations [437]. oriented [284]. origin [271, 174, 187, 205, 552, 636]. Oryzias [145]. other [613, 192]. Otolith [10, 99, 95, 299, 593, 452, 346, 315, 217, 359, 267, 23, 308, 169, 368, 511, 59, 129, 230, 25]. otolith-bound [23]. otolith-derived [169]. otoliths [664, 605, 24, 701, 14, 636]. otter [682]. outbreak [429]. outcomes [696]. Outmigration [329]. outputs [511]. overestimated [230]. overexploited [427]. Overlap [562, 91, 237]. overlapping [362]. Oversummer [35]. overview [255, 585]. overwhelms [570]. overwinter [234, 211]. oxygen [70]. Oxythermal [374]. oyster [438]. Oysters [436].

Pacific [167, 81, 375, 4, 63, 413, 691, 303, 666, 597, 133, 165, 419, 387, 390, 470, 118, 578, 223, 466, 696, 56, 658, 82, 665, 307, 635, 161, 417, 699, 323, 247, 97, 513, 119, 712]. pacificus [175]. Paleolimnological [378, 430, 567]. pallasii [578, 307, 635]. Pandalus [332, 361, 112]. panel [564]. panels [39, 564]. paradigm [213]. Paradox [230]. Paralichthys [79]. parameters [468, 72, 26, 38]. parametric [204, 433]. parasitic [165]. Parentage [201, 125]. Parentage-based [201, 125]. parental [31, 399]. Park [657, 567, 87]. parr [422]. part [618, 708, 617]. Partial [260]. partially [24]. participation [626]. particle [400]. particle-tracking [400]. Partitioning [255, 267, 222, 373]. partner [25]. pass [615]. Passage [210, 581, 642, 615, 206, 270, 499]. past [155]. pathogen [44]. pathways [451]. pattern [127, 226]. Patterns [392, 303, 299, 491, 607, 44, 477, 451, 217, 104, 565, 547, 91, 643, 290, 352, 552, 298, 464, 699, 323, 41, 25]. peak [362]. pelagic [244, 215, 682, 710, 677, 448, 495, 718, 709, 96]. pelagics [717]. pelican [445]. pellucida [283]. penaeid [507]. Peninsula [219]. Penobscot [405, 270, 502]. Perca [338]. perceived [311]. perception [439]. perch [338]. performance [541, 486, 687, 682, 251, 407, 180, 67, 15, 276, 171, 88, 247, 59, 591, 573, 22]. peri [482]. peri-Alpine [482]. period [76]. Periodic [21]. permafrost [430]. Persistence [622]. persistent [602, 667, 380]. Personality [486]. personatus [466]. perspective [460, 168, 431]. perspectives [545, 155, 430, 301, 269]. Petromyzon [679, 200, 341, 206, 612, 510]. phase [14]. phenological [460]. phenology [303]. phenotype [713, 588]. phenotypes [170]. Phenotypic [139]. phosphorus [325]. photosynthesis [648]. phylogenetic [514]. Physical [418, 246, 98, 583, 117]. Physiological [372, 664, 146, 610, 58, 238]. physiology [219, 612, 105, 503]. Phytoplankton [584, 313, 78]. pike [396, 638]. pink [591]. piscivore [700, 710]. piscivorous [350, 224]. pit [353, 98, 260, 130]. placement [564]. plaice [140, 177]. planktivores [96]. planktivorous [192]. plankton [440]. planning [106, 232]. plans [550, 716]. plant [604]. plasticity [173]. platessoides [140, 177]. platform [304]. platforms [457]. plume [371]. point [578, 509, 675, 320, 134]. points [423, 493, 72, 555]. policies [715]. policy [92]. pollock [464, 481]. pollution [618, 708, 617]. polymorpha [101]. polymorphisms [110, 291, 342, 175]. ponds [287, 523]. Poor [499, 340]. poorly [480]. pop [104, 171]. pop-up [104, 171]. Population [298, 175, 513, 413, 291, 520, 397, 579, 329, 485, 98, 167, 81, 92, 245, 145, 241, 239, 209, 190, 359, 667, 297, 424, 194, 210, 429, 14, 310, 362, 574, 109, 324, 50, 3, 43, 100, 688, 320, 306]. population-specific [14]. populations [491, 263, 212, 620, 139, 253, 283, 376, 235, 39, 181, 601, 612, 111, 699, 357, 695, 386, 96, 45, 285, 282]. portfolio [568]. portion [190]. position [557]. possession [434]. possible [168]. Post [413, 719, 377, 678, 492, 394]. post-collapse [719]. Post-settlement [413]. post-smolt [394]. post-smolts [678, 492]. post-spawned [377]. posthatching [212]. postlarvae [249]. postlarval [69]. Postrelease [148]. postsmolts [192]. Potential [402, 129, 624, 445, 387, 655, 447, 669, 336, 237, 421, 109, 454, 45, 57]. Powell [534]. power [559, 195, 604]. powered [40]. powerhouse [642]. practical [345]. practitioners [680]. prairie [523]. pre [212, 202, 356, 394]. pre- [212, 394]. pre-industrial [202]. pre-spawning [356]. Precambrian [218]. precautionary [587]. precipitation [494]. precision [561, 519]. precocious [347]. Predation [620, 445, 478, 678, 204, 429, 649]. predator [456, 714, 311, 668, 353]. predators [678, 613, 268, 714]. predatory [182, 500]. predict [278, 652, 42]. predicted [372]. Predicting [263, 243, 646, 507, 52, 331, 198, 49, 656]. prediction [128]. predictions [13, 517]. predictive [559, 50]. predictor [207]. predictors [251]. preeminence [378]. preference [276, 500, 583]. preferences [422, 135, 684]. preliminary [393]. prerecruits [56]. prerelease [187]. presence [25]. pressure [83, 16, 334]. prevalence [401, 220]. prevalent [159]. previous [581]. prey [134, 215, 242, 456, 35, 311, 317, 715]. preyed [224]. primary [7, 202, 374, 603, 137, 196]. Prince [307, 95]. Prionace [448]. prior [202, 180, 26]. priorities [459, 318, 337]. priority [463]. priors [483]. probabilistic [705]. probabilities [498]. probability [377, 347, 554]. problem [611]. procedures [221]. process [106, 389]. processes [664, 614, 408, 603]. produce [376]. producer [202]. production [263, 711, 137, 135, 196, 431, 57, 698, 221]. productivity [478, 603, 72, 654, 529, 372]. profiles [468, 102, 594, 703]. program [570, 264]. programs [407, 284, 71]. progress [273]. project [575]. projected [695]. projects [288, 631, 630]. promote [436]. promotes [238]. propagated [282]. propagation [478]. propagule [334]. properties [583]. proportion [342]. Prosopium [287, 398]. protected [445, 509, 675]. protection [286]. provenance [600]. provide [530]. provides [702]. Provincial [657, 567]. provisioning [227]. proxies [356]. proxy [567]. PRV [514]. PRV-1 [514]. PS752 [37]. pseudoharengus [64, 543]. Pseudopleuronectes [100]. Ptychobranchus [282]. pufferfish [574]. pugnose [277, 280]. pulse [574]. pulsed [182]. pups [703]. pygmy [398].

qPCR [2]. quality [440, 103, 572, 534, 164, 465, 275]. Quantification [366]. Quantifying [491, 579, 171, 308, 398]. quarter [619]. quarter-century [619]. Quebec [218, 558]. quota [118].

R. [673]. racemization [548]. radiocarbon [576]. rainbow [234, 263, 242, 120, 354, 442, 633, 352, 95, 462]. Range [696, 526, 469, 476, 419, 219, 276, 366, 193]. Range-wide [696]. ranges [338, 650]. Ranking [191]. Rapid [494, 43, 118]. rapidly [515]. rate [486, 47]. rates [405, 700, 549, 570, 603, 382, 400, 703, 553, 592, 326, 705, 1]. ratio [648]. rations [247]. ratios [664, 508]. rays [6]. Re [90, 283]. re-establishment [283]. Re-evaluating [90]. Reach [34, 519]. Reach- [34]. reach-scale [519]. reaches [28]. reaction [309]. real [461, 143, 647, 604]. real-life [461]. real-time [143, 604]. really [102]. reared [67, 621, 512]. Rearing [170, 216, 656, 646, 625, 27, 394]. rebuilding [550]. recapture [483, 93, 29, 498, 457, 384, 203]. recede [259]. reconnaissance [487]. reconstructing [516]. reconstructions [161]. record [664]. records [660]. recovering [138, 194]. recovery [133, 391, 286, 702, 146, 535, 669, 454, 471, 427, 20, 274]. recreational [425, 152, 263, 66, 337, 190, 336, 17, 204, 203, 434, 126, 414]. recruit [532, 525, 97, 320]. Recruitment [214, 614, 248, 575, 227, 493, 72, 162, 404, 19, 428, 82, 47, 26, 324, 427, 21, 386, 591, 573, 712]. recruits [194]. Red [457, 371, 269, 531]. Redd [561, 57, 562]. redds [213]. Redefining [100]. redfish [459, 142, 60]. redside [285]. reduce [478, 449, 4, 566, 434, 193]. Reduced [248, 118, 648]. reduces [474, 440, 350, 407]. reduction [74, 361].

Reductions [651, 73]. reef [169, 576, 551, 436, 333]. reference [423, 578, 493, 72, 555, 509, 675, 320]. reflect [219]. reflecting [294]. reflex [610]. refuges [661]. refugia [586, 526]. regard [583]. regime [691]. regimes [243, 144, 380]. region [73, 134, 153, 197]. Regional [262, 491, 677, 323, 589, 467, 420]. regression [621, 62]. regularization [575]. regulated [530]. regulation [424]. regulations [450]. reintroduced [80, 237, 310]. reintroduction [406, 198, 181, 529]. reintroductions [285]. related [245, 378, 625, 565, 547, 204, 534, 39, 14, 197, 45]. relates [373]. relation [83, 562]. relationship [473, 657, 251, 493, 250, 324, 471]. relationships [405, 144, 149, 650, 52, 479]. relative [614, 478, 264, 382, 108, 136, 343]. release [17, 356, 662, 284, 87, 126]. released [204, 514, 394]. releases [553]. releasing [546]. relevance [44]. relevant [276]. reliant [415]. Remarkable [121]. remediation [202]. remote [365, 475]. removal [644]. repeated [205]. replace [402]. replacement [562]. Reply [631]. report [107]. reproduction [390]. reproductive [468, 687, 408, 674, 164, 543, 181, 109, 495, 394]. rescue [411]. Research [318, 286, 273, 613, 707, 706, 426]. reserves [234]. Reservoir [297, 312, 122, 53, 534, 636]. reservoirs [393, 640]. residence [543]. residency [666, 360, 628]. resident [30, 147, 317]. residential [657]. resilience [568, 495]. resistant [505]. resolution [365]. resolve [39]. resolved [385]. resource [339, 603, 117, 372]. resource-based [372]. resource-limited [603]. respond [684]. Response [533, 707, 429, 31, 520, 68, 673, 121, 200, 288, 631, 630, 310, 362, 153, 312, 409]. Responses [279, 691, 218, 58, 206, 154, 718, 357, 668, 420, 653]. responsiveness [188]. restoration [131, 407, 268, 229, 124, 172]. restored [186]. restricted [213]. results [356, 193]. retention [8, 116]. Retrospective [405, 607]. return [405, 377, 382]. returns [1]. reveal [468, 563, 8, 346, 336, 237, 663, 298, 697, 95, 335]. revealed [508, 384]. Revealing [452]. reveals [157, 581, 378, 580, 447, 669, 411, 29, 421, 210, 342, 323, 599, 653]. revenue [647]. review [686]. revisited [598]. Revisiting [213]. rich [505, 153]. Richards [592]. richness [392, 84, 208, 52]. right [368]. riparian [524]. Risk [273, 392, 453, 618, 708, 669, 643, 278]. risks [257]. River [327, 103, 216, 690, 577, 418, 447, 61, 379, 270, 502, 672, 280, 27, 115, 246, 174, 597, 397, 92, 28, 239, 483, 603, 186, 490, 116, 288, 631, 108, 630, 661, 20, 528, 530, 278, 41, 639, 331, 405, 435, 704, 478, 155, 273, 260, 701, 580, 201, 328, 89, 220, 338, 442, 633, 308, 654, 159]. riverine [322, 622, 224, 314, 410, 496]. rivers [84, 623, 642, 19, 34, 699, 650, 117]. Riverscape [19, 365, 85]. RNA [220]. road [455]. road-salt [455]. roadway [178]. Roanoke [32]. robust [637]. robustness [712]. rockfish [468, 118]. rockfishes [82]. Rockies [584]. Rocky [489]. Role [568, 8, 408, 116, 522, 609, 416, 358, 334]. ropes [402]. roughy [370]. Round [41, 287, 628, 571, 661, 357]. routine [230]. routing [367]. ruberrimus [468]. rule [608]. rules [441, 712]. run [249, 288, 631, 410, 89, 630, 161]. run-of-river [288, 631, 630]. runoff [178]. runs [249, 503]. Russia [462].

sablefish [696, 266, 573]. Sacramento [704, 214]. sagittal [24]. Saint [379]. salamanders [539]. salar [122, 406, 486, 678, 407, 180, 395, 15, 649, 512, 356, 492, 515, 192]. salinity [316, 510]. Salish [274, 275, 466, 196, 417]. Salmo [122, 292, 406, 486, 678, 407, 180, 395, 596, 569, 15, 649, 512, 310, 356, 492, 515, 192, 305, 599]. Salmon [225, 691, 303, 666, 491, 174, 122, 405, 157, 435, 110, 125, 291, 520, 133, 178, 377, 397, 262, 165, 406, 549, 689, 626, 44, 103, 216, 248, 419, 170, 486, 365, 167, 81, 92, 12, 506, 245, 387, 390, 213, 146, 30, 594, 678, 212, 620, 477, 580, 407, 239, 198, 180, 80, 395, 201, 296, 224, 67, 186, 490, 147, 226, 656, 646, 187, 669, 553, 267, 415, 330, 99, 488, 347, 220, 15, 376, 252, 35, 429, 399, 665, 39, 46, 649, 349, 154, 181, 512, 502, 514, 416, 654, 49, 500, 311]. salmon [527, 362, 552, 529, 109, 433, 356, 662, 161, 5, 544, 20, 40, 417, 492, 238, 149, 644, 515, 446, 97, 320, 42, 591, 57, 467, 228, 192, 126, 616, 409, 503, 275, 159]. salmonid [577, 339, 131, 288, 631, 458, 586, 268, 625, 61, 630, 672, 274, 27]. salmonids [469, 597, 114, 418, 328, 176, 344, 428, 105, 562, 295]. salmonines [91, 373]. salt [170, 455]. Salvelinus [687, 422, 401, 655, 540, 623, 516, 451, 210, 258, 87]. sample [188, 220]. samplers [613]. samples [613, 38]. sampling [519, 651, 514, 511, 69]. San [214]. sand [466, 283]. sandbar [454]. Sander [700, 47, 115]. Santa [447]. sapi [699]. sapidissima [690]. sapidus [191]. SARA [272]. SARA-listed [272]. sardine [557, 223, 583, 712]. SASI [231]. Saskatchewan [525]. satellite [104, 171]. scabbardfish [484]. scale [371, 327, 125, 689, 519, 226, 341, 667, 34, 609, 661, 301, 280, 274]. scale-dependent [609]. scale-derived [689]. scales [472, 93, 426]. scan [704, 89, 75]. scenarios [575, 5]. schedule [228]. Schindler [240]. schools [393, 437]. Science [272, 124, 444, 459, 304, 168, 709]. scientific [560, 619]. Scomber [579]. scombrus [579]. scorpionfish [694]. Scotia [423]. Scotian [94, 199]. sculpin [628]. Sea [400, 466, 573, 425, 165, 219, 118, 304, 679, 200, 341, 389, 206, 612, 635, 5, 366, 644, 510, 531, 467, 228, 8, 196, 716, 554, 464, 417, 481]. Sea-level [573]. Seabed [231, 137]. seal [703]. seals [508]. season [560, 143, 186, 11]. Seasonal [508, 360, 217, 595, 503]. seasonality [628]. seasons [434]. Seawater [440]. Sebastes [468, 142, 82]. sectors [111]. sediment [7, 202, 235]. sedimentary [294]. sediments [379]. seepage [455, 420]. segments [718]. seining [443]. select [242]. selecting [578, 542]. selection [31, 474, 365, 339, 540, 18, 332, 402, 305]. selective [327, 145, 681]. selectivity [432, 142, 244, 36, 638, 295, 22]. selenium [73]. self [682]. self-adjusting [682]. selfisher [432]. Semi [682, 67]. Semi-pelagic [682]. semi-wild [67]. seminatural [148]. sensing [365]. sensitivities [129]. Sensitivity [485]. sentinel [240]. separate [451, 257]. sequencing [110, 291, 236, 175]. series [299, 439, 559, 580, 376, 78]. seroprevalence [383]. settlement [413]. seven [497]. Seventy [676]. Sex [146, 700, 20, 197, 228, 653]. Sex-specific [146, 228, 653]. sexual [380, 450]. sexually [700]. shad [690, 14, 366, 699]. shallow [101, 455]. shape [474, 99, 59, 230]. shaped [671]. shapes [408]. shark [138, 476, 127, 448, 454]. sharks [360, 595]. Shelf [94, 199]. shell [364, 436]. Shield [188, 123, 218]. shift [138, 463]. shifting [430, 57]. shifts [439, 80, 64, 307, 442, 633, 109, 113, 335]. shiner [277, 280]. shoreline [657]. short [146, 574, 507]. short-lived [507]. short-term [146, 574]. shortened [402]. Shorter [18, 434]. shortly [662]. shrimp [18, 332, 361, 112, 4, 507]. side [704, 89, 25]. side-scan [704, 89]. signals [664]. signatures [395]. significance [561]. significant [607]. Silurus [482]. simplified [641]. Simulated [356, 172, 461, 68, 173, 662]. simulation [321, 385]. simulations [400, 415]. since [491]. single [110, 291, 342, 269, 175]. single-species [269]. siscowet [655, 540]. sisters [170]. site [365, 3]. situ [157, 35]. six [597]. Size [348, 553, 261, 469, 474, 188, 700, 437, 689, 142, 249, 145, 580, 211, 36, 224, 137, 144, 18, 332, 61, 89, 220, 344, 252, 399, 76, 311, 20, 403, 71, 539, 117, 562, 295, 373, 497, 394]. size-at-age [580, 403]. size-selective [145]. size-selectivity [295]. size-structured [137]. sizes [469, 396, 684]. skates [563]. skills [50]. slick [68]. Slope [400]. slow [103, 427]. Small [224, 709, 327, 710, 401, 200, 85, 670, 344, 612, 455, 717]. smallmouth [314, 27, 172]. smelt [323, 69]. smelter [294]. smolt [15, 649, 500, 40, 42, 394]. smolts [122, 678, 642, 502, 492, 238]. Snake [201, 642]. snapper [371, 457]. SNP [39, 115]. sockeye [157, 435, 262, 167, 81, 146, 30, 594, 580, 220, 433, 42]. soft [364]. soft-shell [364]. software [432]. softwater [455]. solutions [168]. Somatic [385, 152, 151, 603, 79]. Some [103, 168]. Somniosus [360]. sonar [704, 89]. sorting [66]. Sound [307, 193]. soundscape [85]. source [435, 432, 7, 283, 285]. sources [593, 613, 129]. south [473, 524]. southeastern [454]. Southern [160, 317, 614, 623, 139, 104, 163, 495, 384, 528, 489, 166, 115, 289, 659]. southernmost [235]. Southwest [636]. Southwestern [563]. sp [275]. Space [338, 607, 360, 710, 136, 29, 344, 290, 527, 79, 257, 591]. Spartina [202]. spasmodic [60]. Spatial [423, 547, 643, 528, 271, 152, 685, 218, 575, 215, 490, 451, 93, 136, 692, 232, 566, 386, 50, 100, 426, 565]. Spatially [370, 613, 658, 362, 711]. Spatiotemporal [134, 690, 683, 268, 91, 345, 54, 140, 397, 44, 151, 84, 367, 463, 5, 13, 119, 479, 177]. spatula [530]. spawn [174, 674]. spawned [377]. spawner [641, 532, 97, 320]. spawner-recruit [532]. spawners [180]. Spawning [186, 324, 691, 246, 560, 704, 262, 690, 365, 244, 339, 260, 594, 546, 116, 400, 493, 656, 646, 10, 410, 89, 164, 56, 64, 307, 356, 662, 166, 386, 562, 11, 426]. spawns [597]. specie [366]. Species [135, 650, 295, 639, 261, 272, 327, 425, 459, 541, 94, 199, 273, 375, 233, 628, 236, 190, 447, 268, 410, 208, 522, 643, 46, 14, 311, 637, 362, 588, 555, 111, 718, 113, 278, 274, 269, 13, 668, 562, 507, 11, 128, 321, 63, 52, 65, 273]. Species-specific [135, 295, 639]. specific [468, 146, 557, 135, 330, 347, 250, 14, 228, 295, 639, 653]. spectra [437, 348, 497]. spectroscopy [605]. spectrum [188]. Speed [193]. Sperm [212, 164]. spill [442, 633]. spills [68]. Spirinchus [323]. sponge [449]. sporadic [74]. sport [445]. spotted [623]. spp [246]. sprat [716]. Sprattus [716]. spread [356, 388]. spring [408, 642, 529, 324]. springtime [510]. Sr [267, 664, 267, 636]. Sr/ [267, 636]. Sr/Ca [664]. SSN [680]. SSURBA [136]. St [593, 614, 273, 104]. St. [701, 308, 280]. stabilize [65]. Stable [114, 702, 677, 508, 7, 557, 703, 316]. stage [461, 485, 359, 330, 389, 601, 504]. stage-based [461]. stages [620, 629, 654]. staircase [229]. standard [700]. State [29, 292, 377, 607, 363, 136, 290, 79, 257, 591]. State-space [29, 136, 79, 257, 591]. States [454]. stations [343]. Statistical [680, 432, 614, 685, 59, 537]. status [138, 32, 494, 610, 163, 161, 221, 398]. steelhead [86, 329, 24, 478, 201, 676, 229, 394, 462]. steep [494]. stenotherm [185]. still} [619]. stimuli [200]. Stock [271, 347, 461, 110, 125, 291, 607, 24, 683, 107, 143, 545, 456, 74, 141, 493, 72, 205, 415, 136, 220, 505, 124, 529, 324, 471, 523, 88, 697, 132, 528, 431, 269, 150, 386, 100, 511, 97, 591, 383, 221]. Stock-of-origin [271]. Stock-specific [347]. stocked [621, 53, 600]. stocking [535]. stocks [461, 125, 578, 60, 441, 59, 11, 414]. stomach [710]. storage [506, 462]. stores [515]. strain [148]. strains [305]. strategies [60, 638, 111, 639]. strategy [618, 708, 617, 717]. stratification [193]. stratified [38]. straying [596, 20]. Stream [469, 691, 261, 538, 262, 183, 519, 673, 422, 7, 131, 195, 200, 147, 182, 85, 93, 208, 344, 428, 229, 35, 442, 633, 259, 680, 71, 539, 562, 295]. stream-dwelling [428]. stream-resident [147]. streamflow [520]. streams [148, 572, 656, 646, 571, 524, 530, 207]. strength [621, 516, 344, 349, 54]. strengths [660]. stress [468, 331, 159]. stressed [413]. stressors [690, 279, 567]. striped [715]. strong [599]. stronghold [209]. strontium [222, 636]. structural [107]. structure [134, 425, 33, 501, 559, 677, 210, 109, 595, 301, 189, 149, 323, 100, 521, 175, 71, 513, 373]. structured [137, 106, 601]. structures [131, 172]. studies [613, 106]. study [453, 418, 141, 266, 376, 60, 713, 705, 269, 383, 63, 193, 345, 140, 177, 385, 414]. sturgeon [704, 485, 570, 315, 53, 410, 89, 354, 164, 58, 222, 386]. sturgeons [686]. sub [365]. sub-Arctic [365]. subadult [508, 338]. subarctic [440, 189]. subbasin [478]. submerged [393, 609]. subsidies [147, 182, 524]. subsidy [322]. substantial [632, 627]. subsurface [590]. success [86, 581, 264, 198, 180, 67, 116, 164, 543, 21, 394]. successful [390]. successive [581]. sucker [571, 206, 274, 275]. Sudbury [73, 294]. suitability [287, 127, 466, 692]. suite [221]. Summer [422, 94, 199, 79, 317, 228, 193]. superimposition [561, 562]. Superior [102, 655, 398]. supersaturation [354]. supplied [366]. supply [117]. Support [400, 272, 28, 409]. suppression [264, 121, 87]. Surface [208, 238, 425, 536]. surpass [534]. surplus [221]. surplus-production-based [221]. Surprising [288, 631, 630]. survey [560, 683, 94, 199, 84, 536, 613, 136, 637]. survey-based [136]. surveys [77, 519, 613, 443, 141, 336, 2, 611, 517, 537, 119, 398]. Survival [490, 413, 157, 329, 256, 160, 678, 212, 202, 211, 679, 367, 80, 185, 187, 669, 553, 330, 204, 475, 35, 502, 498, 427, 42, 467]. susceptibility [173]. sustainability [531]. sustainable [28, 601, 709, 436, 11]. SVM [588]. Swept [231]. Swimming [343, 67, 171, 247, 172]. sympatric [352, 295, 11]. symptoms [663]. Synchronous [233]. Synchrony [397, 580, 467]. syndrome [57]. synthesis [19, 388]. system [360, 158, 483, 53, 108, 465, 502, 301, 130]. system-wide [502]. systematic [686, 615]. systemic [719]. systems [28, 99, 600, 15].

T90 [142]. tactics [292]. tag [157, 98, 260, 674, 679, 415, 252, 1, 130]. tag-reliant [415]. tagged [678, 382]. tagging [125, 581, 201, 671, 50]. tags [104, 171, 697, 3, 130, 193]. tailwater [546]. Taking [376]. tale [438, 500, 96]. talking [518]. target [582, 349, 555, 343]. targeted [181]. taxonomy [236]. technique [144]. technology [355]. telemetric [611]. telemetry [444, 679, 483, 89, 15, 475, 338, 498, 293, 95, 3]. telemetry-based [483]. teleost [468, 595]. temperate [378, 460, 243]. Temperature [381, 261, 469, 520, 700, 425, 339, 219, 121, 209, 465, 222, 64, 35, 522, 326, 713, 235, 277, 76, 680, 597]. temperature-dependent [339, 64]. temperatures [673, 542, 58, 276, 250, 238, 357, 278, 539]. template [430]. Temporal [303, 358, 595, 471, 373, 423, 152, 218, 16, 250, 643, 528, 467]. temporally [127]. Tenacibaculum [435]. tench [308]. tendency [599]. term [77, 216, 244, 146, 195, 80, 185, 574, 380]. terms [412]. Territories [527]. test [227, 229]. textures [588]. Thaleichthys [175, 323]. their [461, 476, 170, 215, 219, 251, 586, 163, 176, 662, 471, 715]. theoretically [461]. Thermal [489, 684, 691, 243, 173, 144, 526, 276, 277, 695, 650, 285, 193, 503]. thermally [493]. thiamine [181]. thickens [242]. things [368]. Thompson [577, 672]. threat [328]. Threatened [275, 329, 490, 277, 280]. threats [643]. three [691, 541, 144, 615, 389, 34, 181, 588, 668, 22]. three-stage [389]. throughout [521]. throughput [236, 368]. Thunnus [271, 55, 205, 204, 588, 385]. thynnus [271, 205, 385]. tidal [329]. Tide [246]. tides [8, 269]. Time [634, 299, 439, 143, 559, 346, 580, 710, 367, 72, 376, 344, 543, 604, 78, 22]. time-varying [346, 72, 22]. timed [546]. timeframe [525]. times [64]. timing [408, 674, 553, 307]. Tinca [308]. tissue [23]. today [90]. tolerance [277, 489, 510, 285, 684]. tool [99]. tools [389, 132]. top [268, 668]. total [354]. Tournament [306]. track [567, 455, 420]. Tracking [581, 679, 400, 15, 671, 1]. trade [92, 283, 105]. trade-off [283]. trade-offs [92, 105]. tradeoffs [483]. trait [9]. trait-based [9]. traits [245, 477, 625, 235, 181, 471, 462]. trans [25]. trans-lateral [25]. transboundary [696]. transferability [692]. transform [605]. transgenic [399]. transient [509, 675]. transitions [241, 438]. translation [545]. translational [337]. translocation [283]. transmission [5]. transmitter [185]. transmontanus [354, 58]. transport [8, 116, 69]. transported [186]. transverse [639]. trap [173, 499]. trapping [341]. Trash [24]. trauma [354, 48, 302]. travel [367]. trawl [142, 682, 449, 251, 173, 361, 4, 402, 517]. trawling [137]. trawls [18, 332, 564]. treasure [24]. treat [634]. trees [393]. trend [163]. Trends [313, 468, 710, 136, 79, 427, 398]. tributaries [538, 701]. Tributary [530, 546, 61, 117]. tribute [240]. trigger [428]. triggerfish [3]. triggers [412]. triploid [58]. triploidy [350]. Trojan [687]. troll [626]. Trophic [655, 557, 237, 316, 91, 640, 668, 412, 659, 335]. tropical [496]. trout [261, 31, 148, 561, 86, 292, 234, 263, 102, 452, 687, 264, 242, 401, 304, 198, 180, 572, 382, 621, 516, 121, 546, 120, 209, 147, 610, 85, 596, 16, 534, 354, 411, 194, 276, 442, 633, 258, 352, 500, 310, 417, 695, 489, 95, 207, 43, 539, 87, 305, 599, 462]. truncata [364]. truncate [364]. trutta [292, 180, 596, 569, 310, 305, 599]. tshawytscha [553, 99, 347, 20, 57, 409]. tuna [271, 55, 582, 400, 205, 204, 342, 385]. tunicate [526]. tunnel [681]. turbulent [206]. Two [599, 11, 461, 579, 212, 622, 354, 326, 601, 14, 500, 96, 117, 562]. Two-generation [599]. two-stage-structured [601]. types [119].

U.S. [718]. unanticipated [663]. uncertainty [107, 651]. undatum [521]. Undercurrent [658]. underestimated [187]. understand [204, 671]. Understanding [48, 545, 128, 302]. understood [480, 260]. Unionidae [282]. unique [115]. unit [110, 291]. United [454]. unlikely [120]. unregulated [454]. Unveiling [391]. updated [131]. upon [242, 224, 168]. Upstream [180, 538, 581]. uptake [222]. upwelling [465, 82]. urban [216]. urbanization [51]. urchins [219]. USA [91, 502]. use [360, 686, 367, 53, 451, 108, 85, 176, 338, 232, 76, 352, 509, 675, 417, 504, 632, 627, 681, 230, 128, 426, 24]. user [49]. Using [689, 519, 575, 127, 339, 62, 567, 455, 704, 110, 291, 393, 453, 9, 264, 94, 199, 260, 370, 532, 202, 710, 629, 337, 231, 443, 116, 493, 236, 542, 703, 144, 415, 475, 191, 338, 308, 290, 498, 79, 441, 342, 132, 257, 59, 13, 130, 45, 175, 52, 479, 70, 372, 331]. utility [586]. utilized [548].

validates [169]. validating [217]. validation [55, 548, 576]. Validity [6]. Valley [676]. valuable [445]. value [582, 239, 266, 109, 659]. variability [474, 77, 151, 508, 460, 195, 557, 223, 490, 330, 526, 384, 573, 712, 385]. variable [437, 338, 574]. variables [179, 294]. variants [462]. Variation [2, 311, 134, 397, 689, 367, 139, 16, 411, 136, 347, 344, 544, 632, 627, 403, 285, 373, 412, 140, 177, 503]. variations [160, 54]. varied [362]. varies [121]. various [176, 294]. varying [346, 309, 613, 72, 658, 22]. vaterite [24]. vateritic [25]. vegetated [338]. velocities [172]. velocity [34]. ventilation [158]. versa [436]. versus [102, 236, 254, 464, 305]. vessel [536, 582, 251, 448]. vessel-based [536]. via [669, 514]. viability [202, 662, 688]. vice [436]. Victims [37]. viral [220]. virginica [436]. visual [641]. visualization [132]. visualize [107]. vitreus [700, 47, 115]. volume [695]. vulnerability [624, 563, 295, 482]. vulnerable [297, 279]. walleye [700, 391, 162, 404, 663, 522, 250, 47, 290, 464, 481, 426, 115]. warm [408, 615]. warm-water [408, 615]. Warmer [396]. Warming [464, 174, 408, 250, 695, 510, 684, 331]. warming-induced [331]. warranted [411]. wartmanni [244]. wastewater [364]. water [664, 520, 396, 300, 170, 218, 657, 360, 219, 408, 701, 243, 209, 288, 631, 586, 361, 615, 208, 590, 534, 379, 640, 35, 250, 630, 277, 153, 238, 357, 420, 275]. water-level [640]. waters [382, 621, 75, 240]. watershed [452, 577, 379, 442, 633, 258, 672, 149]. watersheds [51]. waterway [693]. Watson [217]. Wavyrayed [281]. way [107]. ways [176]. weather [671]. web [406, 68, 677]. webs [183, 137]. weight [473, 464, 479]. weight-at-age [464]. weighting [63, 22]. well [265]. West [718, 266, 626]. western [419, 463, 320, 521, 539, 513]. westslope [209, 411, 276]. wetlands [698]. whales [317, 513]. Whaling [714]. whelk [521]. Where [359, 597, 90, 355]. whether [33]. Whirling [328]. whisker [703]. white [138, 476, 570, 315, 623, 571, 354, 58, 206]. white-spotted [623]. whitefish [287, 244, 482, 398]. whitetip [448]. Who [597, 518]. wide [696, 502, 323, 510]. widely [484]. widespread [411]. width [242]. Wild [57, 148, 256, 506, 67, 621, 185, 669, 328, 194, 399, 552, 635, 356, 5, 644, 282, 616, 333]. wild-caught [333]. wild-reared [621]. wilderness [57]. Wildfire [539]. wildfires [189]. will [615, 523]. William [307]. wind [8]. window [691, 510]. Winnipeg [290]. winter [234, 67, 424, 247, 100]. Wintertime [296]. wire [382]. Wisconsin [386]. witch [683]. Within [20, 597, 313, 214, 447, 667, 527, 454, 697]. Within-river [20]. Wolastoq [379]. won’t [523]. workforce [651]. world [671, 640, 647].

Year [628, 621, 516, 525, 166]. year-around [166]. year-class [621, 516]. Year-round [628]. years [273, 676, 502, 149, 78, 385]. yellow [338, 504]. yellow-stage [504]. yelloweye [468]. yellowfin [55, 204]. Yellowstone [87, 261, 120, 87]. yield [566, 436]. young [508, 171, 349, 464]. young-adult [171]. Yukon [654, 159]. YY [264].

Zealand [541, 139, 34]. zone [443, 10]. zones [116]. Zooplankton [218, 188, 606, 536, 590, 301, 78, 372]. zooplankton-feeding [606]. Zostera [202].

References

Scheffel:2020:CAT

[1] Trevor K. Scheffel, Joseph E. Hightower, Jeffrey A. Buckel, Jacob R. Krause, and Frederick S. Scharf. Coupling acoustic tracking with conventional tag returns to estimate mortality for a coastal flatfish with high rates of emigration. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(1):1–22, January 2020. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/cjfas-
2018-0174.

Lance:2020:VIE

[2] Richard F. Lance and Xin Guan. Variation in inhibitor effects on qPCR assays and implications for eDNA surveys. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(1):23–33, January 2020. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/cjfas-2018-0263.

Shertzer:2020:EPA

[3] Kyle W. Shertzer, Nathan M. Bacheler, William E. Pine, Brendan J. Runde, Jeffrey A. Buckel, Paul J. Rudershausen, and Jamie H. MacMahan. Estimating population abundance at a site in the open ocean: combining information from conventional and telemetry tags with application to gray triggerfish (Balistes capriscus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(1):34–43, January 2020. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/doi/

full/10.1139/cjfas-2018-0356.

Lomeli:2020:EIR

[4] Mark J. M. Lomeli, Scott D. Groth, Matthew T. O. Blume, Bent Herrmann, and W. Waldo Wakefield. The efficacy of illumination to reduce bycatch of eulachon and groundfishes before trawl capture in the eastern North Pacific Ocean shrimp fishery. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(1):44–54, January 2020. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/doi/

full/10.1139/cjfas-2018-0497.

Peacock:2020:EST

[5] Stephanie J. Peacock, Martin Krkošek, Andrew W. Bateman, and Mark A. Lewis. Estimation of spatiotemporal transmission dynamics and analysis of management scenarios for sea lice of farmed and wild salmon. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(1):55–68, January 2020. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2019-0036.

Crane:2020:VAE

[6] Derek P. Crane, Marinda R. Cornett, Cory J. Bauerlien, Michael L. Hawkins, Daniel A. Isermann, Jeff L. Hansbarger, Kevin L. Kapuscinski, Jonathan R. Meerbeek, Timothy D. Simonson, and Jeffrey M. Kampa. Validity of age estimates from muskellunge (Esox masquinongy) fin rays and associated effects on estimates of growth. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(1):69–80, January 2020. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/cjfas-2018-0404.

Evans:2020:MSI

[7] Thomas M. Evans and Amy M. Weber. Multiple stable isotopes identify sediment organic matter as the primary nutritional source to American brook lamprey larvae along a stream gradient. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(1):81–89, January 2020. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/cjfas-2018-0270.

Deschepper:2020:BMR

[8] Inge Deschepper, Kieran Lyons, Olga Lyashevska, and Deirdre Brophy. Biophysical models reveal the role of tides, wind, and larval behaviour in early transport and retention of Atlantic herring (Clupea harengus) in the Celtic Sea. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(1):90–107, January 2020. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/cjfas-
2018-0491.

Cormier:2020:MTB

[9] Emily C. Cormier, Danielle R. Sisson, Kathleen M. Rühland, John P. Smol, and Joseph R. Bennett. A morphological trait-based approach to environmental assessment models using diatoms. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(1):108–112, January 2020. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/cjfas-2018-0376.

Irgens:2020:OGZ

[10] Christian Irgens, Arild Folkvord, Håkon Otterå, and Olav S. Kjesbu. Otolith growth and zone formation during first maturity and spawning of Atlantic cod (Gadus morhua). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(1):113–123, January 2020. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/doi/

full/10.1139/cjfas-2019-0033.

Vasconcelos:2020:TSE

[11] Joana Vasconcelos, Ricardo Sousa, Paulo Henriques, Antonieta Amorim, João Delgado, and Rodrigo Riera. Two sympatric, not externally discernible, and heavily exploited deepwater species with coastal migration during spawning season: implications for sustainable stocks management of Aphanopus carbo and Aphanopus intermedius around Madeira. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(1):124–131, January 2020. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2018-0423.

Cramer:2020:BCA

[12] Steven P. Cramer and Lucius Caldwell. Bias and consequences in attempts to estimate historical salmon abundance. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(1):132–145, January 2020. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/cjfas-2018-0467.

Stock:2020:CPF

[13] Brian C. Stock, Eric J. Ward, Tomoharu Eguchi, Jason E. Jannot, James T. Thorson, Blake E. Feist, and Brice X. Semmens. Comparing predictions of fisheries bycatch using multiple spatiotemporal species distribution model frameworks. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77 (1):146–163, January 2020. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2018-0281.

Nachon:2020:PSC

[14] David José Nachón, Gilles Bareille, Hilaire Drouineau, Hélène Tabouret, Catherine Taverny, Catherine Boisneau, Sylvain Berail, Christophe Pécheyran, Fanny Claverie, and Françoise Daverat. 1980s populationspecific compositions of two related anadromous shad species during the oceanic phase determined by microchemistry of archived otoliths. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(1):164–176, January 2020. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2018-0444.

Leander:2020:ONE

[15] Johan Leander, Jonatan Klaminder, Micael Jonsson, Tomas Brodin, Kjell Leonardsson, and Gustav Hellström. The old and the new: evaluating performance of acoustic telemetry systems in tracking migrating Atlantic salmon (Salmo salar) smolt and European eel (Anguilla anguilla) around hydropower facilities. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(1): 177–187, January 2020. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2019-0058.

Koeck:2020:BAF

[16] Barbara Koeck, Magnus Lovén Wallerius, Robert Arlinghaus, and Jörgen I. Johnsson. Behavioural adjustment of fish to temporal variation in fishing pressure affects catchability: an experiment with angled trout. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(1):188–193, January 2020. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2019-0064.

Kaemingk:2020:HRD

[17] Mark A. Kaemingk, Keith L. Hurley, Christopher J. Chizinski, and Kevin L. Pope. Harvest–release decisions in recreational fisheries. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(1):194–201, January 2020. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2019-0119.

Ingolfsson:2020:STI

[18] Ólafur Arnar Ingólfsson and Terje Jørgensen. Shorter trawls improve size selection of northern shrimp. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77 (1):202–211, January 2020. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2018-0443.

Humphries:2020:RRC

[19] Paul Humphries, Alison King, Nicole McCasker, R. Keller Kopf, Rick Stoffels, Brenton Zampatti, and Amina Price. Riverscape recruitment: a conceptual synthesis of drivers of fish recruitment in rivers. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(2):213–225, February 2020. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2018-0138.

Pollock:2020:WRS

[20] Amanda M. M. Pollock, Maryam Kamran, Andrew H. Dittman, Marc A. Johnson, and David L. G. Noakes. Within-river straying: sex and size influence recovery location of hatchery Chinook salmon (Oncorhynchus tshawytscha). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(2):226–235, February 2020. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2018-0384.

Rindorf:2020:PFR

[21] Anna Rindorf, Noel Cadigan, Daniel Howell, Margit Eero, and Henrik Gislason. Periodic fluctuations in recruitment success of Atlantic cod. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(2):236–246, February 2020. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2018-0496.

Xu:2020:CPT

[22] Haikun Xu, James T. Thorson, and Richard D. Methot, Jr. Comparing the performance of three data-weighting methods when allowing for time-varying selectivity. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(2): 247–263, February 2020. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2019-0107.

Lueders-Dumont:2020:CIC

[23] Jessica A. Lueders-Dumont, Daniel M. Sigman, Beverly J. Johnson, Olaf P. Jensen, Sergey Oleynik, and Bess B. Ward. Comparison of the isotopic composition of fish otolith-bound organic N with host tissue. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(2):264–275, February 2020. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2018-0360.

Budnik:2020:TTU

[24] Richard R. Budnik, John R. Farver, Joel E. Gagnon, and Jeffrey G. Miner. Trash or treasure? Use of sagittal otoliths partially composed of vaterite for hatchery stock discrimination in steelhead. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(2):276–284, February 2020. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/cjfas-2018-0387.

Vignon:2020:WPV

[25] Matthias Vignon. When the presence of a vateritic otolith has morphological effect on its aragonitic partner: trans-lateral compensation induces bias in microecological patterns in one-side-only vateritic otolith. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(2):285–294, February 2020. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2019-0066.

Munyandorero:2020:IPD

[26] Joseph Munyandorero. Inferring prior distributions of recruitment compensation metrics from life-history parameters and allometries. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(2):295–313, February 2020. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2018-0463.

Rubenson:2020:ISR

[27] Erika S. Rubenson and Julian D. Olden. An invader in salmonid rearing habitat: current and future distributions of smallmouth bass (Micropterus dolomieu) in the Columbia River Basin. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(2):314–325, February 2020. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/cjfas-2018-0357.

Dallaire:2020:MCC

[28] Camille Ouellet Dallaire, Bernhard Lehner, and Irena Creed. Multidisciplinary classification of Canadian river reaches to support the sustainable management of freshwater systems. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(2):326–341, February 2020. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/doi/

full/10.1139/cjfas-2018-0284.

Lecomte:2020:SSM

[29] Jean-Baptiste Lecomte, Olivier Le Pape, Hélène Baillif, Marie Nevoux, Youen Vermard, Marie Savina, Matthieu Veron, Sigrid Lehuta, Ewan Hunter, and Etienne Rivot. State-space modeling of multidecadal mark– recapture data reveals low adult dispersal in a nursery-dependent fish metapopulation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(2):342– 354, February 2020. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2019-0037.

Elliott:2020:KSS

[30] Lucas D. Elliott, Hillary G. M. Ward, and Michael A. Russello. Kokanee– sockeye salmon hybridization leads to intermediate morphology and resident life history: implications for fisheries management. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(2):355–364, February 2020. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2019-0034.

Alioravainen:2020:BEJ

[31] Nico Alioravainen, Pekka Hyvärinen, and Anssi Vainikka. Behavioural effects in juvenile brown trout in response to parental angling selection. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(2):365–374, February 2020. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2018-0424.

Eschenroeder:2020:HLF

[32] Jackman C. Eschenroeder and James H. Roberts. Habitat loss, fragmentation, and the genetic status of Roanoke bass. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(2):375–387, February 2020. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/cjfas-2019-0103.

Bowlby:2020:EWM

[33] Heather D. Bowlby and A. Jamie F. Gibson. Evaluating whether metapopulation structure benefits endangered diadromous fishes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(2):388–400, February 2020. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2019-0001.

McAllister:2020:RMS

[34] Tara G. McAllister, Susanna A. Wood, Emma M. MacKenzie, and Ian Hawes. Reach- and mat-scale differences in Microcoleus autumnalis (cyanobacterium) accrual along velocity and nitrate gradients in three New Zealand rivers. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(2): 401–412, February 2020. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2019-0133.

Lusardi:2020:OGS

[35] Robert A. Lusardi, Bruce G. Hammock, Carson A. Jeffres, Randy A. Dahlgren, and Joseph D. Kiernan. Oversummer growth and survival of juvenile Coho salmon (Oncorhynchus kisutch) across a natural gradient of stream water temperature and prey availability: an in situ enclosure experiment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(2):413–424, February 2020. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/cjfas-
2018-0484.

Gwinn:2020:BEI

[36] Daniel C. Gwinn, Gavin Butler, Brett Ingram, Scott Raymond, Mark Lintermans, and Qifeng Ye. Borrowing external information to estimate angler size selectivity: model development and application to Murray cod. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(2):425–437, February 2020. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2019-0045.

Germida:2020:MMV

[37] James J. Germida and Suzanne Kettley. In memoriam/in memoriam: Victims of Flight PS752. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(3), March 2020. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/cjfas-
2020-0049.

Perreault:2020:EGP

[38] Andrea M. J. Perreault, Nan Zheng, and Noel G. Cadigan. Estimation of growth parameters based on length-stratified age samples. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(3):439–450, March 2020. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2019-0129.

McKinney:2020:DSP

[39] Garrett J. McKinney, Carita E. Pascal, William D. Templin, Sara E. Gilk-Baumer, Tyler H. Dann, Lisa W. Seeb, and James E. Seeb. Dense SNP panels resolve closely related Chinook salmon populations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(3):451–461, March 2020. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2019-0067.

Pulkkinen:2020:BAM

[40] Henni Pulkkinen, Panu Orell, Jaakko Erkinaro, and Samu Mäntyniemi. Bayesian arrival model for Atlantic salmon smolt counts powered by environmental covariates and expert knowledge. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(3):462–474, March 2020. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/cjfas-2018-0352.

Slapansky:2020:RGM

[41] Luděk Šlapanský, Michal Janáč, Kevin Roche, and Pavel Jurajda. Round goby movement patterns in a non-navigable river. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(3):475–483, March 2020. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2018-0488.

Stevenson:2020:IAG

[42] Christine F. Stevenson, Arthur L. Bass, Nathan B. Furey, Kristina M. Miller, Shaorong Li, Erin L. Rechisky, Aswea D. Porter, David W. Welch, and Scott G. Hinch. Infectious agents and gene expression differ between sockeye salmon (Oncorhynchus nerka) smolt age classes but do not predict migration survival. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(3):484–495, March 2020. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/cjfas-
2019-0113.

Simard:2020:RED

[43] Lee G. Simard, J. Ellen Marsden, Robert E. Gresswell, and Megan Euclide. Rapid early development and feeding benefits an invasive population of lake trout. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(3):496–504, March 2020. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/cjfas-
2019-0122.

Cantrell:2020:RLB

[44] Danielle Cantrell, Ramón Filgueira, Crawford W. Revie, Erin E. Rees, Raphael Vanderstichel, Ming Guo, Michael G. G. Foreman, Di Wan, and Jon Grant. The relevance of larval biology on spatiotemporal patterns of pathogen connectivity among open-marine salmon farms. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(3):505–519, March 2020. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2019-0040.

Sullivan:2020:APM

[45] Christopher J. Sullivan, Daniel A. Isermann, Kaitlin E. Whitlock, and Jonathan F. Hansen. Assessing the potential to mitigate climate-related expansion of largemouth bass populations using angler harvest. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(3):520–533, March 2020. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2019-0035.

Meek:2020:IHB

[46] Mariah H. Meek, Molly R. Stephens, Alisha Goodbla, Bernie May, and Melinda R. Baerwald. Identifying hidden biocomplexity and genomic diversity in Chinook salmon, an imperiled species with a history of anthropogenic influence. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(3):534–547, March 2020. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/cjfas-
2019-0171.

May:2020:ILG

[47] Cassandra J. May, Stuart A. Ludsin, David C. Glover, and Elizabeth A. Marschall. The influence of larval growth rate on juvenile recruitment in Lake Erie walleye (Sander vitreus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(3):548–555, March 2020. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/doi/

full/10.1139/cjfas-2019-0059.

Pleizier:2020:UGB

[48] Naomi K. Pleizier, Charlotte Nelson, Steven J. Cooke, and Colin J. Brauner. Understanding gas bubble trauma in an era of hydropower expansion: how do fish compensate at depth? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(3):556–563, March 2020. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/cjfas-2019-0243. See correction [302].

Murphy:2020:GOM

[49] Christina A. Murphy, Chee Sing Lee, Brent Johnson, Ivan Arismendi, and Sherri L. Johnson. GrowChinook: an optimized multimodel and graphic user interface for predicting juvenile Chinook salmon growth in lentic ecosystems. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(3):564–575, March 2020. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/cjfas-
2018-0420.

Senina:2020:ITF

[50] Inna Senina, Patrick Lehodey, John Sibert, and John Hampton. Integrating tagging and fisheries data into a spatial population dynamics model to improve its predictive skills. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77 (3):576–593, March 2020. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2018-0470.

Pierce:2020:EUC

[51] Rita Monteiro Pierce, Karin E. Limburg, Daniella Hanacek, and Ivan Valiela. Effects of urbanization of coastal watersheds on growth and condition of juvenile alewives in New England. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(3):594–601, March 2020. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/doi/

full/10.1139/cjfas-2018-0434.

White:2020:PFS

[52] Shannon White, Evan Faulk, Caleb Tzilkowski, Andrew Weber, Matthew Marshall, and Tyler Wagner. Predicting fish species richness and habitat relationships using Bayesian hierarchical multispecies occupancy models. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(3):602–610, March 2020. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2019-0125.

Hegna:2020:MHU

[53] Jonathan Hegna, Kim Scribner, and Edward Baker. Movements, habitat use, and entrainment of stocked juvenile lake sturgeon in a hydroelectric reservoir system. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(3):611–624, March 2020. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/cjfas-
2018-0407.

Zhang:2020:SVJ

[54] Fan Zhang, Rick M. Rideout, and Noel G. Cadigan. Spatiotemporal variations in juvenile mortality and cohort strength of Atlantic cod (Gadus morhua) off Newfoundland and Labrador. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(3):625–635, March 2020. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/cjfas-2019-0156.

Andrews:2020:AVY

[55] Allen H. Andrews, Ashley Pacicco, Robert Allman, Brett J. Falterman, Erik T. Lang, and Walter Golet. Age validation of yellowfin (Thunnus albacares) and bigeye (Thunnus obesus) tuna of the northwestern Atlantic Oceano. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(4):637–643, April 2020. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/cjfas-
2019-0328.

Laurel:2020:LSH

[56] Benjamin J. Laurel and Lauren A. Rogers. Loss of spawning habitat and prerecruits of Pacific cod during a Gulf of Alaska heatwave. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(4):644–650, April 2020. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2019-0238.

Thurow:2020:WSS

[57] Russell F. Thurow, Timothy Copeland, and Bryce N. Oldemeyer. Wild salmon and the shifting baseline syndrome: application of archival and contemporary redd counts to estimate historical Chinook salmon (Oncorhynchus tshawytscha) production potential in the central Idaho wilderness. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(4):651–665, April 2020. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2019-0111.

Leal:2020:DTW

[58] Michaiah J. Leal, Joel P. Van Eenennaam, Andrea D. Schreier, and Anne E. Todgham. Diploid and triploid white sturgeon (Acipenser transmontanus) differ in magnitude but not kinetics of physiological responses to exhaustive exercise at ambient and elevated temperatures. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(4):666–673, April 2020. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2019-0289.

Smolinski:2020:APS

[59] Szymon Smoliński, Franziska Maria Schade, and Florian Berg. Assessing the performance of statistical classifiers to discriminate fish stocks using Fourier analysis of otolith shape. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(4):674–683, April 2020. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/doi/

full/10.1139/cjfas-2019-0251.

Licandeo:2020:MSS

[60] Roberto Licandeo, Daniel E. Duplisea, Caroline Senay, Julie R. Marentette, and Murdoch K. McAllister. Management strategies for spasmodic stocks: a Canadian Atlantic redfish fishery case study. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(4):684–702, April 2020. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2019-0210.

Kaylor:2020:CAI

[61] Matthew J. Kaylor, Seth M. White, Edwin R. Sedell, and Dana R. Warren. Carcass additions increase juvenile salmonid growth, condition, and size in an interior Columbia River Basin tributary. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(4):703–715, April 2020. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/cjfas-2019-0215.

Mace:2020:UCR

[62] Marvin M. Mace III and Michael J. Wilberg. Using censored regression when estimating abundance with CPUE data to account for daily catch limits. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(4):716–722, April 2020. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/cjfas-
2019-0093.

Wang:2020:OWS

[63] Jintao Wang, Robert Boenish, and Xinjun Chen. Optimal weighting in species habitat modeling: a case study from Ommastrephes bartramii in the Northwest Pacific Ocean. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77 (4):723–740, April 2020. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2019-0204.

Lombardo:2020:ETD

[64] Steven M. Lombardo, Jeffrey A. Buckel, Ernie F. Hain, Emily H. Griffith, and Holly White. Evidence for temperature-dependent shifts in spawning times of anadromous alewife (Alosa pseudoharengus) and blueback herring (Alosa aestivalis). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(4):741–751, April 2020. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/cjfas-
2019-0140.

vanRijn:2020:ASS

[65] Itai van Rijn, Moshe Kiflawi, and Jonathan Belmaker. Alien species stabilize local fisheries catch in a highly invaded ecosystem. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(4):752–761, April 2020. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2019-0065.

Dassow:2020:EDC

[66] Colin J. Dassow, Alexander J. Ross, Olaf P. Jensen, Greg G. Sass, Brett T. van Poorten, Christopher T. Solomon, and Stuart E. Jones. Experimental demonstration of catch hyperstability from habitat aggregation, not effort sorting, in a recreational fishery. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(4):762–769, April 2020. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/doi/

full/10.1139/cjfas-2019-0245.

Hatanpaa:2020:EWF

[67] Aurora Hatanpää, Hannu Huuskonen, Jukka Kekäläinen, Raine Kortet, Pekka Hyvärinen, Maria Letizia Vitelletti, and Jorma Piironen. Early winter foraging success, swimming performance, and morphology of juvenile landlocked Atlantic salmon reared under semi-wild and hatchery conditions. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(4):770–778, April 2020. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/cjfas-
2019-0079.

Cederwall:2020:LUO

[68] Jeffrey Cederwall, Tyler A. Black, Jules M. Blais, Mark L. Hanson, Bruce P. Hollebone, Vince P. Palace, José Luis Rodríguez-Gil, Charles W. Greer, Christine Maynard, Alice C. Ortmann, Rebecca C. Rooney, and Diane M. Orihel. Life under an oil slick: response of a freshwater food web to simulated spills of diluted bitumen in field mesocosms. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(5):779–788, May 2020. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2019-0224.

Smith:2020:BHM

[69] William E. Smith, Ken B. Newman, and Lara Mitchell. A Bayesian hierarchical model of postlarval delta smelt entrainment: integrating transport, length composition, and sampling efficiency in estimates of loss. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(5):789–813, May 2020. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2019-0148.

Yuan:2020:MHD

[70] Lester L. Yuan and John R. Jones. Modeling hypolimnetic dissolved oxygen depletion using monitoring data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(5):814–823, May 2020. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/doi/

full/10.1139/cjfas-2019-0294.

Sutton:2020:MFC

[71] Isaac A. Sutton and Nicholas E. Jones. Measures of fish community size structure as indicators for stream monitoring programs. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(5):824–835, May 2020. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2019-0268.

Holt:2020:ITV

[72] Carrie A. Holt and Catherine G. J. Michielsens. Impact of time-varying productivity on estimated stock–recruitment parameters and biological reference points. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(5):836–847, May 2020. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/cjfas-
2019-0104.

Azdajic:2020:EDS

[73] Mija Aždajić, Nelson Belzile, John M. Gunn, Jules M. Blais, and Alexandre J. Poulain. Effects of a decade of selenium emission reductions on mercury accumulation in aquatic biota in the Sudbury region of Ontario. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(5):848–856, May 2020. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2019-0196.

Guan:2020:IDL

[74] Lisha Guan, Yong Chen, Robert Boenish, Xianshi Jin, and Xiujuan Shan. Improving data-limited stock assessment with sporadic stock index information in stock reduction analysis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(5):857–868, May 2020. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/doi/

full/10.1139/cjfas-2018-0500.

Perez-Jvostov:2020:HSC

[75] Felipe Pérez-Jvostov, William J. Sutherland, Rowan D. H. Barrett, Catherine A. Brown, Jeffrey A. Cardille, Steven J. Cooke, Melania E. Cristescu, Nicolas Fortin St-Gelais, Gregor F. Fussmann, Katherine Griffiths, Andrew P. Hendry, Nicolas W. R. Lapointe, Elizabeth A. Nyboer, Ralph L. Pentland, Andrea J. Reid, Anthony Ricciardi, Jennifer M. Sunday, and Irene Gregory-Eaves. Horizon scan of conservation issues for inland waters in Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77 (5):869–881, May 2020. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2019-0105.

Mulder:2020:MTD

[76] Ingeborg M. Mulder, Corey J. Morris, J. Brian Dempson, Ian A. Fleming, and Michael Power. Marine temperature and depth use by anadromous Arctic char correlates to body size and diel period. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(5):882–893, May 2020. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2019-0097.

Cauvy-Fraunie:2020:IIV

[77] Sophie Cauvy-Fraunié, Verena M. Trenkel, Martin Daufresne, Anthony Maire, Hervé Capra, Jean-Michel Olivier, Jérémy Lobry, Bernard Cazelles, and Nicolas Lamouroux. Interpretation of interannual variability in longterm aquatic ecological surveys. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(5):894–903, May 2020. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/doi/

full/10.1139/cjfas-2019-0146.

Schindler:2020:CPZ

[78] Eva U. Schindler, Bahman Shafii, Paul J. Anders, William J. Price, Charlie Holderman, Ken I. Ashley, and Marley Bassett. Characterizing the phytoplankton and zooplankton communities in Kootenay Lake: a time series analysis of 24 years of nutrient addition. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(5):904–916, May 2020. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/doi/

full/10.1139/cjfas-2018-0429.

OLeary:2020:DSG

[79] Cecilia A. O’Leary, Christine C. Stawitz, and Janet A. Nye. Detecting somatic growth trends for summer flounder (Paralichthys dentatus) using a state-space approach. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(5):917–930, May 2020. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/cjfas-
2019-0213.

Harbicht:2020:MST

[80] Andrew B. Harbicht, Dylan J. Fraser, and William R. Ardren. Minor shifts towards more natural conditions in captivity improve long-term survival among reintroduced Atlantic salmon. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(5):931–942, May 2020. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/doi/

full/10.1139/cjfas-2019-0201.

Connors:2020:CCI

[81] Brendan Connors, Michael J. Malick, Gregory T. Ruggerone, Pete Rand, Milo Adkison, James R. Irvine, Robert Campbell, and Kristen Gorman. Climate and competition influence sockeye salmon population dynamics across the Northeast Pacific Ocean. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(6):943–949, June 2020. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/doi/

full/10.1139/cjfas-2019-0422. See correction [167].

Markel:2020:CEC

[82] Russell W. Markel and Jonathan B. Shurin. Contrasting effects of coastal upwelling on growth and recruitment of nearshore Pacific rockfishes (genus Sebastes). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(6):950–962, June 2020. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/cjfas-
2019-0179.

Dempsey:2020:ANN

[83] Danielle P. Dempsey, Pierre Pepin, Mariano Koen-Alonso, and Wendy C. Gentleman. Application of neural networks to model changes in fish community biomass in relation to pressure indicators and comparison with a linear approach. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(6):963–977, June 2020. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/cjfas-
2018-0411.

Dunn:2020:FSD

[84] Corey G. Dunn and Craig P. Paukert. A flexible survey design for monitoring spatiotemporal fish richness in nonwadeable rivers: optimizing efficiency by integrating gears. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77 (6):978–990, June 2020. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2019-0315.

Kacem:2020:RMS

[85] Zaccaria Kacem, Marco A. Rodríguez, Irene T. Roca, and Raphaël Proulx. The riverscape meets the soundscape: acoustic cues and habitat use by brook trout in a small stream. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77 (6):991–999, June 2020. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2019-0311.

Berejikian:2020:DCA

[86] Barry A. Berejikian, Christopher P. Tatara, Donald M. Van Doornik, Michael A. Humling, Matt R. Cooper, Chris R. Pasley, and Jeffrey J. Atkins. Duration in captivity affects competitive ability and breeding success of male but not female steelhead trout (Oncorhynchus mykiss). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(6):1000–1009, June 2020. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2019-0255.

Syslo:2020:CER

[87] John M. Syslo, Travis O. Brenden, Christopher S. Guy, Todd M. Koel, Patricia E. Bigelow, Philip D. Doepke, Jeffrey L. Arnold, and Brian D. Ertel. Could ecological release buffer suppression efforts for non-native lake trout (Salvelinus namaycush) in Yellowstone Lake, Yellowstone National Park? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(6):1010–1025, June 2020. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2019-0306.

Pons:2020:CPC

[88] Maite Pons, Jason M. Cope, and Laurence T. Kell. Comparing performance of catch-based and length-based stock assessment methods in datalimited fisheries. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(6):1026–1037, June 2020. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/cjfas-
2019-0276.

Kazyak:2020:ISS

[89] David C. Kazyak, Amy M. Flowers, Nathan J. Hostetter, John A. Madsen, Matthew Breece, Amanda Higgs, Lori M. Brown, J. Andrew Royle, and Dewayne A. Fox. Integrating side-scan sonar and acoustic telemetry to estimate the annual spawning run size of Atlantic sturgeon in the Hudson River. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(6):1038–1048, June 2020. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/cjfas-
2019-0398.

Boenish:2020:REA

[90] Robert Boenish and Yong Chen. Re-evaluating Atlantic cod mortality including lobster bycatch: where could we be today? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(6):1049–1058, June 2020. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2019-0313.

Kornis:2020:SPT

[91] Matthew S. Kornis, David B. Bunnell, Heidi K. Swanson, and Charles R. Bronte. Spatiotemporal patterns in trophic niche overlap among five salmonines in Lake Michigan, USA. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(6):1059–1075, June 2020. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/doi/

full/10.1139/cjfas-2019-0288.

Connors:2020:IHP

[92] Brendan M. Connors, Benjamin Staton, Lewis Coggins, Carl Walters, Mike Jones, Daniel Gwinn, Matt Catalano, and Steve Fleischman. Incorporating harvest–population diversity trade-offs into harvest policy analyses of salmon management in large river basins. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(6):1076–1089, June 2020. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/cjfas-2019-0282.

Kanno:2020:MMR

[93] Yoichiro Kanno, Naoki Yui, Wataru Mamiya, Rei Sakai, Yuri Yabuhara, Tohru Miyazaki, Shunsuke Utsumi, Osamu Kishida, and Hiromi Uno. A multistate mark–recapture approach to characterize stream fish movement at multiple spatial scales. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77 (6):1090–1100, June 2020. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2019-0329.

Debertin:2020:EBM

[94] Allan J. Debertin. Estimating the biomass of a mixed species complex using hydroacoustics and catch data from the Bay of Fundy and Scotian Shelf summer ecosystem survey. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(7):1101–1116, July 2020. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/doi/

full/10.1139/cjfas-2019-0152. See correction [199].

Roloson:2020:OMA

[95] Scott D. Roloson, Sean J. Landsman, Raymond Tana, Brendan J. Hicks, Jon W. Carr, Fred Whoriskey, and Michael R. van den Heuvel. Otolith microchemistry and acoustic telemetry reveal anadromy in non-native rainbow trout (Oncorhynchus mykiss) in Prince Edward Island, Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(7):1117–1130, July 2020. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2019-0229.

Stone:2020:HIP

[96] Joshua P. Stone, Kevin L. Pangle, Steven A. Pothoven, Henry A. Vanderploeg, Stephen B. Brandt, Tomas O. Höök, Thomas H. Johengen, and Stuart A. Ludsin. Hypoxia’s impact on pelagic fish populations in Lake Erie: a tale of two planktivores. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77 (7):1131–1148, July 2020. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2019-0265.

Staton:2020:EMS

[97] Benjamin A. Staton, Matthew J. Catalano, Brendan M. Connors, Lewis G. Coggins, Michael L. Jones, Carl J. Walters, Steven J. Fleischman, and Daniel C. Gwinn. Evaluation of methods for spawner–recruit analysis in mixed-stock Pacific salmon fisheries. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(7):1149–1162, July 2020. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/doi/

full/10.1139/cjfas-2019-0281.

Conner:2020:EPA

[98] Mary M. Conner, Phaedra E. Budy, Richard A. Wilkison, Michael Mills, David Speas, Peter D. Mackinnon, and Mark C. Mckinstry. Estimating population abundance with a mixture of physical capture and PIT tag antenna detection data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77 (7):1163–1171, July 2020. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2019-0326.

Koeberle:2020:OSC

[99] Alexander L. Koeberle, Ivan Arismendi, Whitney Crittenden, Cecilia Di Prinzio, Daniel Gomez-Uchida, David L. G. Noakes, and Shannon Richardson. Otolith shape as a classification tool for Chinook salmon (Oncorhynchus tshawytscha) discrimination in native and introduced systems. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(7):1172–1188, July 2020. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2019-0280.

Siskey:2020:RSP

[100] Matthew R. Siskey, Michael G. Frisk, Robert M. Cerrato, and Karin E. Limburg. Redefining spatial population structure of winter flounder (Pseudopleuronectes americanus): implications for stock assessment and management. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(7):1189–1200, July 2020. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/cjfas-
2019-0279.

Harrow-Lyle:2020:IMN

[101] Tyler Harrow-Lyle and Andrea E. Kirkwood. The invasive macrophyte Nitellopsis obtusa may facilitate the invasive mussel Dreissena polymorpha and Microcystis blooms in a large, shallow lake. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(7):1201–1208, July 2020. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/cjfas-2019-0337.

Chavarie:2020:DVN

[102] L. Chavarie, J. Hoffmann, A. M. Muir, C. C. Krueger, C. R. Bronte, K. L. Howland, C. P. Gallagher, S. P. Sitar, M. J. Hansen, M. R. Vinson, L. F. Baker, L. L. Loseto, W. M. Tonn, and H. K. Swanson. Dietary versus nondietary fatty acid profiles of lake trout ecotypes from Lake Superior and Great Bear Lake: Are fish really what they eat? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(7):1209–1220, July 2020. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2019-0343.

Carmichael:2020:SLI

[103] Richard A. Carmichael, Daniele Tonina, Ernest R. Keeley, Rohan M. Benjankar, and Kevin E. See. Some like it slow: a bioenergetic evaluation of habitat quality for juvenile Chinook salmon in the Lemhi River, Idaho. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(7):1221–1232, July 2020. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2019-0136.

James:2020:MPA

[104] Travis L. James, Sean J. Landsman, Laura L. Ramsay, Melanie D. Giffin, Arnault Le Bris, and Michael R. van den Heuvel. Migration patterns of Atlantic halibut captured in the southern Gulf of St. Lawrence as assessed with pop-up satellite archival and floy tags. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(7):1233–1242, July 2020. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/cjfas-2019-0262.

Rosenfeld:2020:ATO

[105] Jordan Rosenfeld, Jeff Richards, Dave Allen, Travis Van Leeuwen, and Gauthier Monnet. Adaptive trade-offs in fish energetics and physiology: insights from adaptive differentiation among juvenile salmonids. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(7):1243–1255, July 2020. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2019-0350.

Koehn:2020:CSD

[106] Laura E. Koehn, Timothy E. Essington, Phillip S. Levin, Kristin N. Marshall, Lee G. Anderson, Alida Bundy, Courtney Carothers, Felicia Coleman, Jonathan H. Grabowski, Edward Houde, Olaf P. Jensen, Christian Möllmann, and Anthony D. M. Smith. Case studies demonstrate capacity for a structured planning process for ecosystem-based fisheries management. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(7):1256–1274, July 2020. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/cjfas-
2019-0202.

Cope:2020:NWV

[107] Jason Cope and Vladlena Gertseva. A new way to visualize and report structural and data uncertainty in stock assessments. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(8):1275–1280, August 2020. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2020-0082.

Hubbell:2020:FAE

[108] Joshua P. Hubbell, Jake F. Schaefer, Peter Flood, Melvin L. Warren, and Ken A. Sterling. Fragmentation alters ecological gradients and headwater fish assemblage composition relative to land use in a dendritic river system. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(8):1281–1291, August 2020. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/cjfas-
2019-0080.

Ohlberger:2020:RVL

[109] Jan Ohlberger, Daniel E. Schindler, Randy J. Brown, Joel M. S. Harding, Milo D. Adkison, Andrew R. Munro, Lara Horstmann, and Joe Spaeder. The reproductive value of large females: consequences of shifts in demographic structure for population reproductive potential in Chinook salmon. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(8):1292–1301, August 2020. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/cjfas-
2020-0012.

Beacham:2020:AEC

[110] Terry D. Beacham, Colin Wallace, Kim Jonsen, Brenda McIntosh, John R. Candy, Eric B. Rondeau, Jean-Sébastien Moore, Louis Bernatchez, and Ruth E. Withler. Accurate estimation of conservation unit contribution to Coho salmon mixed-stock fisheries in British Columbia, Canada, using direct DNA sequencing for single nucleotide polymorphisms. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(8):1302–1315, August 2020. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2019-0339.

Punt:2020:EMS

[111] André E. Punt, Margaret Siple, Gudjón Már Sigurdsson, Gísli Víkingsson, Tessa B. Francis, Sandra M. Granquist, Philip S. Hammond, Dennis Heinemann, Kristy J. Long, Jeffrey E. Moore, Maritza Sepúlveda, Randall R. Reeves, Paul R. Wade, Rob Williams, and Alexandre N. Zerbini. Evaluating management strategies for marine mammal populations: an example for multiple species and multiple fishing sectors in Iceland. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(8):1316–1331, August 2020. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2019-0386.

LeCorre:2020:LCN

[112] Nicolas Le Corre, Pierre Pepin, AnnDorte Burmeister, Wojciech Walkusz, Katherine Skanes, Zeliang Wang, David Brickman, and Paul V. R. Snelgrove. Larval connectivity of northern shrimp (Pandalus borealis) in the Northwest Atlantic. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(8): 1332–1347, August 2020. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2019-0454.

Rogosch:2020:IIC

[113] Jane S. Rogosch and Julian D. Olden. Invaders induce coordinated isotopic niche shifts in native fish species. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(8):1348–1358, August 2020. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/doi/

full/10.1139/cjfas-2019-0346.

Cucherousset:2020:SIN

[114] Julien Cucherousset, Libor Závorka, Sergine Ponsard, Régis Céréghino, and Frédéric Santoul. Stable isotope niche convergence in coexisting native and non-native salmonids across age classes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(8):1359–1365, August 2020. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/cjfas-2019-0186.

Zhao:2020:SAH

[115] Honggang Zhao, Katherine Silliman, Matthew Lewis, Sarah Johnson, Garret Kratina, Steve J. Rider, Carol A. Stepien, Eric M. Hallerman, Benjamin Beck, Adam Fuller, and Eric Peatman. SNP analyses highlight a unique, imperiled southern walleye (Sander vitreus) in the Mobile River Basin. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(8):1366–1378, August 2020. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/cjfas-
2019-0351.

Heer:2020:MGC

[116] Tej Heer, Mathew G. Wells, P. Ryan Jackson, and Nicholas E. Mandrak. Modelling grass carp egg transport using a 3-D hydrodynamic river model: the role of egg retention in dead zones on spawning success. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(8):1379–1392, August 2020. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2019-0344.

Tavernini:2020:ETS

[117] David A. Tavernini and John S. Richardson. Effects of tributary size on the resource supply and physical habitat at tributary junctions along two mainstem rivers. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(8): 1393–1408, August 2020. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2019-0435.

Forrest:2020:ERA

[118] Robyn E. Forrest, Ian J. Stewart, Cole C. Monnahan, Katherine H. Bannar-Martin, and Lisa C. Lacko. Evidence for rapid avoidance of rockfish habitat under reduced quota and comprehensive at-sea monitoring in the British Columbia Pacific halibut fishery. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(8):1409–1420, August 2020. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/cjfas-2019-0444.

Webster:2020:MCD

[119] Raymond A. Webster, Eric Soderlund, Claude L. Dykstra, and Ian J. Stewart. Monitoring change in a dynamic environment: spatiotemporal modelling of calibrated data from different types of fisheries surveys of Pacific halibut. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(8):1421–1432, August 2020. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/cjfas-
2019-0240.

Heim:2020:ACU

[120] Kurt C. Heim, Thomas E. McMahon, Steven T. Kalinowski, Brian D. Ertel, and Todd M. Koel. Abiotic conditions are unlikely to mediate hybridization between invasive rainbow trout and native Yellowstone cutthroat trout in a high-elevation metapopulation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(9):1433–1445, September 2020. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2019-0317.

Healy:2020:RRN

[121] Brian D. Healy, Robert C. Schelly, Charles B. Yackulic, Emily C. Omana Smith, and Phaedra Budy. Remarkable response of native fishes to invasive trout suppression varies with trout density, temperature, and annual hydrology. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(9):1446– 1462, September 2020. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2020-0028.

Babin:2020:MAS

[122] Amanda B. Babin, Mouhamed Ndong, Katy Haralampides, Stephan Peake, Ross Jones, R. Allen Curry, and Tommi Linnansaari. Migration of Atlantic salmon (Salmo salar) smolts in a large hydropower reservoir. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(9):1463–1476, September 2020. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2019-0395.

Chapelsky:2020:ICN

[123] Andrew J. Chapelsky, Matthew M. Guzzo, Lee E. Hrenchuk, and Paul J. Blanchfield. Invertebrate colonization of a newly constructed diversion channel in the Canadian Shield. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(9):1477–1486, September 2020. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/cjfas-2020-0026.

Munsch:2020:SIM

[124] Stuart H. Munsch, Correigh M. Greene, Rachel C. Johnson, William H. Satterthwaite, Hiroo Imaki, Patricia L. Brandes, and Michael R. O’Farrell. Science for integrative management of a diadromous fish stock: interdependencies of fisheries, flow, and habitat restoration. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(9):1487–1504, September 2020. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2020-0075.

Beacham:2020:LSP

[125] Terry D. Beacham, Kim Jonsen, Brenda McIntosh, Ben J. G. Sutherland, David Willis, Cheryl Lynch, and Colin Wallace. Large-scale parentagebased tagging and genetic stock identification applied in assessing mixedstock fisheries and hatchery brood stocks for Coho salmon in British Columbia, Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(9): 1505–1517, September 2020. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2020-0035.

VanLeeuwen:2020:MAS

[126] Travis E. Van Leeuwen, J. Brian Dempson, Chantelle M. Burke, Nicholas I. Kelly, Martha J. Robertson, Robert J. Lennox, Torgeir B. Havn, Martin Svenning, Ross Hinks, Matthew M. Guzzo, Eva B. Thorstad, Craig F. Purchase, and Amanda E. Bates. Mortality of Atlantic salmon after catch and release angling: assessment of a recreational Atlantic salmon fishery in a changing climate. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(9):1518–1528, September 2020. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2019-0400.

DeWysiecki:2020:UTE

[127] Agustín María De Wysiecki, Noela Sánchez-Carnero, Alejo Joaquín Irigoyen, Andrés Conrado Milessi, Jorge Horacio Colonello, Nelson Darío Bovcon, Federico Cortés, Santiago Aldo Barbini, Paula Victoria Cedrola, Nidia Marina Coller, and Andrés Javier Jaureguizar. Using temporally explicit habitat suitability models to infer the migratory pattern of a large mobile shark. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(9): 1529–1539, September 2020. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2020-0036.

Wagner:2020:IUP

[128] Tyler Wagner, Gretchen J. A. Hansen, Erin M. Schliep, Bethany J. Bethke, Andrew E. Honsey, Peter C. Jacobson, Benjamen C. Kline, and Shannon L. White. Improved understanding and prediction of freshwater fish communities through the use of joint species distribution models. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(9):1540–1551, September 2020. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2019-0348.

Smolinski:2020:PSB

[129] Szymon Smoliński, John Morrongiello, Peter van der Sleen, Bryan A. Black, and Steven E. Campana. Potential sources of bias in the climate sensitivities of fish otolith biochronologies. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(9):1552–1563, September 2020. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2019-0450.

Stout:2020:KIC

[130] J. Benjamin Stout, Mary Conner, Phaedra Budy, Peter Mackinnon, and Mark McKinstry. Keeping it classy: classification of live fish and ghost PIT tags detected with a mobile PIT tag interrogation system using an innovative analytical approach. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(9): 1564–1573, September 2020. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2019-0403.

Foote:2020:ISR

[131] Kyleisha J. Foote, Pascale M. Biron, and James W. A. Grant. Impact of in-stream restoration structures on salmonid abundance and biomass: an updated meta-analysis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(9): 1574–1591, September 2020. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2019-0327.

Regular:2020:ICA

[132] Paul M. Regular, Gregory J. Robertson, Robert Rogers, and Keith P. Lewis. Improving the communication and accessibility of stock assessment using interactive visualization tools. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(9):1592–1600, September 2020. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/cjfas-2019-0424.

Benjamin:2020:CNA

[133] Joseph R. Benjamin, J. Ryan Bellmore, Emily Whitney, and Jason B. Dunham. Can nutrient additions facilitate recovery of Pacific salmon? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(10):1601–1611, October 2020. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2019-0438.

Barton:2020:SVE

[134] Mark B. Barton, Johanna J. Vollenweider, Ron A. Heintz, Brenda L. Norcross, and Kevin M. Boswell. Spatiotemporal variation of environmental conditions and prey availability that drive Arctic nearshore fish community structure in the Point Barrow, Alaska, region. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(10):1612–1624, October 2020. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2019-0068.

Jarvis:2020:SSP

[135] Lauren A. Jarvis, Bailey C. McMeans, Henrique Corrêa Giacomini, and Cindy Chu. Species-specific preferences drive the differential effects of lake factors on fish production. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(10): 1625–1637, October 2020. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2020-0105.

Kumar:2020:SSS

[136] Rajeev Kumar, Noel G. Cadigan, Nan Zheng, Divya A. Varkey, and M. Joanne Morgan. A state-space spatial survey-based stock assessment (SSURBA) model to inform spatial variation in relative stock trends. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(10):1638–1658, October 2020.

CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2019-0427.

Howarth:2020:ETP

[137] Leigh M. Howarth, Paul J. Somerfield, Julia L. Blanchard, James J. Waggitt, Susan Allender, and Jan G. Hiddink. The effects of trawling and primary production on size-structured food webs in seabed ecosystems. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(10):1659–1665, October 2020. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2020-0025.

Bastien:2020:IRN

[138] G. Bastien, A. Barkley, J. Chappus, V. Heath, S. Popov, R. Smith, T. Tran, S. Currier, D. C. Fernandez, P. Okpara, V. Owen, B. Franks, R. Hueter, D. J. Madigan, C. Fischer, B. McBride, and N. E. Hussey. Inconspicuous, recovering, or northward shift: status and management of the white shark (Carcharodon carcharias) in Atlantic Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(10):1666–1677, October 2020. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2020-0055.

Ingram:2020:PEG

[139] Travis Ingram, Ludovic Dutoit, Pavel Mikheev, Samiullah Khan, and Marc Schallenberg. Phenotypic, ecological, and genomic variation in common bully (Gobiomorphus cotidianus) populations along depth gradients in New Zealand’s southern Great Lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(10):1678–1687, October 2020. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2020-0015.

Zheng:2020:SVM

[140] Nan Zheng, Matthew Robertson, Noel Cadigan, Fan Zhang, Joanne Morgan, and Laura Wheel. Spatiotemporal variation in maturation: a case study with American plaice (Hippoglossoides platessoides) on the Grand Bank off Newfoundland. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(10): 1688–1699, October 2020. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2020-0017. See correction [177].

Hodgdon:2020:FIE

[141] Cameron T. Hodgdon, Kisei R. Tanaka, Jocelyn Runnebaum, Jie Cao, and Yong Chen. A framework to incorporate environmental effects into stock assessments informed by fishery-independent surveys: a case study with American lobster (Homarus americanus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(10):1700–1710, October 2020. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2020-0076.

Cheng:2020:ONT

[142] Zhaohai Cheng, Paul D. Winger, Shannon M. Bayse, Gebremeskel Eshetu Kebede, Harold DeLouche, Haraldur Arnar Einarsson, Michael V. Pol, David Kelly, and Stephen J. Walsh. Out with the old and in with the new: T90 codends improve size selectivity in the Canadian redfish (Sebastes mentella) trawl fishery. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(10): 1711–1720, October 2020. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2020-0063.

DeFilippo:2020:BHA

[143] Lukas B. DeFilippo, Daniel E. Schindler, Kyle Shedd, and Kevin L. Schaberg. A Bayesian hierarchical approach to integrating historical and in-season genetic data for real-time assessment of a mixed stock fishery. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(10):1721–1732, October 2020. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2020-0092.

Huntsberger:2020:ASR

[144] Carl J. Huntsberger, Raouf Kilada, William G. Ambrose, and Richard A. Wahle. Age-at-size relationships of the American lobster (Homarus americanus) from three contrasting thermal regimes using gastric mill band counts as a direct aging technique. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(10):1733–1740, October 2020. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/cjfas-2020-0018.

Evangelista:2020:DDC

[145] Charlotte Evangelista, Leif Asbjørn Vøllestad, Beatriz Diaz Pauli, and Eric Edeline. Density-dependent consequences of size-selective induced life-history changes to population fitness in medaka (Oryzias latipes). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(10):1741–1748, October 2020. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2019-0406.

Eliason:2020:SSD

[146] Erika J. Eliason, Melissa Dick, David A. Patterson, Kendra A. Robinson, Jeremy Lotto, Scott G. Hinch, and Steven J. Cooke. Sex-specific differences in physiological recovery and short-term behaviour following fisheries capture in adult sockeye salmon (Oncorhynchus nerka). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(11):1749–1757, November 2020. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2019-0258.

Hermann:2020:ECG

[147] Nathan T. Hermann, Dominic T. Chaloner, Brandon S. Gerig, and Gary A. Lamberti. Ecological consequences of Great Lakes salmon subsidies for stream-resident brook and brown trout. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(11):1758–1771, November 2020. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2020-0086.

Alioravainen:2020:PED

[148] Nico Alioravainen, Jenni M. Prokkola, Alexandre Lemopoulos, Laura Härkönen, Pekka Hyvärinen, and Anssi Vainikka. Postrelease exploration and diel activity of hatchery, wild, and hybrid strain brown trout in seminatural streams. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(11): 1772–1779, November 2020. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2019-0436.

Reid:2020:ILW

[149] David A. Reid, Robin Pike, Stephen Bird, Peter Tschaplinski, and David Wilford. Implications of legacy watershed disturbances for channel structure and salmon habitat availability under different low-flow levels: an analysis of 45 years of discharge–habitat relationships at Carnation Creek, British Columbia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(11):1780– 1793, November 2020. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2020-0120.

Saul:2020:HFB

[150] Steven Saul, Elizabeth N. Brooks, and David Die. How fisher behavior can bias stock assessment: insights from an agent-based modeling approach. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(11):1794–1809, November 2020. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2019-0025.

Correa:2020:IEA

[151] Giancarlo M. Correa, Lorenzo Ciannelli, Lewis A. K. Barnett, Stan Kotwicki, and Claudio Fuentes. Improved estimation of age composition by accounting for spatiotemporal variability in somatic growth. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(11):1810–1821, November 2020. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2020-0166.

Cahill:2020:STA

[152] Christopher L. Cahill, Sean C. Anderson, Andrew J. Paul, Laura MacPherson, Michael G. Sullivan, Brett van Poorten, Carl J. Walters, and John R. Post. A spatial–temporal approach to modeling somatic growth across inland recreational fisheries landscapes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(11):1822–1835, November 2020. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2019-0434.

Perales:2020:LWL

[153] K. Martin Perales, Catherine L. Hein, Noah R. Lottig, and M. Jake Vander Zanden. Lake water level response to drought in a lake-rich region explained by lake and landscape characteristics. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(11):1836–1845, November 2020. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2019-0270.

Minkoff:2020:NAA

[154] David Minkoff, Nathan F. Putman, Jelle Atema, and William R. Ardren. Nonanadromous and anadromous Atlantic salmon differ in orientation responses to magnetic displacements. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(11):1846–1852, November 2020. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/cjfas-2020-0094.

Doretto:2020:RCC

[155] Alberto Doretto, Elena Piano, and Courtney E. Larson. The River Continuum Concept: lessons from the past and perspectives for the future. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(11):1853–1864, November 2020. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2020-0039.

Anonymous:2020:NA

[156] Anonymous. Note of appreciation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(12), December 2020. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/doi/

full/10.1139/cjfas-2020-0390.

Bass:2020:SEE

[157] Arthur L. Bass, Christine F. Stevenson, Aswea D. Porter, Erin L. Rechisky, Nathan B. Furey, Steve J. Healy, Adam M. Kanigan, Andrew G. Lotto, David W. Welch, and Scott G. Hinch. In situ experimental evaluation of tag burden and gill biopsy reveals survival impacts on migrating juvenile sockeye salmon. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(12): 1865–1869, December 2020. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2020-0134.

Eom:2020:LIS

[158] Junho Eom and Chris M. Wood. A less invasive system for the direct measurement of ventilation in fish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77 (12):1870–1877, December 2020. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/doi/

full/10.1139/cjfas-2020-0177.

vonBiela:2020:EPH

[159] Vanessa R. von Biela, Lizabeth Bowen, Stephen D. McCormick, Michael P. Carey, Daniel S. Donnelly, Shannon Waters, Amy M. Regish, Sarah M. Laske, Randy J. Brown, Sean Larson, Stanley Zuray, and Christian E.

Zimmerman. Evidence of prevalent heat stress in Yukon River Chinook salmon. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(12):1878–1892, December 2020. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/cjfas-
2020-0209.

DiSalvatore:2020:SKC

[160] Pablo Di Salvatore, Hernán J. Sacristán, M. Paula Sotelano, Federico Tapella, María Gowland-Sainz, and Gustavo A. Lovrich. Southern king crab larval survival: from intra- and interfemale variations to a fisheryinduced mortality. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(12):1893– 1903, December 2020. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2020-0040.

Peacock:2020:ECC

[161] Stephanie J. Peacock, Eric Hertz, Carrie A. Holt, Brendan Connors, Cameron Freshwater, and Katrina Connors. Evaluating the consequences of common assumptions in run reconstructions on Pacific salmon biological status assessments. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(12): 1904–1920, December 2020. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2019-0432.

Honsey:2020:DWR

[162] Andrew E. Honsey, Zachary S. Feiner, and Gretchen J. A. Hansen. Drivers of walleye recruitment in Minnesota’s large lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(12):1921–1933, December 2020. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2019-0453. See cor-

rection [404].

Kovach:2020:GMI

[163] Ryan P. Kovach, Andrew R. Whiteley, Matthew E. Jaeger, Sally Painter, Angela Lodmell, and Robb F. Leary. Genetic monitoring informs conservation status and trend of Arctic grayling at the southern edge of their distribution. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(12):1934–1942, December 2020. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/cjfas-
2020-0196.

Larson:2020:MLS

[164] Douglas L. Larson, Jacob G. Kimmel, Joseph J. Riedy, Jonathan Hegna, Edward A. Baker, and Kim T. Scribner. Male lake sturgeon (Acipenser fulvescens) migratory and spawning behaviors are associated with sperm quality and reproductive success. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(12):1943–1959, December 2020. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/cjfas-2020-0124.

Brookson:2020:DIJ

[165] Cole B. Brookson, Martin Krkošek, Brian P. V. Hunt, Brett T. Johnson, Luke A. Rogers, and Sean. C. Godwin. Differential infestation of juvenile Pacific salmon by parasitic sea lice in British Columbia, Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(12):1960–1968, December 2020. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2020-0160.

Rose:2020:ENY

[166] George A. Rose and Sherrylynn Rowe. Evidence of near year-around spawning in Atlantic cod off southern Newfoundland: implications for management. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(12):1969– 1976, December 2020. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2020-0206.

Connors:2020:CCC

[167] Brendan Connors, Michael J. Malick, Gregory T. Ruggerone, Pete Rand, Milo Adkison, James R. Irvine, Robert Campbell, and Kristen Gorman. Correction: Climate and competition influence sockeye salmon population dynamics across the Northeast Pacific Ocean. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 77(12):1977–1978, December 2020. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2020-0401. See [81].

Link:2021:NFS

[168] Jason S. Link, Francisco E. Werner, Kevin Werner, John F. Walter, III, Mark Strom, Michael P. Seki, Franklin Schwing, Jeremy Rusin, Clay E.

Porch, Kenric Osgood, Karl Moline, Richard D. Methot, Jr., Patrick D. Lynch, Douglas Lipton, Kristen Koch, Evan A. Howell, Jonathan A. Hare, Robert J. Foy, David Detlor, Lisa Desfosse, John Crofts, and Nicole Cabana. A NOAA fisheries science perspective on the conditions during and after COVID-19: challenges, observations, and some possible solutions, or why the future is upon us. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(1):1–12, January 2021. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/doi/

full/10.1139/cjfas-2020-0346.

Patterson:2021:ELV

[169] William F. Patterson, Beverly K. Barnett, Thomas C. TinHan, and Susan K. Lowerre-Barbieri. Eye lens δ ^{14}C validates otolith-derived age estimates of Gulf of Mexico reef fishes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(1):13–17, January 2021. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/doi/

full/10.1139/cjfas-2020-0237.

Church:2021:RCD

[170] Kathleen D. W. Church, Kevyn Janisse, Lida Nguyen-Dang, John W. Heath, Daniel D. Heath, and Christina A. D. Semeniuk. Rearing conditions differentially affect behavioural phenotypes of male “jack” and “hooknose” Chinook salmon and their sisters in both fresh and salt water. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(1):18–31, January 2021. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2020-0052.

McGuigan:2021:QEP

[171] Charles J. McGuigan, Lela S. Schlenker, John D. Stieglitz, Daniel D. Benetti, and Martin Grosell. Quantifying the effects of pop-up satellite archival tags on the swimming performance and behavior of young-adult mahi-mahi (Coryphaena hippurus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(1):32–39, January 2021. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/doi/

full/10.1139/cjfas-2020-0030.

Strailey:2021:SIR

[172] Katherine K. Strailey, Ryan T. Osborn, Rafael O. Tinoco, Piotr Cienciala, Bruce L. Rhoads, and Cory D. Suski. Simulated instream restoration structures offer smallmouth bass (Micropterus dolomieu) swimming and energetic advantages at high flow velocities. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(1):40–56, January 2021. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/cjfas-2020-0032.

Hollins:2021:DTP

[173] J. Hollins, B. Koeck, A. Crespel, D. M. Bailey, and S. S. Killen. Does thermal plasticity affect susceptibility to capture in fish? Insights from a simulated trap and trawl fishery. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(1):57–67, January 2021. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/doi/

full/10.1139/cjfas-2020-0125.

Austin:2021:WRN

[174] Catherine S. Austin, Timothy E. Essington, and Thomas P. Quinn. In a warming river, natural-origin Chinook salmon spawn later but hatcheryorigin conspecifics do not. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78 (1):68–77, January 2021. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2020-0060.

Sutherland:2021:PSE

[175] Ben J. G. Sutherland, John Candy, Kayla Mohns, Olivia Cornies, Kim Jonsen, Khai Le, Richard G. Gustafson, Krista M. Nichols, and Terry D. Beacham. Population structure of eulachon (Thaleichthys pacificus) from Northern California to Alaska using single nucleotide polymorphisms from direct amplicon sequencing. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78 (1):78–89, January 2021. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2020-0200.

Lennox:2021:VWA

[176] Robert J. Lennox, Ulrich Pulg, Brendan Malley, Sven-Erik Gabrielsen, Erlend M. Hanssen, Steven J. Cooke, Kim Birnie-Gauvin, Bjørn T. Barlaup, and Knut Wiik Vollset. The various ways that anadromous salmonids use lake habitats to complete their life history. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(1):90–100, January 2021. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/cjfas-2020-0225.

Zheng:2021:CSV

[177] Nan Zheng, Matthew Robertson, Noel Cadigan, Fan Zhang, Joanne Morgan, and Laura Wheeland. Correction: Spatiotemporal variation in maturation: a case study with American plaice (Hippoglossoides platessoides) on the Grand Bank off Newfoundland. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(1):101–102, January 2021. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/doi/

full/10.1139/cjfas-2020-0404. See [140].

Blair:2021:ACE

[178] Stephanie I. Blair, Clyde H. Barlow, and Jenifer K. McIntyre. Acute cerebrovascular effects in juvenile Coho salmon exposed to roadway runoff. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(2):103–109, February 2021. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2020-0240.

Cohen:2021:EVA

[179] Rachel S. Cohen, Derek K. Gray, Jasmina M. Vucic, Alyssa D. Murdoch, and Sapna Sharma. Environmental variables associated with littoral macroinvertebrate community composition in Arctic lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(2):110–123, February 2021. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2020-0065.

Hagelin:2021:UFP

[180] Anna Hagelin, Jon Museth, Larry Greenberg, Morten Kraabøl, Olle Calles, and Eva Bergman. Upstream fishway performance by Atlantic salmon (Salmo salar) and brown trout (Salmo trutta) spawners at complex hydropower dams — is prior experience a success criterion? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(2):124–134, February 2021. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2019-0271.

Mitchell:2021:ELT

[181] Kimberly T. Mitchell, Shawn R. Garner, Aimee Lee Houde, Chris C. Wilson, Trevor E. Pitcher, and Bryan D. Neff. Effects of a low-thiamine diet on reproductive traits in three populations of Atlantic salmon targeted for reintroduction into Lake Ontario. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(2):135–143, February 2021. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/doi/

full/10.1139/cjfas-2019-0379.

Itakura:2021:LCP

[182] Hikaru Itakura, Yoichi Miyake, Takashi Kitagawa, Takuya Sato, and Shingo Kimura. Large contribution of pulsed subsidies to a predatory fish inhabiting large stream channels. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(2):144–153, February 2021. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/doi/

full/10.1139/cjfas-2020-0004.

Clancy:2021:EDE

[183] Niall G. Clancy, Janice Brahney, James Dunnigan, and Phaedra Budy. Effects of a diatom ecosystem engineer (Didymosphenia geminata) on stream food webs: implications for native fishes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(2):154–164, February 2021. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/cjfas-2020-0121.

Massie:2021:FAA

[184] Danielle L. Massie, Yan Li, and Tyler Wagner. A framework for assessing the ability to detect macroscale effects on fish growth. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(2):165–172, February 2021. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2019-0296.

Hubbard:2021:NLT

[185] Justin A. G. Hubbard, Brendan E. Hickie, Jeff Bowman, Lee E. Hrenchuk, Paul J. Blanchfield, and Michael D. Rennie. No long-term effect of intracoelomic acoustic transmitter implantation on survival, growth, and body condition of a long-lived stenotherm in the wild. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(2):173–183, February 2021. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2020-0106.

Hatanpaa:2021:SSM

[186] Aurora Hatanpää, Hannu Huuskonen, Matti Janhunen, Raine Kortet, and Jorma Piironen. Spawning season movements of transported landlocked Atlantic salmon in a newly restored river habitat. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(2):184–192, February 2021. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2019-0334.

Irvine:2021:BHO

[187] James R. Irvine. Bias in hatchery-origin Coho salmon survival estimates due to underestimated prerelease mortality. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(3):193–196, March 2021. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/cjfas-2020-0287.

Barth:2021:ERC

[188] Lauren Emily Barth, Brian John Shuter, William Gary Sprules, Charles Kenneth Minns, and James Anthony Rusak. Evaluation of the responsiveness of the crustacean zooplankton community size spectrum to environmental change and an exotic invader in a sample of Canadian Shield lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(3):197–217, March 2021. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/cjfas-
2020-0038.

Pretty:2021:FIS

[189] Thomas J. Pretty, Charles-Matthew Chanyi, Catherine Kuhn, and Derek K. Gray. Factors influencing the structure of macroinvertebrate communities in subarctic lakes affected by wildfires. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(3):218–231, March 2021. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2020-0141.

Hunt:2021:IHP

[190] Len M. Hunt, Daniel J. Phaneuf, Joshua K. Abbott, Eli P. Fenichel, Jennifer A. Rodgers, Jeffrey D. Buckley, D. Andrew R. Drake, and Timothy B.

Johnson. The influence of human population change and aquatic invasive species establishment on future recreational fishing activities to the Canadian portion of the Laurentian Great Lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(3):232–244, March 2021. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/cjfas-2020-0159.

Liang:2021:REI

[191] Dong Liang, Geneviève M. Nesslage, Michael J. Wilberg, and Thomas J. Miller. Ranking ecosystem impacts on Chesapeake Bay blue crab (Callinectes sapidus) using empirical Gaussian Graphical Models. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(3):245–254, March 2021. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2019-0439.

Utne:2021:FIB

[192] Kjell Rong Utne, Katie Thomas, Jan Arge Jacobsen, Johanna Fall, Niall Ó. Maoiléidigh, Cecilie Thorsen Broms, and Webjørn Melle. Feeding interactions between Atlantic salmon (Salmo salar) postsmolts and other planktivorous fish in the Northeast Atlantic. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(3):255–268, March 2021. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/cjfas-2020-0037.

Wells:2021:SSG

[193] Mathew G. Wells, Jingzhi Li, Bryan Flood, Yulong Kuai, Jill L. Brooks, Steven J. Cooke, Patricia Semcesen, and Jonathan D. Midwood. Speed of sound gradients due to summer thermal stratification can reduce the detection range of acoustic fish tags: results from a field study in Hamilton Harbour, Ontario. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(3):269–285, March 2021. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/cjfas-
2020-0078.

Larson:2021:GCW

[194] Wesley A. Larson, Matthew S. Kornis, Keith N. Turnquist, Charles R. Bronte, Mark E. Holey, S. Dale Hanson, Theodore J. Treska, Wendylee Stott, and Brian L. Sloss. The genetic composition of wild recruits in a recovering lake trout population in Lake Michigan. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(3):286–300, March 2021. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2020-0072.

George:2021:CVP

[195] Scott D. George, Daniel S. Stich, and Barry P. Baldigo. Considerations of variability and power for long-term monitoring of stream fish assemblages. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(3):301–311, March 2021. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2020-0013.

Johannessen:2021:PPD

[196] Sophia C. Johannessen, Robie W. Macdonald, and Jonathan E. Strivens. Has primary production declined in the Salish Sea? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(3):312–321, March 2021. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2020-0115.

Rundio:2021:DGC

[197] David E. Rundio, John Carlos Garza, Steven T. Lindley, Thomas H. Williams, and Devon E. Pearse. Differences in growth and condition of juvenile Oncorhynchus mykiss related to sex and a migration-associated genomic region. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(3):322–331, March 2021. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/cjfas-
2020-0073.

Hagelin:2021:CAJ

[198] Anna Hagelin and Eva Bergman. Competition among juvenile brown trout, grayling, and landlocked Atlantic salmon in flumes — predicting effects of interspecific interactions on salmon reintroduction success. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(3):332–338, March 2021. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2020-0155.

Debertin:2021:CEB

[199] Allan J. Debertin. Correction: Estimating the biomass of a mixed-species complex using hydroacoustics and catch data from the Bay of Fundy and Scotian Shelf summer ecosystem survey. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(3):339, March 2021. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/doi/

full/10.1139/cjfas-2021-0021. See [94].

Heath:2021:BRS

[200] Victoria L. S. Heath, Scott Miehls, Nicholas Johnson, and Dennis M. Higgs. Behavioural response of sea lamprey (Petromyzon marinus) to acoustic stimuli in a small stream. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(4):341–348, April 2021. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/doi/

full/10.1139/cjfas-2020-0301.

Hargrove:2021:PBT

[201] John S. Hargrove, Carlos A. Camacho, William C. Schrader, John H. Powell, Thomas A. Delomas, Jon E. Hess, Shawn R. Narum, and Matthew R. Campbell. Parentage-based tagging improves escapement estimates for ESA-listed adult Chinook salmon and steelhead in the Snake River basin. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(4):349–360, April 2021. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2020-0152.

Fraser:2021:AVP

[202] Megan R. Fraser, Tyler Winsor, Jim Williams, Russell C. Wyeth, and David J. Garbary. Assessing the viability of pre-industrial sediment prior to remediation using primary producer (Zostera marina and Spartina alterniflora) growth and survival. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(4):361–370, April 2021. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/doi/

full/10.1139/cjfas-2019-0415.

Semmler:2021:IBC

[203] Robert Francis Semmler, Matthew Bryan Ogburn, Robert Aguilar, Elizabeth Watkins North, Marjorie Lindquist Reaka, and Anson Hemingway Hines. The influence of blue crab movement on mark–recapture estimates of recreational harvest and exploitation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(4):371–385, April 2021. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/cjfas-2020-0112.

Kneebone:2021:APS

[204] Jeff Kneebone, Hugues P. Benoît, Diego Bernal, and Walt Golet. Application of a parametric survival model to understand capture-related mortality and predation of yellowfin tuna (Thunnus albacares) released in a recreational fishery. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(4):386–399, April 2021. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/cjfas-
2020-0266.

Jansen:2021:ABT

[205] Teunis Jansen, Einar Eg Nielsen, Naiara Rodriguez-Ezpeleta, Haritz Arrizabalaga, Søren Post, and Brian R. MacKenzie. Atlantic bluefin tuna (Thunnus thynnus) in Greenland — mixed-stock origin, diet, hydrographic conditions, and repeated catches in this new fringe area. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(4):400–408, April 2021. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2020-0156.

Lewandoski:2021:BRS

[206] Sean A. Lewandoski, Peter Hrodey, Scott Miehls, Paul P. Piszczek, and Daniel P. Zielinski. Behavioral responses of sea lamprey (Petromyzon marinus) and white sucker (Catostomus commersonii) to turbulent flow during fishway passage attempts. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78 (4):409–421, April 2021. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2020-0223.

Sepulveda:2021:CLE

[207] Adam J. Sepulveda, Robert Al-Chokhachy, Matthew B. Laramie, Kyle Crapster, Ladd Knotek, Brian Miller, Alexander V. Zale, and David S. Pilliod. It’s complicated … environmental DNA as a predictor of trout and char abundance in streams. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78 (4):422–432, April 2021. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2020-0182.

King:2021:SWC

[208] Katelyn B. S. King, Mary Tate Bremigan, Dana Infante, and Kendra Spence Cheruvelil. Surface water connectivity affects lake and stream fish species richness and composition. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(4):433–443, April 2021. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/cjfas-2020-0090.

Heinle:2021:IWT

[209] Kadie B. Heinle, Lisa A. Eby, Clint C. Muhlfeld, Amber Steed, Leslie Jones, Vincent D’Angelo, Andrew R. Whiteley, and Mark Hebblewhite. Influence of water temperature and biotic interactions on the distribution of westslope cutthroat trout (Oncorhynchus clarkii lewisi) in a population stronghold under climate change. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(4):444–456, April 2021. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/doi/

full/10.1139/cjfas-2020-0099.

Li:2021:GDG

[210] Peiwen Li, Peter van Coeverden de Groot, Rute B. G. Clemente-Carvalho, and Stephen C. Lougheed. ddRAD genotyping reveals hierarchical genetic population structure in anadromous Arctic char (Salvelinus alpinus) in the Lower Northwest Passage, Nunavut. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(4):457–471, April 2021. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/doi/

full/10.1139/cjfas-2020-0069.

Geissinger:2021:FIS

[211] Emilie A. Geissinger, Robert S. Gregory, Benjamin J. Laurel, and Paul V. R. Snelgrove. Food and initial size influence overwinter survival and condition of a juvenile marine fish (age-0 Atlantic cod). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(4):472–482, April 2021. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2020-0142.

Eronen:2021:SMO

[212] A. Eronen, J. Kekäläinen, J. Piironen, P. Hyvärinen, H. Huuskonen, M. Janhunen, S. Yaripour, and R. Kortet. Sperm motility and offspring pre- and posthatching survival in hybridization crosses among a landlocked and two anadromous Atlantic salmon populations: implications for conservation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(4):483–492, April 2021. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/cjfas-
2020-0158.

Eisenhauer:2021:RRM

[213] Zachary J. Eisenhauer, Paul M. Christman, Jean-Michel Matte, William R. Ardren, Dylan J. Fraser, and James W. A. Grant. Revisiting the restricted movement paradigm: the dispersal of Atlantic salmon fry from artificial redds. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(4):493–503, April 2021. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/cjfas-
2020-0162.

Huntsman:2021:RDN

[214] Brock M. Huntsman, Frederick Feyrer, Matthew J. Young, James A. Hobbs, Shawn Acuña, Joseph E. Kirsch, Brian Mahardja, and Swee Teh. Recruitment dynamics of non-native largemouth bass within the Sacramento–San Joaquin Delta. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(5):505–521, May 2021. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/doi/

full/10.1139/cjfas-2020-0241.

Dillon:2021:BHA

[215] Rebecca A. Dillon, Joseph D. Conroy, Kathryn J. Lang, Kevin L. Pangle, and Stuart A. Ludsin. Bottom hypoxia alters the spatial distribution of pelagic intermediate consumers and their prey. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(5):522–538, May 2021. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/doi/

full/10.1139/cjfas-2020-0001.

Chalifour:2021:CSE

[216] Lia Chalifour, David C. Scott, Misty MacDuffee, Steven Stark, John F. Dower, Terry D. Beacham, Tara G. Martin, and Julia K. Baum. Chinook salmon exhibit long-term rearing and early marine growth in the Fraser River, British Columbia, a large urban estuary. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(5):539–550, May 2021. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/doi/

full/10.1139/cjfas-2020-0247.

Hussy:2021:EMD

[217] Karin Hüssy, Maria Krüger-Johnsen, Tonny Bernt Thomsen, Benjamin Dominguez Heredia, Tomas Næraa, Karin E. Limburg, Yvette Heimbrand, Kate McQueen, Stefanie Haase, Uwe Krumme, Michele Casini, Monica Mion, and Krzysztof Radtke. It’s elemental, my dear Watson: validating seasonal patterns in otolith chemical chronologies. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(5):551–566, May 2021. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2020-0388.

Couture:2021:ZCP

[218] Suzanne Couture, Christiane Hudon, Pierre Gagnon, Zofia E. Taranu, Bernadette Pinel-Alloul, Daniel Houle, Lama Aldamman, Conrad Beauvais, and Maude Lachapelle. Zooplankton communities in Precambrian Shield lakes (Quebec, Canada): responses to spatial and temporal gradients in water chemistry and climate. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(5):567–579, May 2021. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/doi/

full/10.1139/cjfas-2020-0139.

Feng:2021:IWT

[219] Wenping Feng, Nobuyasu Nakabayashi, Eri Inomata, Masakazu N. Aoki, and Yukio Agatsuma. Impacts of water temperature on the physiology and behaviours of the sea urchins Heliocidaris crassispina and Mesocentrotus nudus that reflect their range extension and disappearance in the Oga Peninsula, northern Honshu, Japan. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(5):580–588, May 2021. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/doi/

full/10.1139/cjfas-2020-0327.

Laurin:2021:ISS

[220] Emilie Laurin, Julia Bradshaw, Laura Hawley, Ian A. Gardner, Kyle Garver, Stewart C. Johnson, and Krishna K. Thakur. Importance of sample size for estimating prevalence: a case example of infectious hematopoietic necrosis viral RNA detection in mixed-stock Fraser River sockeye salmon (Oncorhynchus nerka), British Columbia, Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(5):589–598, May 2021. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2020-0279.

Xu:2021:CSS

[221] Luoliang Xu, Cameron Hodgdon, Ming Sun, Mackenzie Mazur, Xinjun Chen, and Yong Chen. Comparing a suite of surplus-production-based stock status identification approaches and management procedures. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(5):599–611, May 2021. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2020-0149.

Loeppky:2021:EIU

[222] Alison R. Loeppky and W. Gary Anderson. Environmental influences on uptake kinetics and partitioning of strontium in age-0 lake sturgeon (Acipenser fulvescens): effects of temperature and ambient calcium activities. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(5):612–622, May 2021. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2020-0335.

Giron-Nava:2021:EVF

[223] Alfredo Giron-Nava, Exequiel Ezcurra, Antoine Brias, Enriqueta Velarde, Ethan Deyle, Andrés M. Cisneros-Montemayor, Stephan B. Munch, George Sugihara, and Octavio Aburto-Oropeza. Environmental variability and fishing effects on the Pacific sardine fisheries in the Gulf of California. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(5):623–630, May 2021. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2020-0010.

Hasegawa:2021:SBB

[224] Koh Hasegawa, Kentaro Honda, Taku Yoshiyama, Kengo Suzuki, and Sho Fukui. Small biased body size of salmon fry preyed upon by piscivorous fish in riverine and marine habitats. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(5):631–638, May 2021. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/doi/

full/10.1139/cjfas-2020-0339.

Ardren:2021:ICE

[225] William R. Ardren and Larry Greenberg. Introduction to “Conservation, Ecology, and Evolution of Nonanadromous Atlantic Salmon”. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(6), June 2021. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2021-0035.

Hinch:2021:EHM

[226] Scott G. Hinch, Nolan N. Bett, Erika J. Eliason, Anthony P. Farrell, Steven J. Cooke, and David A. Patterson. Exceptionally high mortality of adult female salmon: a large-scale pattern and a conservation concern. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(6):639–654, June 2021. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2020-0385.

Friedland:2021:TPH

[227] Kevin D. Friedland. A test of the provisioning hypothesis of recruitment control in Georges Bank haddock. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(6):655–658, June 2021. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/doi/

full/10.1139/cjfas-2020-0453.

Trehin:2021:GDF

[228] Cécile Tréhin, Etienne Rivot, Ludivine Lamireau, Lisa Meslier, Anne-Laure Besnard, Stephen D. Gregory, and Marie Nevoux. Growth during the first summer at sea modulates sex-specific maturation schedule in Atlantic salmon. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(6):659–669, June 2021. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/cjfas-
2020-0236.

Loughin:2021:CSA

[229] Thomas M. Loughin, Stephen N. Bennett, and Nicolaas Bouwes. Comparison of staircase and asymmetrical before–after, control–impact (aBACI) experimental designs to test the effectiveness of stream restoration at increasing juvenile steelhead density. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(6):670–680, June 2021. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/doi/

full/10.1139/cjfas-2020-0096.

Tuset:2021:POS

[230] Víctor Manuel Tuset, José Luis Otero-Ferrer, Carolina Siliprandi, Amalia Manjabacas, Pere Marti-Puig, and Antoni Lombarte. Paradox of otolith shape indices: routine but overestimated use. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(6):681–692, June 2021. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/doi/

full/10.1139/cjfas-2020-0369.

Goode:2021:EBI

[231] Andrew G. Goode, Jonathan H. Grabowski, and Damian C. Brady. Evaluating benthic impact of the Gulf of Maine lobster fishery using the swept area seabed impact (SASI) model. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(6):693–703, June 2021. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/doi/

full/10.1139/cjfas-2020-0305.

McKee:2021:MAL

[232] Anne McKee, Jon Grant, and Jeffrey Barrell. Mapping American lobster (Homarus americanus) habitat for use in marine spatial planning. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(6):704–720, June 2021. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2020-0051.

Gislason:2021:SCL

[233] Davíd Gíslason, Robert B. McLaughlin, and Beren W. Robinson. Synchronous changes in length at maturity across four species of Lake Erie fish with different harvest histories. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(6):721–737, June 2021. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/doi/

full/10.1139/cjfas-2019-0408.

Biro:2021:ALR

[234] Peter A. Biro, John R. Post, and Christa Beckmann. Autumn lipid reserves, overwinter lipid depletion, and high winter mortality of rainbow trout in experimental lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78 (6):738–743, June 2021. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2020-0276.

Mari:2021:ETD

[235] Lisandrina Mari, Martin Daufresne, Jean Guillard, Guillaume Evanno, and Emilien Lasne. Elevated temperature and deposited sediment jointly affect early life history traits in southernmost Arctic char populations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(6):744–751, June 2021. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2020-0256.

Hoffman:2021:CLF

[236] Joel Christopher Hoffman, Christy Meredith, Erik Pilgrim, Anett Trebitz, Chelsea Hatzenbuhler, John Russell Kelly, Gregory Peterson, Julie Lietz, Sara Okum, and John Martinson. Comparison of larval fish detections using morphology-based taxonomy versus high-throughput sequencing for invasive species early detection. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(6):752–764, June 2021. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/doi/

full/10.1139/cjfas-2020-0224.

Larocque:2021:TNO

[237] Sarah M. Larocque, Timothy B. Johnson, and Aaron T. Fisk. Trophic niche overlap and abundance reveal potential impact of interspecific interactions on a reintroduced fish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(6):765–774, June 2021. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/doi/

full/10.1139/cjfas-2020-0204.

Regish:2021:SWM

[238] Amy M. Regish, William R. Ardren, Nicholas R. Staats, Henry Bouchard, Jonah L. Withers, Theodore Castro-Santos, and Stephen D. McCormick. Surface water with more natural temperatures promotes physiological and endocrine changes in landlocked Atlantic salmon smolts. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(6):775–786, June 2021. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2020-0295.

Greenberg:2021:LAS

[239] Larry A. Greenberg, Johnny R. Norrgård, Pär Gustafsson, and Eva Bergman. Landlocked Atlantic salmon in a large river–lake ecosystem: managing an endemic, large-bodied population of high conservation value. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(6):787–796, June 2021. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2020-0163.

Vinebrooke:2021:TDS

[240] Rolf Vinebrooke. A tribute to David Schindler: the sentinel of fresh waters. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(7), July 2021. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2021-0106.

Gatto:2021:LDT

[241] John V. Gatto, Jeffrey L. Kline, William F. Loftus, and Joel C. Trexler. Linking demographic transitions to population dynamics in a fluctuating environment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(7):797–808, July 2021. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/cjfas-
2020-0101.

Dodrill:2021:PTR

[242] Michael J. Dodrill, Charles B. Yackulic, Theodore A. Kennedy, Michael D. Yard, and Josh Korman. As the prey thickens: rainbow trout select prey based upon width not length. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(7):809–819, July 2021. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/doi/

full/10.1139/cjfas-2020-0113.

Gillis:2021:POW

[243] Daniel P. Gillis, Charles K. Minns, and Brian J. Shuter. Predicting open-water thermal regimes of temperate North American lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(7):820–840, July 2021. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2020-0140.

DeWeber:2021:LTC

[244] J. Tyrell DeWeber, Roland Rösch, Jan Baer, and Alexander Brinker. Long-term changes in body condition and gillnet selectivity in lake constance pelagic spawning whitefish (Coregonus wartmanni). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(7):841–851, July 2021. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2020-0231.

Devlin:2021:EGI

[245] Robert H. Devlin, Janine Supernault, Salvador A. Gezan, Michelle T. T. Chan, Mike Wetklo, Carlo A. Biagi, Dionne Sakhrani, Stu Barnetson, Glen Dixon, Evelyn Tattersall, William S. Davidson, Ben F. Koop, David M. Willis, and Ruth E. Withler. Environmental and genetic influences on fitness-related traits in a hatchery Coho salmon population. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(7):852–868, July 2021. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2020-0383.

Alcott:2021:TGF

[246] Derrick Alcott, Elsa Goerig, Christopher Rillahan, Pingguo He, and Theodore Castro-Santos. Tide gates form physical and ecological obstacles to river herring (Alosa spp.) spawning migrations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(7):869–880, July 2021. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2020-0347.

Sewall:2021:GCS

[247] Fletcher Sewall, Brenda Norcross, and Ron Heintz. Growth, condition, and swimming performance of juvenile Pacific herring with winter feeding rations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(7):881–893, July 2021. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2020-0293.

Chen:2021:RRC

[248] Emily K. Chen and Mark J. Henderson. Reduced recruitment of Chinook salmon in a leveed bar-built estuary. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(7):894–904, July 2021. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/doi/

full/10.1139/cjfas-2020-0122.

Engman:2021:BRL

[249] Augustin C. Engman, Thomas J. Kwak, and Jesse R. Fischer. Big runs of little fish: first estimates of run size and exploitation in an amphidromous postlarvae fishery. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(7):905–912, July 2021. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/cjfas-
2020-0093.

Massie:2021:DLS

[250] Danielle L. Massie, Gretchen J. A. Hansen, Yan Li, Greg G. Sass, and Tyler Wagner. Do lake-specific characteristics mediate the temporal relationship between walleye growth and warming water temperatures? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(7):913–923, July 2021. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2020-0169.

Gillis:2021:AND

[251] Darren M. Gillis, Adriaan D. Rijnsdorp, and Jan Jaap Poos. Association networks in the Dutch offshore beam trawl fleet: their predictors and relationship to vessel performance. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(7):924–942, July 2021. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/doi/

full/10.1139/cjfas-2019-0353.

Liss:2021:MNA

[252] Stephanie A. Liss, Katherine R. Znotinas, Shannon E. Blackburn, Eric S. Fischer, James S. Hughes, Ryan A. Harnish, Huidong Li, and Zhiqun Daniel Deng. From 95 to 59 millimetres: a new active acoustic tag size guideline for salmon. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(7):943–957, July 2021. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/doi/

full/10.1139/cjfas-2020-0222.

Jorde:2021:MGC

[253] Per Erik Jorde, Mats B. O. Huserbråten, Bjørghild B. Seliussen, Mari S. Myksvoll, Frode B. Vikebø, Geir Dahle, Asgeir Aglen, and Torild Johansen. The making of a genetic cline: introgression of oceanic genes into coastal cod populations in the Northeast Atlantic. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(7):958–968, July 2021. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2020-0380.

Leech:2021:IVD

[254] Dina M. Leech, Troy L. Clift, Jessica L. Littlefield, Nicholas R. Ravagli, and Jacob E. Spain. Indirect versus direct effects of freshwater browning on larval fish foraging. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78 (7):969–983, July 2021. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2020-0379.

Benoit:2021:PFC

[255] David M. Benoit, Donald A. Jackson, and Cindy Chu. Partitioning fish communities into guilds for ecological analyses: an overview of current approaches and future directions. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(7):984–993, July 2021. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/doi/

full/10.1139/cjfas-2020-0455.

Chapman:2021:DEW

[256] Jacqueline M. Chapman, Lisa A. Kelly, Amy K. Teffer, Kristi M. Miller, and Steven J. Cooke. Disease ecology of wild fish: opportunities and challenges for linking infection metrics with behaviour, condition, and survival. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(8):995–1007, August 2021. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2020-0315.

Smith:2021:DRE

[257] William E. Smith, Leo Polansky, and Matthew L. Nobriga. Disentangling risks to an endangered fish: using a state-space life cycle model to separate natural mortality from anthropogenic losses. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(8):1008–1029, August 2021. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/cjfas-2020-0251.

Mochnacz:2021:FED

[258] Neil J. Mochnacz, Darryl I. MacKenzie, Nicola Koper, Margaret F. Docker, and Dan J. Isaak. Fringe effects: detecting bull trout (Salvelinus confluentus) at distributional boundaries in a montane watershed. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(8):1030–1044, August 2021. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2020-0219.

Rossi:2021:FMM

[259] Gabriel J. Rossi, Mary E. Power, Shelley Pneh, Jason R. Neuswanger, and Timothy J. Caldwell. Foraging modes and movements of Oncorhynchus mykiss as flow and invertebrate drift recede in a California stream. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(8):1045–1056, August 2021. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2020-0398.

Dzul:2021:PMS

[260] Maria Christina Dzul, William Louis Kendall, Charles Brandon Yackulic, Dana L. Winkelman, David Randall Van Haverbeke, and Michael Yard. Partial migration and spawning movements of humpback chub in the Little Colorado River are better understood using data from autonomous PIT tag antennas. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(8):1057–1072, August 2021. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/cjfas-
2020-0291.

Al-Chokhachy:2021:IES

[261] Robert Al-Chokhachy, Mike Lien, Bradley B. Shepard, and Brett High. The interactive effects of stream temperature, stream size, and non-native species on Yellowstone cutthroat trout. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(8):1073–1083, August 2021. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/doi/

full/10.1139/cjfas-2020-0408.

Bradford:2021:RLE

[262] Michael J. Bradford and Douglas C. Braun. Regional and local effects drive changes in spawning stream occupancy in a sockeye salmon metapopulation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(8):1084–1095, August 2021. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/cjfas-
2020-0463.

Cantin:2021:PDP

[263] Ariane Cantin, Anne Farineau, Darren J. Bender, and John R. Post. Predicting distribution, production, recreational fishing demand, and conservation concern of indigenous rainbow trout populations based on landscape characteristics. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(8): 1096–1108, August 2021. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2020-0323.

Day:2021:EMF

[264] Casey C. Day, Erin L. Landguth, Ryan K. Simmons, William P. Baker, Andrew R. Whiteley, Paul M. Lukacs, Kaeli A. Davenport, and Andrew R. Bearlin. Evaluation of management factors affecting the relative success of a brook trout eradication program using YY male fish and electrofishing suppression. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(8):1109–1119, August 2021. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/cjfas-
2020-0433.

Boyce:2021:LLH

[265] Daniel G. Boyce, Susanna Fuller, Chelsey Karbowski, Katie Schleit, and Boris Worm. Leading or lagging: How well are climate change considerations being incorporated into Canadian fisheries management? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(8):1120–1129, August 2021. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2020-0394.

Krigbaum:2021:IVT

[266] Melissa J. Krigbaum and Christopher M. Anderson. Increasing value through gear flexibility: a case study of US west coast sablefish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(8):1130–1145, August 2021. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2020-0399.

Janak:2021:PME

[267] Jill M. Janak, Timothy J. Linley, Ryan A. Harnish, and Steve D. Shen. Partitioning maternal and exogenous diet contributions to otolith ^{87}Sr/ 86Sr in Kokanee salmon (Oncorhynchus nerka). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(8):1146–1157, August 2021. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/cjfas-2020-0242.

Ivanova:2021:SIN

[268] Silviya V. Ivanova, Sarah M. Larocque, Aaron T. Fisk, and Timothy B. Johnson. Spatiotemporal interactions of native and introduced salmonid top predators in a large lake: implications for species restoration. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(8):1158–1167, August 2021. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2020-0447.

Sagarese:2021:ESS

[269] Skyler R. Sagarese, Nathan R. Vaughan, John F. Walter, and Mandy Karnauskas. Enhancing single-species stock assessments with diverse ecosystem perspectives: a case study for Gulf of Mexico red grouper (Epinephelus morio) and red tides. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(8): 1168–1180, August 2021. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2020-0257.

Mensinger:2021:CDP

[270] Matthew A. Mensinger, Erik J. Blomberg, and Joseph D. Zydlewski. The consequences of dam passage for downstream-migrating American eel in the Penobscot River, Maine. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78 (8):1181–1192, August 2021. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2020-0402.

Aalto:2021:SOC

[271] Emilius A. Aalto, Francesco Ferretti, Matthew V. Lauretta, John F. Walter, Michael J. W. Stokesbury, Robert J. Schallert, and Barbara A. Block. Stock-of-origin catch estimation of Atlantic bluefin tuna (Thunnus thynnus) based on observed spatial distributions. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(8):1193–1204, August 2021. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/cjfas-2019-0445.

Anonymous:2021:SSC

[272] Anonymous. Science to support Canada’s SARA-listed freshwater species. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(9), September 2021. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2021-0234.

Drake:2021:FYC

[273] D. Andrew R. Drake, Karl A. Lamothe, Kristin E. Thiessen, Todd J. Morris, Marten A. Koops, Thomas C. Pratt, Scott M. Reid, Donald A. Jackson, and Nicholas E. Mandrak. Fifteen years of Canada’s Species at Risk Act: Evaluating research progress for aquatic species in the Great Lakes– St. Lawrence River basin. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(9): 1205–1218, September 2021. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2021-0143.

Rosenfeld:2021:ELS

[274] Jordan Rosenfeld, Michael P. Pearson, Jill Miners, and Kaitlyn Zinn. Effects of landscape-scale hypoxia on salish sucker and salmonid habitat associations: implications for endangered species recovery and management. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(9):1219–1233, September 2021. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2020-0171.

Zinn:2021:EEF

[275] Kaitlyn R. Zinn, Jordan S. Rosenfeld, and Eric B. Taylor. Effects of experimental flow manipulations on water quality, hypoxia, and growth of threatened salish sucker (Catostomus sp. cf. catostomus) and juvenile Coho salmon (Oncorhynchus kisutch). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(9):1234–1246, September 2021. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/cjfas-2020-0135.

Macnaughton:2021:MPT

[276] Camille J. Macnaughton, Travis C. Durhack, Neil J. Mochnacz, and Eva C. Enders. Metabolic performance and thermal preference of westslope cutthroat trout (Oncorhynchus clarkii lewisi) and non-native trout across an ecologically relevant range of temperatures. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(9):1247–1256, September 2021. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2020-0173.

McDonnell:2021:EAE

[277] Laura H. McDonnell, Nicholas E. Mandrak, Saravpreet Kaur, and Lauren J. Chapman. Effects of acclimation to elevated water temperature and hypoxia on thermal tolerance of the threatened pugnose shiner (Notropis anogenus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(9): 1257–1267, September 2021. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2020-0362.

Rosencranz:2021:DLM

[278] Jordan Rosencranz, Kim Cuddington, Madison Brook, Marten A. Koops, and D. Andrew Drake. Data-limited models to predict river temperatures for aquatic species at risk. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(9): 1268–1277, September 2021. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2020-0294.

Rodriguez:2021:RVF

[279] Marco A. Rodríguez, Geoffrey Marselli, and Nicholas E. Mandrak. Responses of vulnerable fishes to environmental stressors in the Canadian Great Lakes basin. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(9):1278– 1292, September 2021. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2020-0314.

Potts:2021:FSD

[280] Lindsay B. Potts, Nicholas E. Mandrak, and Lauren J. Chapman. Finescale distribution and occupancy modelling of the threatened pugnose shiner (Notropis anogenus) in the St. Lawrence River, Ontario, Canada.

Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(9):1293–1304, September 2021. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2020-0456.

Lamothe:2021:ODW

[281] Karl A. Lamothe, Kelly A. McNichols-O’Rourke, and Todd J. Morris. Occupancy and detection of wavyrayed lampmussel (Lampsilis fasciola) in Ontario, Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(9): 1305–1311, September 2021. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2020-0325.

VanTassel:2021:GDM

[282] Nichelle M. VanTassel, Todd J. Morris, Christopher G. Wilson, and David T. Zanatta. Genetic diversity maintained in comparison of captivepropagated and wild populations of Lampsilis fasciola and Ptychobranchus fasciolaris (Bivalvia: Unionidae). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(9):1312–1320, September 2021. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/cjfas-2020-0373.

Lamothe:2021:TTE

[283] Karl A. Lamothe, Adam S. van der Lee, D. Andrew R. Drake, and Marten A. Koops. The translocation trade-off for eastern sand darter (Ammocrypta pellucida): balancing harm to source populations with the goal of re-establishment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(9): 1321–1331, September 2021. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2020-0288.

Rytwinski:2021:WEE

[284] Trina Rytwinski, Lisa A. Kelly, Lisa A. Donaldson, Jessica J. Taylor, Adrienne Smith, D. Andrew R. Drake, André L. Martel, Juergen Geist, Todd J. Morris, Anna L. George, Alan J. Dextrase, Joseph R. Bennett, and Steven J. Cooke. What evidence exists for evaluating the effectiveness of conservation-oriented captive breeding and release programs for imperilled freshwater fishes and mussels? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(9): 1332–1346, September 2021. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2020-0331.

Turko:2021:CSP

[285] Andy J. Turko, Alexandra T. A. Leclair, Nicholas E. Mandrak, D. Andrew R. Drake, Graham R. Scott, and Trevor E. Pitcher. Choosing source populations for conservation reintroductions: lessons from variation in thermal tolerance among populations of the imperilled redside dace. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(9):1347–1355, September 2021. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2020-0377.

Castaneda:2021:ARN

[286] Rowshyra A. Castañeda, Josef D. Ackerman, Lauren J. Chapman, Steven J. Cooke, Kim Cuddington, Alan J. Dextrase, Donald A. Jackson, Marten A. Koops, Martin Krkošek, Kevin K. Loftus, Nicholas E. Mandrak, André L. Martel, Péter K. Molnár, Todd J. Morris, Trevor E. Pitcher, Mark S. Poesch, Michael Power, Thomas C. Pratt, Scott M. Reid, Marco A. Rodríguez, Jordan Rosenfeld, Chris C. Wilson, David T. Zanatta, and D. Andrew R. Drake. Approaches and research needs for advancing the protection and recovery of imperilled freshwater fishes and mussels in Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(9):1356– 1370, September 2021. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2020-0374.

Conley:2021:MHS

[287] Amy K. Conley, Matthew D. Schlesinger, James G. Daley, Lisa K. Holst, and Timothy G. Howard. Modeling habitat suitability and management options for maintaining round whitefish (Prosopium cylindraceum) in Adirondack ponds. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(10): 1371–1382, October 2021. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2020-0303.

Hocking:2021:SSR

[288] M. D. Hocking, S. G. Faulkner, K. Akaoka, A. Harwood, T. Hatfield, and F. J. A. Lewis. Surprising salmonid response to water diversion at four run-of-river hydroelectric projects in British Columbia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(10):1383–1396, October 2021. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2020-0246. See com-

ment [630] and reply [631].

Petersen:2021:MEB

[289] Kelly N. Petersen, Mary C. Freeman, Joseph E. Kirsch, William O. McLarney, Mark C. Scott, and Seth J. Wenger. Mixed evidence for biotic homogenization of Southern Appalachian fish communities. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(10):1397–1406, October 2021. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2020-0372.

Munaweera:2021:AMP

[290] I. Munaweera, S. Muthukumarana, D. M. Gillis, D. A. Watkinson, C. Charles, and E. C. Enders. Assessing movement patterns using Bayesian state space models on Lake Winnipeg walleye. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(10):1407–1421, October 2021. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2020-0262.

Beacham:2021:ECU

[291] Terry D. Beacham, Colin Wallace, Kim Jonsen, Ben J. G. Sutherland, Carrie Gummer, and Eric B. Rondeau. Estimation of conservation unit and population contribution to Chinook salmon mixed-stock fisheries in British Columbia, Canada, using direct DNA sequencing for single nucleotide polymorphisms. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(10): 1422–1434, October 2021. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2020-0462.

Birnie-Gauvin:2021:ESC

[292] Kim Birnie-Gauvin, Martin H. Larsen, and Kim Aarestrup. Energetic state and the continuum of migratory tactics in brown trout (Salmo trutta). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(10):1435–1443, October 2021. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/cjfas-
2021-0025.

Peterson:2021:EME

[293] Lisa K. Peterson, Michael L. Jones, Travis O. Brenden, Christopher S. Vandergoot, and Charles C. Krueger. Evaluating methods for estimating mortality from acoustic telemetry data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(10):1444–1454, October 2021. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2020-0417.

Simmatis:2021:LCS

[294] Brigitte Simmatis, Andrew M. Paterson, Brie A. Edwards, and John P. Smol. Linking changes in sedimentary Cladocera assemblages to limnological variables in 67 Sudbury (Ontario, Canada) lakes reflecting various degrees of metal smelter impacts. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(10):1455–1469, October 2021. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/cjfas-2020-0351.

Tsuboi:2021:SSV

[295] Jun ichi Tsuboi, Kentaro Morita, Genki Sahashi, Mari Kuroki, Shinya Baba, and Robert Arlinghaus. Species-specific vulnerability to angling and its size-selectivity in sympatric stream salmonids. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(10):1470–1478, October 2021. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2020-0428.

Harkonen:2021:WGA

[296] Laura Härkönen, Pauliina Louhi, Riina Huusko, and Ari Huusko. Wintertime growth in Atlantic salmon under changing climate: the importance of ice cover for individual growth dynamics. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(10):1479–1485, October 2021. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2020-0258.

Lackmann:2021:LLP

[297] Alec R. Lackmann, Brandon J. Kratz, Ewelina S. Bielak-Lackmann, Reed I. Jacobson, Derek J. Sauer, Allen H. Andrews, Malcolm G. Butler, and Mark E. Clark. Long-lived population demographics in a declining, vulnerable fishery — bigmouth buffalo (Ictiobus cyprinellus) of Jamestown Reservoir, North Dakota. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(10):1486–1496, October 2021. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/cjfas-2020-0485.

OToole:2021:PGR

[298] Ciar O’Toole, Karl P. Phillips, Caroline Bradley, Jamie Coughlan, Eileen Dillane, Ian A. Fleming, Thomas E. Reed, Peter A. H. Westley, Tom F. Cross, Philip McGinnity, and Paulo A. Prodöhl. Population genetics reveal patterns of natural colonisation of an ecologically and commercially important invasive fish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(10): 1497–1511, October 2021. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2020-0255.

Albertsen:2021:EMP

[299] Christoffer Moesgaard Albertsen, Karin Hüssy, Simon Hansen Serre, Jakob Hemmer-Hansen, and Tonny Bernt Thomsen. Estimating migration patterns of fish from otolith chemical composition time series. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(10):1512–1523, October 2021. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2020-0356.

Bunnell:2021:CCW

[300] David B. Bunnell, Stuart A. Ludsin, Roger L. Knight, Lars G. Rudstam, Craig E. Williamson, Tomas O. Höök, Paris D. Collingsworth, Barry M. Lesht, Richard P. Barbiero, Anne E. Scofield, Edward S. Rutherford, Layne Gaynor, Henry A. Vanderploeg, and Marten A. Koops. Consequences of changing water clarity on the fish and fisheries of the Laurentian Great Lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(10): 1524–1542, October 2021. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2020-0376.

Pinel-Alloul:2021:MSM

[301] Bernadette Pinel-Alloul, Alain Patoine, and Jérôme Marty. Multi-scale and multi-system perspectives of zooplankton structure and function in Canadian freshwaters. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(10):

1543–1562, October 2021. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2020-0474.

Pleizier:2021:CUG

[302] Naomi K. Pleizier, Charlotte Nelson, Steven J. Cooke, and Colin J. Brauner. Correction: Understanding gas bubble trauma in an era of hydropower expansion: how do fish compensate at depth? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(10):1563, October 2021. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2021-0207. See [48].

Agha:2021:TPE

[303] Mickey Agha, James P. Losee, Marisa N. C. Litz, Craig Smith, Jason J. Schaffler, William S. Patton, Aaron M. Dufault, and Gabe M. Madel. Temporal patterns and ecosystem correlates of chum salmon (Oncorhynchus keta) migration phenology in the Pacific Northwest. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(11):1565–1575, November 2021. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2020-0392.

Gundelund:2021:ECS

[304] Casper Gundelund, Paul Venturelli, Bruce W. Hartill, Kieran Hyder, Hans Jakob Olesen, and Christian Skov. Evaluation of a citizen science platform for collecting fisheries data from coastal sea trout anglers. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(11):1576–1585, November 2021. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2020-0364.

Vainikka:2021:FIV

[305] Anssi Vainikka, Pekka Hyvärinen, Joni Tiainen, Alexandre Lemopoulos, Nico Alioravainen, Jenni M. Prokkola, Chris K. Elvidge, and Robert Arlinghaus. Fishing-induced versus natural selection in different brown trout (Salmo trutta) strains. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(11): 1586–1596, November 2021. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2020-0313.

Sylvia:2021:TNT

[306] Andrea Sylvia, Stephen J. Dinsmore, and Michael J. Weber. Tournament and non-tournament anglers have little effect on a largemouth bass population compared to natural mortality. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(11):1597–1610, November 2021. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/cjfas-2020-0442.

McGowan:2021:MDS

[307] David W. McGowan, Trevor A. Branch, Stormy Haught, and Mark D. Scheuerell. Multi-decadal shifts in the distribution and timing of Pacific herring (Clupea pallasii) spawning in Prince William Sound, Alaska. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(11):1611–1627, November 2021. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2021-0047.

Morissette:2021:QMC

[308] Olivier Morissette, Frédéric Lecomte, Nathalie Vachon, Annick Drouin, and Pascal Sirois. Quantifying migratory capacity and dispersal of the invasive tench (Tinca tinca) in the St. Lawrence River using otolith chemistry. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(11):1628–1638, November 2021. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2020-0460.

Gobin:2021:MRN

[309] Jenilee Gobin, Michael G. Fox, and Erin S. Dunlop. Maturation reaction norm evolution under varying conditions of eco-evolutionary change. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(11):1639–1649, November 2021. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2020-0274.

Needham:2021:RBT

[310] Robert J. Needham, Martin Gaywood, Angus Tree, Nick Sotherton, Dylan Roberts, Colin W. Bean, and Paul S. Kemp. The response of a brown trout (Salmo trutta) population to reintroduced Eurasian beaver (Castor fiber) habitat modification. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78 (11):1650–1660, November 2021. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/doi/

full/10.1139/cjfas-2021-0023.

Nelson:2021:VPD

[311] Benjamin W. Nelson, Scott F. Pearson, Joseph H. Anderson, Steven J. Jeffries, Austen C. Thomas, William A. Walker, Alejandro Acevedo-Gutiérrez, Iris M. Kemp, Monique M. Lance, Amelia Louden, and Madelyn R. Voelker. Variation in predator diet and prey size affects perceived impacts to salmon species of high conservation concern. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(11):1661–1676, November 2021. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2020-0300.

Wilson:2021:RDN

[312] Sean M. Wilson, Matthew P. Corsi, Darren H. Brandt, and Eric J. Stark. The response of Daphnia to nutrient additions and Kokanee abundance in Dworshak Reservoir, Idaho. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78 (11):1677–1688, November 2021. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/doi/

full/10.1139/cjfas-2020-0427.

Friedland:2021:TPC

[313] Kevin D. Friedland, John R. Moisan, Aurore A. Maureaud, Damian C. Brady, Andrew J. Davies, Steven J. Bograd, Reg A. Watson, and Yannick Rousseau. Trends in phytoplankton communities within large marine ecosystems diverge from the global ocean. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(11):1689–1700, November 2021. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2020-0423.

Humston:2021:NID

[314] Robert Humston, Eric Hallerman, Scott Smith, Juli Sorenson, and Garrett Muckleroy. Natal and intergenerational dispersal of riverine smallmouth bass (Micropterus dolomieu). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78 (11):1701–1711, November 2021. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/doi/

full/10.1139/cjfas-2020-0322.

Frommel:2021:EOM

[315] Andrea Y. Frommel, Vahab Pourfaraj, Colin J. Brauner, and D. Steven O. McAdam. Evaluation of otolith microstructures for ageing of larval white sturgeon. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(11):1712– 1720, November 2021. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2020-0437.

Parzanini:2021:TEE

[316] Camilla Parzanini, Michael T. Arts, Michael Power, Mehis Rohtla, Anne Berit Skiftesvik, Janet Koprivnikar, Howard I. Browman, Dino Milotic, and Caroline M. F. Durif. Trophic ecology of the European eel (Anguilla anguilla) across different salinity habitats inferred from fatty acid and stable isotope analysis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(11):1721–1731, November 2021. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/cjfas-2020-0432.

Sato:2021:SRK

[317] Mei Sato, Andrew W. Trites, and Stéphane Gauthier. Southern resident killer whales encounter higher prey densities than northern resident killer whales during summer. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(11): 1732–1743, November 2021. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2020-0445.

Dey:2021:RPM

[318] Cody J. Dey, Adam I. Rego, Michael J. Bradford, Keith D. Clarke, Katherine McKercher, Neil J. Mochnacz, Alex de Paiva, Karin Ponader, Lisa Robichaud, Amanda K. Winegardner, Court Berryman, Paul J. Blanchfield, Christine M. Boston, Doug Braun, Jacob W. Brownscombe, Christopher Burbidge, Stuart Campbell, Alicia Cassidy, Cindy Chu, Steven J. Cooke, Daniel Coombs, Jenie Cooper, Allen Curry, Maja Cvetkovic, Andréanne Demers, Margaret Docker, Andrea Doherty, Susan E. Doka, Karen Dunmall, Brie Edwards, Eva C. Enders, Neil Fisher, Marika Gauthier-Ouellet, William Glass, Les N. Harris, Caleb Hasler, Jaclyn Hill, Scott G. Hinch, Emma E. Hodgson, Jason Hwang, Ken M. Jeffries, Lonnie King, Rick Kiriluk, Rob Knight, Alex Levy, Jennifer MacDonald, Rob Mackereth, Rob McLaughlin, Charles K. Minns, Jonathan W. Moore, Karine Nantel, Chantal Nessman, Claude Normand, Constance M. O’Connor, Joclyn Paulic, Laura Phalen, John Post, Thomas C. Pratt, Scott M. Reid, C. Alwyn Rose, Jordan Rosenfeld, Karen E. Smokorowski, Darrin Sooley, Mark K. Taylor, Jason Treberg, Jacques Trottier, Tyler D. Tunney, Marie-Pierre Veilleux, Doug A. Watkinson, Dean Watts, Karen Winfield, Jacob P. Ziegler, Jonathan D. Midwood, and Marten A. Koops. Research priorities for the management of freshwater fish habitat in Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(11):1744–1754, November 2021. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2021-0002.

Anonymous:2021:NA

[319] Anonymous. Note of appreciation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(12), December 2021. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/doi/

full/10.1139/cjfas-2021-0251.

Staton:2021:IDI

[320] Benjamin A. Staton, Matthew J. Catalano, Steven J. Fleischman, and Jan Ohlberger. Incorporating demographic information into spawner– recruit analyses alters biological reference point estimates for a western Alaska salmon population. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78 (12):1755–1769, December 2021. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/doi/

full/10.1139/cjfas-2020-0478.

Walsworth:2021:EBS

[321] Timothy E. Walsworth and Phaedra Budy. An empirically based simulation model to inform flow management for endangered species conservation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(12):1770–1781, December 2021. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/cjfas-
2020-0353.

Farly:2021:HCF

[322] Luc Farly, Christiane Hudon, Antonia Cattaneo, and Gilbert Cabana. Hydrological control of a floodplain subsidy to littoral riverine fish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(12):1782–1792, December 2021. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2020-0395.

Saglam:2021:GWA

[323] İsmail K. Sağlam, James Hobbs, Randall Baxter, Levi S. Lewis, Alyssa Benjamin, and Amanda J. Finger. Genome-wide analysis reveals regional patterns of drift, structure, and gene flow in longfin smelt (Spirinchus thaleichthys) in the northeastern Pacific. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(12):1793–1804, December 2021. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2021-0005.

Ojaveer:2021:SSB

[324] Henn Ojaveer, Riina Klais-Peets, Heli Einberg, and Gunta Rubene. Spawning stock biomass modulation of environment–recruitment relationship in a marginal spring spawning herring (Clupea harengus membras) population. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(12):1805– 1815, December 2021. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2021-0018.

Hui:2021:IDM

[325] Yuan Hui, Joseph F. Atkinson, Zhenduo Zhu, Derek Schlea, and Todd Redder. Invasive dreissenid mussel effects on phosphorus dynamics in Lake Ontario: insights from integrated hydrodynamic–ecological modeling. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(12):1816–1832, December 2021. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/cjfas-
2020-0164.

Maitland:2021:ETI

[326] Bryan M. Maitland, Carlos Martínez del Rio, and Frank J. Rahel. Effect of temperature on ^{13}C and ^{15}N incorporation rates and discrimination factors in two North American fishes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(12):1833–1840, December 2021. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/cjfas-2021-0057.

Barrero:2021:ESF

[327] Sandra Jeannette Hernández Barrero, Mauricio Valderrama Barco, Carlos Guillermo Barreto Reyes, and Wolfgang Stotz. Effects of selective fishing on a small scale multi-species and multi-gear freshwater fishery in the Magdalena River Basin (Colombia). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(12):1841–1854, December 2021. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/cjfas-2020-0216.

James:2021:WDC

[328] Clayton T. James, Marie F. Veillard, Amanda M. Martens, Emmanuel A. Pila, Alyssa Turnbull, Patrick Hanington, Andreas Luek, Julie Alexander, and R. Barry Nehring. Whirling disease in the Crowsnest River: an emerging threat to wild salmonids in Alberta. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(12):1855–1868, December 2021. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/cjfas-2020-0484.

Buchanan:2021:OST

[329] Rebecca A. Buchanan, Elissa Buttermore, and Joshua Israel. Outmigration survival of a threatened steelhead population through a tidal estuary. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(12):1869–1886, December 2021. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2020-0467.

Jones:2021:IVL

[330] Kim K. Jones, Trevan J. Cornwell, Daniel L. Bottom, Staci Stein, and Steven Starcevich. Interannual variability in life-stage specific survival and life history diversity of Coho salmon in a coastal Oregon basin. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(12):1887–1899, December 2021. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2020-0306.

Ziegler:2021:PWI

[331] Jacob P. Ziegler, James W. Roy, Matthew J. Bogard, and D. Andrew R. Drake. Predicting warming-induced hypoxic stress for fish in a fragmented river channel using ecosystem metabolism models. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(12):1900–1909, December 2021. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2020-0480.

Ingolfsson:2021:ALI

[332] Ólafur Arnar Ingólfsson, Terje Jørgensen, Manu Sistiaga, and Liz Kvalvik. Artificial light improves size selection for northern shrimp (Pandalus borealis) in trawls. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 78(12):1910– 1917, December 2021. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2020-0458.

Weber:2022:NEA

[333] D. Nick Weber, Andrew T. Fields, William F. Patterson, Beverly K. Barnett, Christopher M. Hollenbeck, and David S. Portnoy. Novel epigenetic age estimation in wild-caught Gulf of Mexico reef fishes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(1):1–5, January 2022. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2021-0240.

Smyth:2022:RPP

[334] Eric R. B. Smyth and D. Andrew R. Drake. The role of propagule pressure and environmental factors on the establishment of a large invasive cyprinid: black carp in the Laurentian Great Lakes basin. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(1):6–20, January 2022. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2020-0187.

Young:2022:ELR

[335] Matthew J. Young, Veronica Larwood, Justin K. Clause, Miranda Bell-Tilcock, George Whitman, Rachel Johnson, and Frederick Feyrer. Eye lenses reveal ontogenetic trophic and habitat shifts in an imperiled fish, Clear Lake hitch (Lavinia exilicauda chi). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(1):21–30, January 2022. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/cjfas-2020-0318.

Johnston:2022:CAC

[336] Fiona D. Johnston, Sean Simmons, Brett van Poorten, and Paul Venturelli. Comparative analyses with conventional surveys reveal the potential for an angler app to contribute to recreational fisheries monitoring. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(1):31–46, January 2022. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2021-0026.

Gervasi:2022:BCU

[337] Carissa L. Gervasi, Rolando O. Santos, Ryan J. Rezek, W. Ryan James, Ross E. Boucek, Christopher Bradshaw, Christopher Kavanagh, Jason Osborne, and Jennifer S. Rehage. Bottom-up conservation: using translational ecology to inform conservation priorities for a recreational fishery. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(1):47–62, January 2022. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2021-0024.

Matley:2022:SUJ

[338] Jordan K. Matley, Natalie V. Klinard, Sarah M. Larocque, Amy A. Weinz, and Scott F. Colborne. Space use of juvenile and subadult yellow perch (Perca flavescens) in the Detroit River using acoustic telemetry: incorporating variable detection ranges in vegetated areas. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(1):63–72, January 2022. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2020-0425.

Dudley:2022:UDS

[339] Peter N. Dudley, Sara N. John, Miles E. Daniels, and Eric M. Danner. Using decades of spawning data and hydraulic models to construct a temperature-dependent resource selection function for management of an endangered salmonid. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79 (1):73–81, January 2022. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2021-0022.

Kourantidou:2022:BAA

[340] Melina Kourantidou, Di Jin, and Andrew Solow. Bioeconomic analysis accounting for environmental effects in data-poor fisheries: the northern Labrador Arctic char. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79 (1):82–96, January 2022. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2021-0077.

Holub:2022:MSI

[341] Rachel Holub, Gale Bravener, and Robert McLaughlin. A managementscale investigation of consistent individual differences in behaviour and trapping bias in sea lamprey (Petromyzon marinus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(1):97–110, January 2022. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2019-0448.

Puncher:2022:IAA

[342] G. N. Puncher, A. Hanke, D. Busawon, E. V. A. Sylvester, W. Golet, L. C. Hamilton, and S. A. Pavey. Individual assignment of Atlantic bluefin tuna in the northwestern Atlantic Ocean using single nucleotide polymorphisms reveals an increasing proportion of migrants from the eastern Atlantic Ocean. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(1):111–123, January 2022. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/cjfas-
2020-0336.

Shi:2022:SCT

[343] Xiaotao Shi, Senfan Ke, Zhiying Tu, Yongmeng Wang, Junjun Tan, and Wentao Guo. Swimming capability of target fish from eight hydropower stations in China relative to fishway design. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(1):124–132, January 2022. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/cjfas-2020-0468.

Letcher:2022:CSB

[344] Benjamin H. Letcher, Keith H. Nislow, Matthew J. O’Donnell, Andrew R. Whiteley, Jason A. Coombs, and Todd L. Dubreuil. Cohort strength and body size in co-occurring salmonids in a small stream network: variation in space and time. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(1): 133–147, January 2022. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2020-0418.

Yan:2022:SMB

[345] Yuan Yan, Eva Cantoni, Chris Field, Margaret Treble, and Joanna Mills Flemming. Spatiotemporal modeling of bycatch data: methods and a practical guide through a case study in a Canadian Arctic fishery. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(1):148–158, January 2022. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2020-0267.

Essington:2022:FGM

[346] Timothy E. Essington, Mary Elizabeth Matta, Bryan A. Black, Thomas E. Helser, and Paul D. Spencer. Fitting growth models to otolith increments to reveal time-varying growth. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79 (1):159–167, January 2022. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2021-0046.

Larsen:2022:SSV

[347] Donald A. Larsen, Abby E. Fuhrman, Deborah L. Harstad, David A. Venditti, and Brian R. Beckman. Stock-specific variation in the probability of precocious male maturation in hatchery Chinook salmon (Oncorhynchus tshawytscha). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79 (1):168–182, January 2022. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2020-0461.

Evans:2022:SSA

[348] Thomas M. Evans, Zachary S. Feiner, Lars G. Rudstam, Doran M. Mason, James M. Watkins, Euan D. Reavie, Anne E. Scofield, Lyubov E. Burlakova, Alexander Y. Karatayev, and W. Gary Sprules. Size spectra analysis of a decade of Laurentian Great Lakes data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(1):183–194, January 2022. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2020-0144.

Minami:2022:ETS

[349] Kenji Minami, Hokuto Shirakawa, Yohei Kawauchi, Huamei Shao, Makoto Tomiyasu, Yuka Iwahara, Yuichi Tsuda, Hideo Takahara, Yanhui Zhu, and Kazushi Miyashita. Estimating target strength of young chum salmon (Oncorhynchus keta). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(2): 195–199, February 2022. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2021-0181.

Farrell:2022:ITR

[350] Collin J. Farrell, Brett M. Johnson, Adam G. Hansen, and Christopher A. Myrick. Induced triploidy reduces mercury bioaccumulation in a piscivorous fish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(2):200–212, February 2022. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/cjfas-
2021-0037.

Bunch:2022:BMA

[351] Aaron J. Bunch, Henry Hershey, Greg R. Reger, Frederick J. Hoogakker, Brady S. Donovan, and Troy M. Farmer. A Bayesian multistate approach to evaluate movements of an invasive freshwater estuarine-opportunist. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(2):213–222, February 2022. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2020-0378.

Naman:2022:DPF

[352] Sean M. Naman, Jordan S. Rosenfeld, and Alecia S. Lannan. Diel patterns of foraging and microhabitat use by sympatric rainbow trout and bull trout: implications for adaptive differentiation and instream flow assessment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(2):223–233, February 2022. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/cjfas-
2020-0475.

Warnock:2022:KLK

[353] Will G. Warnock, Joseph L. Thorley, Steven K. Arndt, Tyler J. Weir, Matthew D. Neufeld, Jeff A. Burrows, and Greg F. Andrusak. Kootenay Lake Kokanee (Oncorhynchus nerka) collapse into a predator pit. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(2):234–248, February 2022. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2020-0410.

Kovac:2022:ETD

[354] Anthony Kovac, Naomi K. Pleizier, and Colin J. Brauner. The effect of total dissolved gas supersaturation on gas bubble trauma in juvenile rainbow trout (Oncorhynchus mykiss), juvenile Kokanee (Oncorhynchus nerka), and two age classes of white sturgeon (Acipenser transmontanus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(2):249–256, February 2022. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2020-0448.

Khokher:2022:ELD

[355] M. Rizwan Khokher, L. Richard Little, Geoffrey N. Tuck, Daniel V. Smith, Maoying Qiao, Carlie Devine, Helen O’Neill, John J. Pogonoski, Rhys Arangio, and Dadong Wang. Early lessons in deploying cameras and artificial intelligence technology for fisheries catch monitoring: where machine learning meets commercial fishing. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(2):257–266, February 2022. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/doi/

full/10.1139/cjfas-2020-0446.

Papatheodoulou:2022:SPS

[356] Magdalene Papatheodoulou, Libor Závorka, Barbara Koeck, Neil B. Metcalfe, and Shaun S. Killen. Simulated pre-spawning catch and release of wild Atlantic salmon (Salmo salar) results in faster fungal spread and opposing effects on female and male proxies of fecundity. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(2):267–276, February 2022. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2021-0089.

Reid:2022:ERE

[357] Heather Bauer Reid and Anthony Ricciardi. Ecological responses to elevated water temperatures across invasive populations of the round goby (Neogobius melanostomus) in the Great Lakes basin. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(2):277–288, February 2022. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2021-0141.

Olson:2022:TDD

[358] Mark H. Olson, Janet M. Fischer, and Masaki Hayashi. Temporal dynamics of dissolved organic matter (DOM) in mountain lakes: the role of catchment characteristics. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79 (2):289–299, February 2022. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2020-0421.

Hussy:2022:WDY

[359] Karin Hüssy, Christoffer Moesgaard Albertsen, Jakob Hemmer-Hansen, Morten Vinther, Simon Hansen Serre, Tonny Bernt Thomsen, and Margit Eero. Where do you come from, where do you go: early life stage drift and migrations of cod inferred from otolith microchemistry and genetic population assignment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79 (2):300–313, February 2022. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2020-0409.

Edwards:2022:SRA

[360] Jena E. Edwards, Kevin J. Hedges, and Nigel E. Hussey. Seasonal residency, activity space, and use of deep-water channels by Greenland sharks (Somniosus microcephalus) in an Arctic fjord system. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(2):314–330, February 2022. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2021-0009.

Jacques:2022:BRD

[361] Nadine Jacques, Hermann Pettersen, Kristine Cerbule, Bent Herrmann, Ólafur Arnar Ingólfsson, Manu Sistiaga, Roger B. Larsen, Jesse Brinkhof, Eduardo Grimaldo, Jure Brčić, and Dagfinn Lilleng. Bycatch reduction in the deep-water shrimp (Pandalus borealis) trawl fishery by increasing codend mesh openness. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(2): 331–341, February 2022. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2021-0045.

Nicol:2022:SOS

[362] Colin L. Nicol, Jeffrey C. Jorgensen, Caleb B. Fogel, Britta Timpane-Padgham, and Timothy J. Beechie. Spatially overlapping salmon species have varied population response to early life history mortality from increased peak flows. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(2):

342–351, February 2022. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2021-0038.

Desforges:2022:ASF

[363] Jessica E. Desforges, Julia Clarke, Evaline J. Harmsen, Alex M. Jardine, Jessica A. Robichaud, Serina Serré, Prosanta Chakrabarty, Joseph R. Bennett, Dalal E. L. Hanna, John P. Smol, Trina Rytwinski, Jessica J. Taylor, André L. Martel, Amanda K. Winegardner, Jerome Marty, Mark K. Taylor, Constance M. O’Connor, Stacey A. Robinson andrea J. Reid, Irena F. Creed, Irene Gregory-Eaves, Nicolas W. R. Lapointe, and Steven J. Cooke. The alarming state of freshwater biodiversity in Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(2):352–365, February 2022. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2021-0073.

Schaefer:2022:TSS

[364] Christina M. Schaefer, David Deslauriers, and Ken M. Jeffries. The truncate soft-shell clam, Mya truncata, as a biomonitor of municipal wastewater exposure and historical anthropogenic impacts in the Canadian Arctic. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(3):367–379, March 2022. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2021-0078.

Clawson:2022:HRR

[365] Chelsea M. Clawson, Jeffrey A. Falke, Larissa L. Bailey, Joshua Rose, Anupma Prakash, and Aaron E. Martin. High-resolution remote sensing and multistate occupancy estimation identify drivers of spawning site selection in fall chum salmon (Oncorhynchus keta) across a sub-Arctic riverscape. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(3):380–394, March 2022. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/cjfas-
2021-0013.

Poulet:2022:QLS

[366] Camille Poulet, Betsy L. Barber-O’Malley, Géraldine Lassalle, and Patrick Lambert. Quantification of land–sea nutrient fluxes supplied by allis shad across the species’ range. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79 (3):395–409, March 2022. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2021-0012.

Hance:2022:DDS

[367] Dalton J. Hance, Russell W. Perry, Adam C. Pope, Arnold J. Ammann, Jason L. Hassrick, and Gabriel Hansen. From drought to deluge: spatiotemporal variation in migration routing, survival, travel time and floodplain use of an endangered migratory fish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(3):410–428, March 2022. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/cjfas-2021-0042.

Quindazzi:2022:EDT

[368] Micah J. Quindazzi, Adam P. Summers, and Francis Juanes. Efficiency is doing things right: high-throughput, automated, 3D methods in the modern era of otolith morphometrics. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(3):429–435, March 2022. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/doi/

full/10.1139/cjfas-2021-0088.

Dzul:2022:IAD

[369] Maria Christina Dzul, Charles Brandon Yackulic, William Louis Kendall, Dana L. Winkelman, Mary M. Conner, and Michael Yard. Incorporating antenna detections into abundance estimates of fish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(3):436–447, March 2022. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2021-0003.

Edwards:2022:SIM

[370] Charles T. T. Edwards, Marie-Julie Roux, and Malcolm R. Clark. Spatially integrated modelling of data-limited orange roughy (Hoplostethus atlanticus) using environmental covariates. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(3):448–457, March 2022. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/cjfas-2020-0265.

Bacheler:2022:FSB

[371] Nathan M. Bacheler, Brendan J. Runde, Kyle W. Shertzer, Jeffrey A. Buckel, and Paul J. Rudershausen. Fine-scale behavior of red snapper (Lutjanus campechanus) around bait: approach distances, bait plume dynamics, and effective fishing area. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(3):458–471, March 2022. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/doi/

full/10.1139/cjfas-2021-0044.

Zhang:2022:PNC

[372] Chen Zhang, Michael T. Brett, Jens M. Nielsen, George B. Arhonditsis, Ashley P. Ballantyne, Jackie L. Carter, Jacob Kann, Dörthe C. Müller-Navarra, Daniel E. Schindler, Jason D. Stockwell, Monika Winder, and David A. Beauchamp. Physiological and nutritional constraints on zooplankton productivity due to eutrophication and climate change predicted using a resource-based modeling approach. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(3):472–486, March 2022. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/cjfas-2021-0071.

Turschak:2022:TVN

[373] Benjamin Turschak, Charles R. Bronte, Sergiusz Czesny, Brandon Gerig, Austin Happel, Tomas O. Höök, Matthew S. Kornis, Benjamin S. Leonhardt, Bryan G. Matthias, Jacques Rinchard, and Harvey Bootsma. Temporal variation in the niche partitioning of Lake Michigan salmonines as it relates to alewife abundance and size structure. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(3):487–502, March 2022. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/cjfas-2021-0027.

Grow:2022:OHP

[374] Ryan C. Grow, Kyle D. Zimmer, Jennifer L. Cruise, Simon K. Emms, Loren M. Miller, Brian R. Herwig, David F. Staples, Angela R. Tipp, Greta M. Gerdes, and Peter C. Jacobson. Oxythermal habitat as a primary driver of ecological niche and genetic diversity in cisco (Coregonus artedi). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(3):503–517, March 2022. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2021-0059.

Evans:2022:ECS

[375] Rhian Evans, Stéphane Gauthier, and Clifford L. K. Robinson. Ecological considerations for species distribution modelling of euphausiids in the Northeast Pacific Ocean. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79 (3):518–532, March 2022. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2020-0481.

Lebot:2022:TFA

[376] Clément Lebot, Marie-Andrée Arago, Laurent Beaulaton, Gaëlle Germis, Marie Nevoux, Etienne Rivot, and Etienne Prévost. Taking full advantage of the diverse assemblage of data at hand to produce time series of abundance: a case study on Atlantic salmon populations of Brittany. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(4):533–547, April 2022. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2020-0368.

Boe:2022:LHC

[377] Kristin Bøe, Michael Power, Martha J. Robertson, J. Brian Dempson, and Ian A. Fleming. Life history contrasts in nutritional state and return probability of post-spawned Atlantic salmon. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(4):548–557, April 2022. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/cjfas-2021-0001.

Erratt:2022:PER

[378] Kevin Erratt, Irena F. Creed, Elizabeth J. Favot, Irina Todoran, Vera Tai, John P. Smol, and Charles G. Trick. Paleolimnological evidence reveals climate-related preeminence of cyanobacteria in a temperate meromictic lake. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(4):558–565, April 2022. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2021-0095.

Labaj:2022:MWS

[379] Andrew L. Labaj, Amber D. LeBlanc, Casey V. Doucet, Christina E. Tardif, and Joshua Kurek. Microlitter in the water, sediments, and mussels of the Saint John River (Wolastoq) watershed, Atlantic Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(4):566–575, April 2022. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2021-0121.

Rorvik:2022:LTI

[380] Carl Jakob Rørvik, Bjarte Bogstad, Geir Ottersen, and Olav Sigurd Kjesbu. Long-term interplay between harvest regimes and biophysical conditions may lead to persistent changes in age at sexual maturity of Northeast Arctic cod (Gadus morhua). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(4):576–586, April 2022. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/doi/

full/10.1139/cjfas-2021-0068.

vonLeesen:2022:TEC

[381] Gotje von Leesen, Bjarte Bogstad, Einar Hjörleifsson, Ulysses S. Ninnemann, and Steven E. Campana. Temperature exposure in cod driven by changes in abundance. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79 (4):587–600, April 2022. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2020-0424.

He:2022:ECC

[382] Ji X. He, Mark P. Ebener, Richard D. Clark, James R. Bence, Charles P. Madenjian, Kevin N. McDonnell, Matthew S. Kornis, and Charles R. Bronte. Estimating catch curve mortality based on relative return rates of coded wire tagged lake trout in US waters of Lake Huron. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(4):601–610, April 2022. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2021-0075.

Trochta:2022:NAD

[383] John T. Trochta, Maya L. Groner, Paul K. Hershberger, and Trevor A. Branch. A novel approach for directly incorporating disease into fish stock assessment: a case study with seroprevalence data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(4):611–630, April 2022. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2021-0094.

Panfili:2022:EVE

[384] Jacques Panfili, Clarisse Boulenger, Camille Musseau, and Alain J. Crivelli. Extreme variability in European eel growth revealed by an extended mark and recapture experiment in southern France and implications for management. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(4):631–641, April 2022. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/cjfas-
2020-0419.

Zhou:2022:SGA

[385] Can Zhou. Somatic growth of Atlantic bluefin tuna (Thunnus thynnus) under global climate variability: evidence from over 60 years of daily resolved growth increments with a simulation study. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(4):642–651, April 2022. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/cjfas-2021-0097.

Scribner:2022:MSA

[386] Kim T. Scribner, Travis O. Brenden, Robert Elliott, Michael Donofrio, Kristin Bott, Jeannette Kanefsy, Jared J. Homola, Iyob Tsehaye, James R. Bence, Edward Baker, and Nancy Auer. Mixed stock analysis of genetic compositions of lake sturgeon (Acipenser fulvescens) mixtures in Lake Michigan: hierarchical spatial heterogeneity and evidence of improving recruitment in Wisconsin spawning populations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(4):652–669, April 2022. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/cjfas-2021-0006.

Devlin:2022:AHP

[387] Robert H. Devlin, Carlo A. Biagi, Dionne Sakhrani, Takafumi Fujimoto, Rosalind A. Leggatt, Jack L. Smith, and Timothy Y. Yesaki. An assessment of hybridization potential between Atlantic and Pacific salmon. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(4):670–676, April 2022. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2021-0083.

Ramazi:2022:EMC

[388] Pouria Ramazi, Samuel M. Fischer, Julie Alexander, Clayton T. James, Andrew J. Paul, Russell Greiner, and Mark A. Lewis. Myxobolus cerebralis establishment and spread: a graphical synthesis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(4):677–691, April 2022. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2020-0352.

Lantz:2022:NGL

[389] Stephen R. Lantz, Robert A. Adair, Jon J. Amberg, Roger Bergstedt, Michael A. Boogaard, Ugo Bussy, Margaret F. Docker, Erin S. Dunlop, Alex Gonzalez, Terrance D. Hubert, Michael J. Siefkes, W. Paul Sullivan, Steve Whyard, Michael P. Wilkie, Bradley Young, and Andrew Muir. Next-generation lampricides: a three-stage process to develop improved control tools for invasive sea lamprey. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(4):692–702, April 2022. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/doi/

full/10.1139/cjfas-2020-0316.

Dunmall:2022:FJC

[390] Karen M. Dunmall, Darcy G. McNicholl, Christian E. Zimmerman, Sara E. Gilk-Baumer, Sean Burril, and Vanessa R. von Biela. First juvenile chum salmon confirms successful reproduction for Pacific salmon in the North American Arctic. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79 (5):703–707, May 2022. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2022-0006.

Cahill:2022:URD

[391] Christopher L. Cahill, Carl J. Walters, Andrew J. Paul, Michael G. Sullivan, and John R. Post. Unveiling the recovery dynamics of walleye after the invisible collapse. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(5):708–723, May 2022. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/cjfas-
2021-0065.

Anas:2022:PDN

[392] M. U. Mohamed Anas and Nicholas E. Mandrak. Patterns and drivers of native, non-native, and at-risk freshwater fish richness in Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(5):724–737, May 2022. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2020-0396.

Blanluet:2022:DBS

[393] Arthur Blanluet, Sven Gastauer, Franck Cattanéo, Chloé Goulon, David Grimardias, and Jean Guillard. Discrimination between schools and submerged trees in reservoirs: a preliminary approach using narrowband and broadband acoustics. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(5):738–748, May 2022. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/cjfas-
2021-0087.

VanDoornik:2022:IPP

[394] Donald M. Van Doornik, Barry A. Berejikian, Megan E. Moore, Andrew M. Claiborne, Mark R. Downen, Joy Lee Waltermire, Katy K. Doctor, and Rob Endicott. The influences of pre- and post-smolt captive rearing environments on growth, maturation, body size, and reproductive success of steelhead (Oncorhynchus mykiss) released as adults. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(5):749–760, May 2022. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2021-0028.

Harder:2022:GSA

[395] Avril M. Harder and Mark R. Christie. Genomic signatures of adaptation to novel environments: hatchery and life history-associated loci in landlocked and anadromous Atlantic salmon (Salmo salar). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(5):761–770, May 2022. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2021-0066.

Berggren:2022:WWI

[396] Terese Berggren, Ulf Bergström, Göran Sundblad, and Örjan Östman. Warmer water increases early body growth of northern pike (Esox lucius), but mortality has larger impact on decreasing body sizes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(5):771–781, May 2022. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2020-0386.

Bouchard:2022:SPD

[397] Colin Bouchard, Mathieu Buoro, Clément Lebot, and Stephanie M. Carlson. Synchrony in population dynamics of juvenile Atlantic salmon: analyzing spatiotemporal variation and the influence of river flow and demography. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(5):782–794, May 2022. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2021-0017.

vanderLee:2022:QST

[398] Adam S. van der Lee, Mark R. Vinson, and Marten A. Koops. Quantifying status and trends from monitoring surveys: application to pygmy whitefish (Prosopium coulterii) in Lake Superior. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(5):795–802, May 2022. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/doi/

full/10.1139/cjfas-2021-0155.

McClelland:2022:AWG

[399] Erin K. McClelland, Breanna Watson, L. Fredrik Sundström, Rosalind A. Leggatt, Dionne Sakhrani, and Robert H. Devlin. Assessing wild genetic background and parental effects on size of growth hormone transgenic Coho salmon. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(5):803–813, May 2022. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/cjfas-
2021-0153.

Hernandez:2022:SSS

[400] Christina M. Hernández, David E. Richardson, Irina I. Rypina, Ke Chen, Katrin E. Marancik, Kathryn Shulzitski, and Joel K. Llopiz. Support for the slope sea as a major spawning ground for Atlantic bluefin tuna: evidence from larval abundance, growth rates, and particle-tracking simulations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(5):814–824, May 2022. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2020-0444.

Gallagher:2022:HPB

[401] Colin P. Gallagher, Matthew M. Guzzo, and Terry A. Dick. High prevalence of basin fidelity and homing by lake trout (Salvelinus namaycush) in a small northern lake. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(5):825–833, May 2022. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/cjfas-
2021-0219.

Sistiaga:2022:PCS

[402] Manu Sistiaga, Jesse Brinkhof, Bent Herrmann, Roger B. Larsen, Eduardo Grimaldo, Kristine Cerbule, Ilmar Brinkhof, and Terje Jørgensen. Potential for codends with shortened lastridge ropes to replace mandated selection devices in demersal trawl fisheries. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(5):834–849, May 2022. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/doi/

full/10.1139/cjfas-2021-0178.

Steele:2022:AAG

[403] Reid W. Steele and Anna B. Neuheimer. Assessing the ability of the growing degree-day metric to explain variation in size-at-age and duration-to-moult of lobsters and crabs. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79 (5):850–860, May 2022. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2020-0310.

Honsey:2022:CDW

[404] Andrew E. Honsey, Zachary S. Feiner, and Gretchen J. A. Hansen. Correction: Drivers of walleye recruitment in Minnesota’s large lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(5):861, May 2022. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2022-0063. See [162].

Barajas:2022:RAM

[405] Miguel F. Barajas, Timothy F. Sheehan, Ruth E. Haas-Castro, Brandon Ellingson, and Katherine E. Mills. Retrospective analysis of marine growth and relationships to return rates of Penobscot River Atlantic salmon. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(6):863–874, June 2022. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2021-0142.

Bryson:2022:FWI

[406] Grace E. Bryson, Karen A. Kidd, and Kurt M. Samways. Food web incorporation of marine-derived nutrients after the reintroduction of endangered inner Bay of Fundy Atlantic salmon (Salmo salar). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(6):875–882, June 2022. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2020-0326.

Grant:2022:ICR

[407] James W. A. Grant and Laura K. Weir. Interspecific competition reduces the performance of Atlantic salmon (Salmo salar): implications for restoration programs. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(6):883–892, June 2022. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/cjfas-
2021-0163.

Fernandes:2022:TSW

[408] Timothy J. Fernandes, Brian J. Shuter, Peter E. Ihssen, and Bailey C. McMeans. The timing of spring warming shapes reproductive effort in a warm-water fish: the role of mismatches between hepatic and gonadal processes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(6):893–911, June 2022. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/cjfas-
2020-0412.

Yochum:2022:ECS

[409] Noëlle Yochum, David R. Bryan, Lyle L. Britt, Barry A. Berejikian, Rebecca Haehn, Scott McEntire, Rick Towler, Jeff Atkins, Brad Gadberry, and Paul Irvin. Evaluating Chinook salmon (Oncorhynchus tshawytscha) response to artificial light in support of bycatch mitigation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(6):912–924, June 2022. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2021-0241.

Izzo:2022:CFL

[410] Lisa K. Izzo, Gayle Barbin Zydlewski, and Donna L. Parrish. Combining fixed-location count data and movement data to estimate abundance of a lake sturgeon spawning run: a framework for riverine migratory species. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(6):925–935, June 2022. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2021-0140.

Kovach:2022:GVW

[411] Ryan P. Kovach, Robb F. Leary, Donovan A. Bell, Sally Painter, Angela Lodmell, and Andrew R. Whiteley. Genetic variation in westslope cutthroat trout reveals that widespread genetic rescue is warranted. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(6):936–946, June 2022. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2021-0102.

Uusi-Heikkila:2022:FTT

[412] Silva Uusi-Heikkilä, Tommi Perälä, and Anna Kuparinen. Fishing triggers trophic cascade in terms of variation, not abundance, in an allometric trophic network model. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79 (6):947–957, June 2022. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2021-0146.

Abookire:2022:PSA

[413] Alisa A. Abookire, Michael A. Litzow, Michael J. Malick, and Benjamin J. Laurel. Post-settlement abundance, condition, and survival in a climatestressed population of Pacific cod. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(6):958–968, June 2022. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/doi/

full/10.1139/cjfas-2021-0224.

Zimmermann:2022:IRF

[414] Fabian Zimmermann, Alf Ring Kleiven, Merete Vik Ottesen, and Guldborg Søvik. Inclusion of recreational fishing in data-limited stocks: a case study on Norway lobster (Nephrops norvegicus) in Norway. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(6):969–978, June 2022. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2021-0152.

Jensen:2022:ETR

[415] Alexander J. Jensen, Benjamin Cox, and James T. Peterson. Evaluating tag-reliant harvest estimators in Chinook salmon mixed-stock fisheries using simulations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(6):979–993, June 2022. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/cjfas-
2021-0197.

Morrison:2022:RHD

[416] Brian P. Morrison and Kathryn S. Peiman. The role of hatcheries in the decline of Lake Ontario Atlantic salmon. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(6):994–1002, June 2022. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/doi/

full/10.1139/cjfas-2021-0253.

Quinn:2022:DCU

[417] Thomas P. Quinn and James P. Losee. Diverse and changing use of the Salish Sea by Pacific salmon, trout, and char. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(6):1003–1021, June 2022. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/cjfas-2021-0162.

FitzGerald:2022:PBC

[418] Alyssa M. FitzGerald, David A. Boughton, Joshua Fuller, Sara N. John, Benjamin T. Martin, Lee R. Harrison, and Nathan J. Mantua. Physical and biological constraints on the capacity for life-history expression of anadromous salmonids: an Eel River, California, case study. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(7):1023–1041, July 2022. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2021-0229.

Chila:2022:IKP

[419] Zander Chila, Karen M. Dunmall, Tracey A. Proverbs, Trevor C. Lantz, Aklavik Hunters and Trappers Committee, Inuvik Hunters and Trappers Committee, Sachs Harbour Hunters and Trappers Committee, Olokhaktomiut Hunters and Trappers Committee, and Paulatuk Hunters and Trappers Committee. Inuvialuit knowledge of Pacific salmon range expansion in the western Canadian Arctic. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(7):1042–1055, July 2022. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/doi/

full/10.1139/cjfas-2021-0172.

Watras:2022:MWB

[420] Carl J. Watras, James R. Michler, and Jeff L. Rubsam. Monitoring the water balance of seepage lakes to track regional responses to an evolving climate. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(7):1056–1063, July 2022. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/cjfas-
2021-0217.

Lejeune:2022:GCM

[421] Benjamin Lejeune, Maud Aline Mouchet, Sonia Mehault, and Dorothée Kopp. Gut content metabarcoding reveals potential importance of fisheries discards consumption in marine fauna. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(7):1064–1073, July 2022. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/doi/

full/10.1139/cjfas-2021-0267.

Dubos:2022:SSH

[422] Véronique Dubos, André St-Hilaire, and Normand E. Bergeron. Summer stream habitat preferences of Nunavik anadromous Arctic char (Salvelinus alpinus) fry and parr. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79 (7):1074–1085, July 2022. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2021-0119.

Barrett:2022:STD

[423] Timothy J. Barrett, Adrian R. Hordyk, Melanie A. Barrett, and Michael R. van den Heuvel. Spatial and temporal differences in fecundity of Atlantic herring (Clupea harengus) off Nova Scotia and consequences for biological reference points. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79 (7):1086–1096, July 2022. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2021-0269.

Langan:2022:DPR

[424] Joseph A. Langan, Jeremy S. Collie, and Cóilín Minto. Detecting population regulation of winter flounder from noisy data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(7):1097–1108, July 2022. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2021-0004.

Blincow:2022:ESS

[425] Kayla M. Blincow and Brice X. Semmens. The effect of sea surface temperature on the structure and connectivity of species landings interaction networks in a multispecies recreational fishery. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(7):1109–1119, July 2022. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/cjfas-2021-0085.

Zentner:2022:MLA

[426] Douglas L. Zentner, Joshua K. Raabe, Timothy K. Cross, and Peter C. Jacobson. Machine learning applied to lentic habitat use by spawning walleye demonstrates the benefits of considering multiple spatial scales in aquatic research. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(7):1120–1137, July 2022. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/cjfas-
2021-0180.

Phillips:2022:OTS

[427] Joseph S. Phillips, Gudni Gudbergsson, and Anthony R. Ives. Opposing trends in survival and recruitment slow the recovery of a historically overexploited fishery. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(7):1138–1144, July 2022. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/cjfas-
2021-0161.

Lobon-Cervia:2022:DRT

[428] Javier Lobón-Cerviá. Does recruitment trigger negative density-dependent feedback loops in stream-dwelling salmonids? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(7):1145–1153, July 2022. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/cjfas-2021-0055.

Madenjian:2022:RAA

[429] Charles P. Madenjian. Response of alewife abundance to the bacterial kidney disease outbreak in the Chinook salmon population of Lake Michigan: importance of predation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79 (7):1154–1161, July 2022. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2021-0244.

Korosi:2022:PPS

[430] Jennifer B. Korosi, Kristen A. Coleman, Grace N. Hoskin, Amanda J. Little, Emily M. Stewart, and Joshua R. Thienpont. Paleolimnological perspectives on the shifting geographic template of permafrost landscapes and its implications for Arctic freshwater biodiversity. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(7):1162–1172, July 2022. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2021-0280.

Rose:2022:CSP

[431] George A. Rose and Sherrylynn Rowe. Congruence of stock production and assessment areas? An historical perspective on Canada’s iconic Northern cod. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(7):1173–1187, July 2022. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/cjfas-
2021-0230.

Brooks:2022:ISO

[432] Mollie E. Brooks, Valentina Melli, Esther Savina, Juan Santos, Russell Millar, Finbarr Gerard O’Neill, Tiago Veiga-Malta, Ludvig Ahm Krag, and Jordan Paul Feekings. Introducing selfisher: open source software for statistical analyses of fishing gear selectivity. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(8):1189–1197, August 2022. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/cjfas-2021-0099.

Ovando:2022:IFS

[433] Daniel Ovando, Curry Cunningham, Peter Kuriyama, Christopher Boatright, and Ray Hilborn. Improving forecasts of sockeye salmon (Oncorhynchus nerka) with parametric and nonparametric models. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(8):1198–1210, August 2022. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2021-0287.

Trudeau:2022:LPL

[434] Ashley Trudeau, Eleanor A. Bochenek, Abigail S. Golden, Michael C. Melnychuk, Douglas R. Zemeckis, and Olaf P. Jensen. Lower possession limits and shorter seasons directly reduce for-hire fishing effort in a multispecies marine recreational fishery. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79 (8):1211–1224, August 2022. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2021-0137.

Bateman:2022:ASF

[435] Andrew W. Bateman, Amy K. Teffer, Arthur Bass, Tobi Ming, Karia Kaukinen, Brian P. V. Hunt, Martin Krkošek, and Kristina M. Miller. Atlantic salmon farms are a likely source of Tenacibaculum maritimum infection in migratory Fraser River sockeye salmon. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(8):1225–1240, August 2022. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/cjfas-2021-0164.

Solinger:2022:OBS

[436] Laura K. Solinger, Kathy A. Ashton-Alcox, Eric N. Powell, Kathleen M. Hemeon, Sara M. Pace, Thomas M. Soniat, and Leanne M. Poussard. Oysters beget shell and vice versa: generating management goals for live oysters and the associated reef to promote maximum sustainable yield of Crassostrea virginica. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(8): 1241–1254, August 2022. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2021-0277.

Bolser:2022:ASS

[437] Derek G. Bolser, Jack P. Egerton, Philip M. Souza, Jr., Kevin M. Boswell, and Brad E. Erisman. Assessing the size spectra of marine fish communities with hydroacoustics: examining the challenges of abundant schools, diverse assemblages, and variable orientations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(8):1255–1268, August 2022. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/cjfas-2021-0189.

Johnson:2022:CTM

[438] Fred A. Johnson, William E. Pine III, and Edward V. Camp. A cautionary tale: management implications of critical transitions in oyster fisheries. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(8):1269–1281, August 2022. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2021-0133.

Claireaux:2022:EDH

[439] Marion Claireaux, Fabian Zimmermann, Bruno Ernande, Mikko Heino, and Katja Enberg. Environmental drivers of herring growth and how the perception shifts with time series length. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(8):1282–1290, August 2022. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/cjfas-2021-0176.

Bandara:2022:SBA

[440] Tharindu Bandara, Sonia Brugel, Agneta Andersson, and Danny Chun Pong Lau. Seawater browning alters community composition and reduces nutritional quality of plankton in a subarctic marine ecosystem. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(8):1291–1301, August 2022. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2021-0118.

Perez-Rodriguez:2022:EHC

[441] Alfonso Pérez-Rodríguez, Ibrahim Umar, Daisuke Goto, Daniel Howell, Iago Mosqueira, and Diana González-Troncoso. Evaluation of harvest control rules for a group of interacting commercial stocks using a multispecies MSE framework. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79 (8):1302–1320, August 2022. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2021-0069.

Medinski:2022:CCM

[442] Nathan A. Medinski, Bryan M. Maitland, Timothy D. Jardine, D. Andrew R. Drake, and Mark S. Poesch. A catastrophic coal mine spill in the Athabasca River watershed induces isotopic niche shifts in stream biota including an endangered rainbow trout ecotype. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(8):1321–1334, August 2022. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/cjfas-2021-0112. See correction [633].

He:2022:FCS

[443] Xiaoping He, Ryan R. E. Stanley, Emily M. Rubidge, Nicholas W. Jeffery, Lorraine C. Hamilton, Kristen M. Westfall, Scott R. Gilmore, Louise-Marie D. Roux, Katie S. P. Gale, Susan G. Heaslip, Royce Steeves, and Cathryn L. Abbott. Fish community surveys in eelgrass beds using both eDNA metabarcoding and seining: implications for biodiversity monitoring in the coastal zone. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(8):

1335–1346, August 2022. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2021-0215.

Brownscombe:2022:ATF

[444] Jacob W. Brownscombe, Lucas P. Griffin, Jill L. Brooks, Andy J. Danylchuk, Steven J. Cooke, and Jonathan D. Midwood. Applications of telemetry to fish habitat science and management. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(8):1347–1359, August 2022. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2021-0101.

Budy:2022:CAP

[445] Phaedra Budy, Gary P. Thiede, Kevin Chapman, and Frank P. Howe. Conflicts among protected native birds and valuable sport fishes: potential pelican predation effects. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(9): 1361–1376, September 2022. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2021-0228.

Shea:2022:EDD

[446] Dylan Shea, Neil Frazer, Kiran Wadhawan, Andrew Bateman, Shaorong Li, Kristina M. Miller, Steven Short, and Martin Krkošek. Environmental DNA dispersal from Atlantic salmon farms. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(9):1377–1388, September 2022. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2021-0216.

Huntsman:2022:JSA

[447] Brock M. Huntsman, Larry R. Brown, Kai Palenscar, Chris Jones, Kerwin Russell, Heather Dyer, Brett Mills, Marissa L. Wulff, and Jason May. Joint-species analysis reveals potential displacement of native fishes by non-native fishes within the Santa Ana River, California. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(9):1389–1406, September 2022. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2021-0210.

Massey:2022:DVM

[448] Yoluène Massey, Philippe S. Sabarros, and Pascal Bach. Drivers of atvessel mortality of the blue shark (Prionace glauca) and oceanic whitetip shark (Carcharhinus longimanus) assessed from monitored pelagic longline experiments. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(9):1407– 1419, September 2022. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2021-0273.

Gale:2022:ECD

[449] Katie S. P. Gale, Robyn E. Forrest, Christopher N. Rooper, Jessica Nephin, Scott Wallace, John Driscoll, and Emily Rubidge. Evaluation of collaboratively developed management measures to reduce coral and sponge bycatch in a fully monitored multispecies trawl fishery. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(9):1420–1434, September 2022. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2021-0274.

Wszola:2022:FRS

[450] Lyndsie S. Wszola, Zachary S. Feiner, Christopher J. Chizinski, Jamilynn B. Poletto, and John P. DeLong. Fishing regulations, sexual dimorphism, and the life history of harvest. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(9):1435–1446, September 2022. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/cjfas-2021-0248.

Hollins:2022:DFM

[451] Jack Hollins, Harri Pettitt-Wade, Colin P. Gallagher, Ellen V. Lea, Lisa L. Loseto, and Nigel E. Hussey. Distinct freshwater migratory pathways in Arctic char (Salvelinus alpinus) coincide with separate patterns of marine spatial habitat-use across a large coastal landscape. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(9):1447–1464, September 2022. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2021-0291.

Cook:2022:REL

[452] Nathan A. Cook and Samuel L. Bourret. Revealing early life history of brown trout in an impaired watershed through otolith geochemistry.

Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(9):1465–1478, September 2022. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2021-0064.

Brown:2022:HRK

[453] Norah E. M. Brown and Thomas W. Therriault. The hidden risk of keystone invaders in Canada: a case study using nonindigenous crayfish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(9):1479–1496, September 2022. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2021-0245.

Peterson:2022:EUI

[454] Cassidy D. Peterson, Michael J. Wilberg, Enric Cortés, Dean L. Courtney, and Robert J. Latour. Effects of unregulated international fishing on recovery potential of the sandbar shark within the southeastern United States. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(9):1497– 1513, September 2022. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2021-0345.

Valleau:2022:UDT

[455] Robin E. Valleau, Kathleen M. Rühland, Andrew M. Paterson, and John P. Smol. Using diatoms to track road-salt seepage into small, shallow, softwater Ontario lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(9): 1514–1528, September 2022. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2021-0072.

Fitzpatrick:2022:BPA

[456] Kimberly B. Fitzpatrick, Brian C. Weidel, Michael J. Connerton, Jana R. Lantry, Jeremy P. Holden, Michael J. Yuille, Brian Lantry, Steven R. LaPan, Lars G. Rudstam, Patrick J. Sullivan, Travis O. Brenden, and Suresh A. Sethi. Balancing prey availability and predator consumption: a multispecies stock assessment for Lake Ontario. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(9):1529–1545, September 2022. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2021-0126.

Osowski:2022:RSE

[457] Ana R. Osowski and Stephen T. Szedlmayer. Red snapper (Lutjanus campechanus) abundance on oil and gas platforms based on mark– recapture methods in the northern Gulf of Mexico. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(9):1546–1560, September 2022. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2021-0227.

Horricks:2022:IBA

[458] Ryan A. Horricks, Leah M. Lewis-McCrea, and Gregor K. Reid. Interactions between American lobster (Homarus americanus) and salmonid aquaculture in the Canadian Maritimes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(9):1561–1571, September 2022. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2021-0252.

Cadigan:2022:NAR

[459] Noel G. Cadigan, D. E. Duplisea, C. Senay, G. J. Parent, P. D. Winger, B. Linton, and K. Kristinsson. Northwest Atlantic redfish science priorities for managing an enigmatic species complex. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(9):1572–1589, September 2022. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2021-0266.

Feiner:2022:PEE

[460] Zachary S. Feiner, Hilary A. Dugan, Noah R. Lottig, Greg G. Sass, and Gretchen A. Gerrish. A perspective on the ecological and evolutionary consequences of phenological variability in lake ice on north-temperate lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(9):1590–1604, September 2022. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2021-0221.

Batts:2022:NMC

[461] Luke Batts, Cóilín Minto, Hans Gerritsen, and Deirdre Brophy. Numbers or mass? Comparison of two theoretically different stage-based stock assessment models and their ability to model simulated and real-life stocks. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(10):1605–1624, October 2022.

CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2021-0213.

Zimmerman:2022:DLH

[462] Mara S. Zimmerman, Matthew R. Sloat, Kirill V. Kuzishchin, Martin C. Arostegui, Marina A. Gruzdeva, Todd R. Seamons, and Thomas P. Quinn. Diversity of life history traits, growth, and lipid storage in migratory variants of steelhead and rainbow trout (Oncorhynchus mykiss) in Kamchatka, Russia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(10):1625–1640, October 2022. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/cjfas-
2021-0357.

Paradinas:2022:ESS

[463] Iosu Paradinas, Joan Giménez, David Conesa, Antonio López-Quílez, and Maria Grazia Pennino. Evidence for spatiotemporal shift in demersal fishery management priority areas in the western Mediterranean. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(10):1641–1654, October 2022. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2021-0327.

Oke:2022:WLO

[464] Krista B. Oke, Franz Mueter, and Michael A. Litzow. Warming leads to opposite patterns in weight-at-age for young versus old age classes of Bering Sea walleye pollock. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79 (10):1655–1666, October 2022. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/doi/

full/10.1139/cjfas-2021-0315.

Li:2022:ITH

[465] Chengxue Li, Lorenzo Ciannelli, Morgan Bancroft, Jay Rooker, Clifford Ryer, and Hui Liu. Interplay of temperature and hypoxia in habitat quality for a juvenile demersal fish in a coastal upwelling system. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(10):1667–1680, October 2022. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2021-0295.

Huard:2022:IBH

[466] Jacqueline R. Huard, Beatrice Proudfoot, Christopher N. Rooper, Tara G. Martin, and Clifford L. K. Robinson. Intertidal beach habitat suitability model for Pacific sand lance (Ammodytes personatus) in the Salish Sea, Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(10):1681–1696, October 2022. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/cjfas-
2021-0335.

Tirronen:2022:ASS

[467] Maria Tirronen, Jeffrey A. Hutchings, Sebastián A. Pardo, and Anna Kuparinen. Atlantic salmon survival at sea: temporal changes that lack regional synchrony. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(10): 1697–1711, October 2022. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2021-0302.

Charapata:2022:LHP

[468] Patrick Charapata, Dion Oxman, Kevin McNeel, Alexa Keith, Farzaneh Mansouri, and Stephen Trumble. Lifetime hormone profiles for a longlived teleost: opercula reveal novel estimates of age-specific reproductive parameters and stress trends in yelloweye rockfish (Sebastes ruberrimus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(10):1712–1728, October 2022. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2022-0048.

Al-Chokhachy:2022:SST

[469] Robert Al-Chokhachy, Benjamin H. Letcher, Clint C. Muhlfeld, Jason B. Dunham, Timothy Cline, Nathaniel P. Hitt, James J. Roberts, and David Schmetterling. Stream size, temperature, and density explain body sizes of freshwater salmonids across a range of climate conditions. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(10):1729–1744, October 2022. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2021-0343.

Fisch:2022:ECA

[470] Nicholas Fisch, Robert Ahrens, Kyle Shertzer, and Ed Camp. An empirical comparison of alternative likelihood formulations for composition data, with application to cobia and Pacific hake. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(10):1745–1764, October 2022. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2022-0036.

Petrie:2022:TEC

[471] Brian Petrie, Kenneth T. Frank, and William C. Leggett. Temporal evolution of critical traits and their relationship to cod stock collapse and recovery. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(11):1765–1779, November 2022. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/cjfas-
2022-0028.

Bailey:2022:FLB

[472] Colin J. Bailey, Kyle L. Wilson, Mikayla More O’Ferrall, and Jonathan W. Moore. Fish length back-calculation from scales: advancing methodology and correction of bias. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(11): 1780–1797, November 2022. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2021-0270.

Cadigan:2022:CRB

[473] Noel Cadigan, Matthew D. Robertson, Kunasekaran Nirmalkanna, and Nan Zheng. The complex relationship between weight and length of Atlantic cod off the south coast of Newfoundland. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(11):1798–1819, November 2022. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2021-0325.

Bak-Jensen:2022:FMS

[474] Zita Bak-Jensen, Bent Herrmann, Juan Santos, Nadine Jacques, Valentina Melli, and Jordan P. Feekings. Fixed mesh shape reduces variability in codend size selection. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(11): 1820–1829, November 2022. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2022-0049.

Lees:2022:ESR

[475] K. J. Lees, M. A. MacNeil, K. J. Hedges, and N. E. Hussey. Estimating survival in a remote community-based fishery using acoustic telemetry. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(11):1830–1842, November 2022. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2021-0288.

Bowlby:2022:CIW

[476] Heather D. Bowlby, Warren N. Joyce, Megan V. Winton, Peterson J. Coates, and Gregory B. Skomal. Conservation implications of white shark (Carcharodon carcharias) behaviour at the northern extent of their range in the Northwest Atlantic. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79 (11):1843–1859, November 2022. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/doi/

full/10.1139/cjfas-2021-0313.

Freshwater:2022:NPD

[477] Cameron Freshwater, Charles K. Parken, Strahan Tucker, Antonio Velez-Espino, and Jackie King. Nonstationary patterns in demographic traits covary with Chinook salmon marine distributions. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(11):1860–1878, November 2022. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2021-0312.

Courter:2022:HPD

[478] Ian I. Courter, Tom Chance, Ryan Gerstenberger, Mark Roes, Sean Gibbs, and Adrian Spidle. Hatchery propagation did not reduce natural steelhead productivity relative to habitat conditions and predation in a mid-Columbia River subbasin. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79 (11):1879–1895, November 2022. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/doi/ full/10.1139/cjfas-2021-0351. See comment [532] and response [532] and response [533].

Yin:2022:IEL

[479] Yihao Yin, Jessica A. Sameoto, David M. Keith, and Joanna Mills Flemming. Improving estimation of length–weight relationships using spatiotemporal models. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(11): 1896–1910, November 2022. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2021-0317.

Boldt:2022:CFF

[480] Jennifer L. Boldt, Hannah M. Murphy, Jean-Martin Chamberland, Allan Debertin, Stéphane Gauthier, Brooke Hackett, Paige S. Hagel, Andrew R.

Majewski, Jenni L. McDermid, David Mérette, Cliff L. K. Robinson, Christopher N. Rooper, Bryanna Sherbo, Elisabeth Van Beveren, and Wojciech Walkusz. Canada’s forage fish: an important but poorly understood component of marine ecosystems. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(11):1911–1933, November 2022. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/cjfas-2022-0060.

Stevenson:2022:IAC

[481] Duane E. Stevenson, Stan Kotwicki, James T. Thorson, Giancarlo M. Correa, and Troy Buckley. The influence of age and cohort on the distribution of walleye pollock (Gadus chalcogrammus) in the eastern Bering Sea. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(11):1934–1949, November 2022. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2021-0300.

Vagnon:2022:VWE

[482] Chloé Vagnon, Franck Cattanéo, Chloé Goulon, Jean Guillard, and Victor Frossard. The vulnerability of whitefish (Coregonus lavaretus) to the invasive European catfish (Silurus glanis) in a large peri-alpine lake. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(11):1950–1960, November 2022. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2021-0310.

Hansen:2022:IPA

[483] Henry H. Hansen, Stephen D. Kachman, Mark A. Pegg, Colin Charles, Douglas A. Watkinson, and Eva C. Enders. Informative priors assess tradeoffs between mark–recapture and telemetry-based fish movement in a large river system. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(11): 1961–1976, November 2022. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2021-0238.

Figueiredo:2022:LCM

[484] Ivone Figueiredo, Isabel Natário, Pascal Lorance, and M. Lucília Carvalho. A life cycle model to assess the abundance of black scabbardfish, a widely distributed fish with cryptic migrations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(11):1977–1991, November 2022. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2021-0272.

Burchfield:2022:SAL

[485] James D. Burchfield, Brian E. McLaren, and Darryl T. McLeod. Sensitivity analysis of a lake sturgeon population with early life stage densitydependent effects. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(11):1992– 2005, November 2022. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2021-0339.

Church:2022:PHA

[486] Kathleen D. W. Church, Lida Nguyen-Dang, Bryan D. Neff, and Christina A. D. Semeniuk. Personality in hatchery Atlantic salmon (Salmo salar) is associated with growth but not metabolic rate or performance. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(11):2006–2024, November 2022. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2022-0041.

Anonymous:2022:NR

[487] Anonymous. Note de reconnaissance. (French) [Editorial note]. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(12), December 2022. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2022-0274.

Lamarins:2022:IDA

[488] Amaïa Lamarins, Florèn Hugon, Cyril Piou, Julien Papaïx, Etienne Prévost, Stephanie M. Carlson, and Mathieu Buoro. Implications of dispersal in Atlantic salmon: lessons from a demo-genetic agent-based model. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(12):2025–2042, December 2022. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2021-0342.

Rogers:2022:TTC

[489] Kevin B. Rogers, Brendon J. Sucher, Brian W. Hodge, and Christopher A. Myrick. Thermal tolerance in cutthroat trout of the southern Rocky Mountains. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(12): 2043–2055, December 2022. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2021-0352.

Hause:2022:STS

[490] Colby L. Hause, Gabriel P. Singer, Rebecca A. Buchanan, Dennis E. Cocherell, Nann A. Fangue, and Andrew L. Rypel. Survival of a threatened salmon is linked to spatial variability in river conditions. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(12):2056–2071, December 2022. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2021-0243.

Atlas:2022:QRP

[491] W. I. Atlas, K. L. Wilson, C. K. Whitney, J. E. Moody, C. N. Service, M. Reid, and M. R. Sloat. Quantifying regional patterns of collapse in British Columbia Central Coast chum salmon (Oncorhynchus keta) populations since 1960. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(12): 2072–2086, December 2022. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2022-0013.

Quinn:2022:MEC

[492] Brady K. Quinn, Marc Trudel, Brent M. Wilson, Jonathan Carr, Jason Daniels, Susan Haigh, David C. Hardie, James P. Hawkes, Christopher W. McKindsey, Mitchell O’Flaherty-Sproul, Émilie Simard, and Fred Page. Modelling the effects of currents and migratory behaviours on the dispersal of Atlantic salmon (Salmo salar) post-smolts in a coastal embayment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(12):2087–2111, December 2022. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2021-0316.

Hodgdon:2022:DFE

[493] Cameron T. Hodgdon, Burton Shank, and Yong Chen. Developing a framework to estimate dynamic reference points for American lobster using a thermally explicit spawning stock biomass/recruitment relationship. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(12):2112–2125, December 2022. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2022-0004.

Houle:2022:RIL

[494] Daniel Houle, Fougère Augustin, and Suzanne Couture. Rapid improvement of lake acid–base status in Atlantic Canada following steep decline in precipitation acidity. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(12): 2126–2137, December 2022. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2021-0349.

Ospina-Alvarez:2022:RRI

[495] Andrés Ospina-Alvarez, Sebastián I. Vásquez, Ignacio A. Catalán, Susan Lowerre-Barbieri, Marcos Arteaga, and Silvia de Juan. A reproductive resilience index for pelagic fish in the southern Humboldt Current Large Marine Ecosystem. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79 (12):2138–2159, December 2022. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/doi/

full/10.1139/cjfas-2021-0263.

Tyler:2022:IDT

[496] Kyle J. Tyler, Dion Wedd, David A. Crook, Mark J. Kennard, and Alison J. King. Identifying drivers of tropical riverine larval fish abundance and diversity. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(12): 2160–2178, December 2022. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2021-0233.

Tuser:2022:CFS

[497] Michal Tušer, Jean Guillard, Atle Rustadbakken, and Thomas Mehner. Comparison of fish size spectra obtained from hydroacoustics and gillnets across seven European natural lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(12):2179–2190, December 2022. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/cjfas-2022-0037.

Munaweera:2022:ESP

[498] Inesh Munaweera, Les N. Harris, Jean-Sébastien Moore, Ross F. Tallman, Aaron T. Fisk, Darren M. Gillis, and Saman Muthukumarana. Estimating survival probabilities of Cambridge Bay Arctic char using acoustic telemetry data and Bayesian multistate capture–recapture models. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(12):2191–2203, December 2022. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2021-0262.

Ohms:2022:PDP

[499] Haley A. Ohms, Dereka N. Chargualaf, Gabriel Brooks, Cory Hamilton, Eric P. Palkovacs, and David A. Boughton. Poor downstream passage at a dam creates an ecological trap for migratory fish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(12):2204–2215, December 2022. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2022-0095.

Nash:2022:TTF

[500] A. J. Nash, K. W. Vollset, E. M. Hanssen, S. Berhe, A. G. Salvanes, T. E. Isaksen, B. T. Barlaup, and R. J. Lennox. A tale of two fishes: depth preference of migrating Atlantic salmon smolt and predatory brown trout in a Norwegian lake. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79 (12):2216–2224, December 2022. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/doi/

full/10.1139/cjfas-2022-0016.

Charbonneau:2022:EFM

[501] Julie A. Charbonneau, David M. Keith, M. Aaron MacNeil, and Jeffrey A. Hutchings. Effects of fishing mortality on the age structure of marine fishes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(12):2225–2236, December 2022. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/cjfas-
2021-0338.

Molina-Moctezuma:2022:EDI

[502] Alejandro Molina-Moctezuma, Daniel S. Stich, and Joseph D. Zydlewski. Effects of dam-induced delays on system-wide survival of Atlantic salmon smolts during high-flow, high-survival years in the Penobscot River, Maine, USA. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 79(12):2237– 2250, December 2022. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2022-0055.

Zillig:2023:IVT

[503] Kenneth W. Zillig, Robert A. Lusardi, Dennis E. Cocherell, and Nann A. Fangue. Interpopulation variation in thermal physiology among seasonal runs of Chinook salmon. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80 (1):1–13, January 2023. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2022-0133.

Rohtla:2023:HUG

[504] Mehis Rohtla, Françoise Daverat, Michael T. Arts, Howard I. Browman, Camilla Parzanini, Anne Berit Skiftesvik, Eva B. Thorstad, Terje van der Meeren, Leif Asbjørn Vøllestad, and Caroline M. F. Durif. Habitat use and growth of yellow-stage European eel in coastal and freshwater ecosystems in Norway. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80(1):14–26, January 2023. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/cjfas-
2022-0033.

Legault:2023:DRM

[505] Christopher M. Legault, John Wiedenmann, Jonathan J. Deroba, Gavin Fay, Timothy J. Miller, Elizabeth N. Brooks, Richard J. Bell, Joseph A. Langan, Jamie M. Cournane, Andrew W. Jones, and Brandon Muffley. Data-rich but model-resistant: an evaluation of data-limited methods to manage fisheries with failed age-based stock assessments. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80(1):27–42, January 2023. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2022-0045.

Crowley:2023:DEA

[506] Samantha E. Crowley, Ian R. Bradbury, Amber M. Messmer, Steven J. Duffy, Christopher C. Parrish, Shahinur S. Islam, and Ian A. Fleming. Differences in energy acquisition and storage of farm, wild, and hybrid Atlantic salmon competing in the wild. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80(1):43–56, January 2023. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/doi/

full/10.1139/cjfas-2021-0326.

Tsai:2023:PND

[507] Cheng-Han Tsai, Stephan B. Munch, Michelle D. Masi, and Adam G. Pollack. Predicting nonlinear dynamics of short-lived penaeid shrimp species in the Gulf of Mexico. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80 (1):57–68, January 2023. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2022-0029.

Drago:2023:SIV

[508] Massimiliano Drago, Celia Llorach, Unai Ormazabal Santa Cruz, Lisette Zenteno-Devaud, Lorena Rebolledo, Diego Rita, Manel Gazo, and Luis Cardona. Seasonal and interannual variability in the diet of young and subadult male Antarctic fur seals as revealed by stable C and N isotope ratios. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80(1):69–84, January 2023. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2022-0145.

Quennessen:2023:ATD

[509] V. I. Quennessen, E. A. Babcock, and J. W. White. Accounting for transient dynamics could improve the use of marine protected areas as a reference point for fisheries management. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80(1):85–104, January 2023. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/doi/

full/10.1139/cjfas-2022-0125. See correction [675].

Shaughnessy:2023:JSL

[510] Ciaran A. Shaughnessy and Stephen D. McCormick. Juvenile sea lamprey (Petromyzon marinus) have a wide window of elevated salinity tolerance that is eventually limited during springtime warming. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80(1):105–114, January 2023. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2022-0097.

Siskey:2023:EIC

[511] Matthew R. Siskey, André E. Punt, Peter-John F. Hulson, Meaghan D. Bryan, James N. Ianelli, and James T. Thorson. The estimated impact of changes to otolith field-sampling and ageing effort on stock assessment inputs, outputs, and catch advice. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80(1):115–131, January 2023. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/doi/

full/10.1139/cjfas-2022-0050.

Mokdad:2023:NBE

[512] A. I. Mokdad, M. Elsheikh, O. M. Sulja, and T. E. Pitcher. Neuromorphological and behavioural effects of early developmental exposure to alarm cue on captive-reared Atlantic salmon (Salmo salar). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80(1):132–141, January 2023. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2022-0100.

Taguchi:2023:PGS

[513] Mioko Taguchi, Mutsuo Goto, Koji Matsuoka, Ralph Tiedemann, and Luis A. Pastene. Population genetic structure of Bryde’s whales (Balaenoptera brydei) on the central and western North Pacific feeding grounds. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80(1):142–155, January 2023. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/cjfas-
2022-0005.

Mordecai:2023:DPA

[514] Gideon Mordecai, Kurt Beardslee, Bonny Glambeck, Neil Frazer, Richard Routledge, and Alexandra Morton. Detection and phylogenetic assessment of PRV-1 via sampling of biological materials released from salmon farms in British Columbia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80 (1):156–169, January 2023. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2022-0019.

Rubenstein:2023:AAS

[515] Sarah R. Rubenstein, Erin Peterson, Paul Christman, and Joseph D. Zydlewski. Adult Atlantic salmon (Salmo salar) delayed below dams rapidly deplete energy stores. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80(1): 170–182, January 2023. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2022-0008.

He:2023:LAC

[516] Ji X. He, Andrew E. Honsey, David F. Staples, James R. Bence, and Tracy Claramunt. Longitudinal analyses of catch-at-age data for reconstructing year-class strength, with an application to lake trout (Salvelinus namaycush) in the main basin of Lake Huron. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80(1):183–194, January 2023. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/doi/

full/10.1139/cjfas-2022-0140.

Thompson:2023:ITL

[517] Patrick L. Thompson, Sean C. Anderson, Jessica Nephin, Carolyn K. Robb, Beatrice Proudfoot, Ashley E. Park, Dana R. Haggarty, and Emily M. Rubidge. Integrating trawl and longline surveys across British Columbia improves groundfish distribution predictions. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80(1):195–210, January 2023. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2022-0108.

Hamelin:2023:LWT

[518] Kayla M. Hamelin, Jeffrey A. Hutchings, and Megan Bailey. Look who’s talking: contributions to evidence-based decision making for commercial fisheries in Atlantic Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80 (2):211–228, February 2023. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2022-0025.

Clark:2023:UCS

[519] Christopher Clark, Robert Al-Chokhachy, and Kai Ross. Using continuous surveys to evaluate precision and bias of inferences from design-based reach-scale sampling of stream habitat. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80(2):229–242, February 2023. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/doi/

full/10.1139/cjfas-2022-0103.

Bellmore:2023:MCS

[520] J. Ryan Bellmore, Christopher J. Sergeant, Rebecca A. Bellmore, Jeffrey A. Falke, and Jason B. Fellman. Modeling Coho salmon (Oncorhynchus kisutch) population response to streamflow and water temperature extremes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80(2): 243–260, February 2023. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2022-0129.

Sturch:2023:HGS

[521] W. H. Sturch and C. C. D’Aloia. Hierarchical genetic structure in a directdeveloping whelk (Buccinum undatum) throughout the western North Atlantic. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80(2):261–272, February 2023. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/cjfas-
2022-0138.

Mahlum:2023:CHS

[522] Shad Mahlum, Kelsey Vitense, Hayley Corson-Dosch, Lindsay Platt, Jordan S. Read, Patrick J. Schmalz, Melissa Treml, and Gretchen J. A. Hansen. Connecting habitat to species abundance: the role of light and temperature on the abundance of walleye in lakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80(2):273–286, February 2023. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2022-0109.

Pfaff:2023:CAP

[523] Peter J. Pfaff, Kristen J. Hase, and Keith B. Gido. Community assembly of prairie farm ponds: build it and they will come, stock it and they won’t. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80(2):287–297, February 2023. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2022-0075.

Larson:2023:EAB

[524] Courtney E. Larson, Patrick Engelken, Deborah G. McCullough, and M. Eric Benbow. Emerald ash borer invasion of riparian forests alters organic matter and bacterial subsidies to south Michigan headwater streams. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80(2):298–312, February 2023. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2022-0127.

Lackmann:2023:BHB

[525] Alec R. Lackmann, Jeff Sereda, Mike Pollock, Reid Bryshun, Michelle Chupik, Katlin McCallum, James Villeneuve, Ewelina S. Bielak-Lackmann, and Mark E. Clark. Bet-hedging bigmouth buffalo (Ictiobus cyprinellus) recruit episodically over a 127-year timeframe in Saskatchewan. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80(2):313–329, February 2023. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533

(electronic). URL https://cdnsciencepub.com/doi/full/10.1139/
cjfas-2022-0122.

Lowen:2023:REE

[526] J. B. Lowen and C. DiBacco. Range expansion and establishment of a non-indigenous tunicate (Diplosoma listerianum) in thermal refugia is mediated by environmental variability in changing coastal environments. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80(2):330–345, February 2023. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2022-0082.

Neuswanger:2023:TWG

[527] Jason R. Neuswanger, Amanda E. Rosenberger, Mark S. Wipfli, and Nicholas F. Hughes. Territories within groups: the dynamic competition of drift-feeding juvenile Chinook salmon in 3-dimensional space. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80(2):346–359, February 2023. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2022-0112.

Reid:2023:STG

[528] Kerry Reid, Jennifer A. Hoey, Benjamin I. Gahagan, Bradley P. Schondelmeier, Daniel J. Hasselman, Alison A. Bowden, Michael P. Armstrong, John Carlos Garza, and Eric P. Palkovacs. Spatial and temporal genetic stock composition of river herring bycatch in southern New England Atlantic herring and mackerel fisheries. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80(2):360–374, February 2023. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/doi/

full/10.1139/cjfas-2022-0144.

Nuetzel:2023:IPN

[529] Hayley M. Nuetzel, Peter F. Galbreath, Benjamin A. Staton, Carrie A. Crump, Leslie M. Naylor, and Gene E. Shippentower. Improved productivity of naturalized spring Chinook salmon following reintroduction from a hatchery stock in Lookingglass Creek, Oregon. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80(2):375–392, February 2023. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/cjfas-2022-0114.

Roberts:2023:TSP

[530] Hayden C. Roberts, Matthew R. Acre, Maximiliaan P. A. Claus, Florian J. Kappen, Kirk O. Winemiller, Daniel J. Daugherty, and Joshuah S. Perkin. Tributary streams provide migratory fish with access to floodplain habitats in a regulated river: evidence from alligator gar, Atractosteus spatula. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80(2):393–407, February 2023. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2022-0076.

Steele:2023:DAG

[531] Reid W. Steele, M. Aaron MacNeil, and Sandie Hankewich. Direct assessment of giant red sea cucumber (Apostichopus californicus) sustainability through experimental fisheries. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80(2):408–419, February 2023. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/doi/

full/10.1139/cjfas-2022-0024.

Falcy:2023:IHE

[532] Matthew R. Falcy. Inferring hatchery effects using spawner-recruit data: comment on Courter et al. (2022). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80(2):420–421, February 2023. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/doi/

full/10.1139/cjfas-2022-0158. See [478] and response [533].

Courter:2023:RCC

[533] Ian I. Courter, Tom Chance, Ryan Gerstenberger, Mark Roes, Sean Gibbs, and Adrian Spidle. Response to Comment on Courter et al. (2022). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80(2):422–423, February 2023. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/cjfas-

2022-0234. See [478, 532].

Korman:2023:DRC

[534] Josh Korman, Bridget R. Deemer, Charles B. Yackulic, Theodore A. Kennedy, and Mariah Giardina. Drought-related changes in water quality surpass effects of experimental flows on trout growth downstream of Lake Powell reservoir. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80(3):424–438, March 2023. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/cjfas-
2022-0142.

Fonken:2023:BSF

[535] Dale R. Fonken, Mary M. Conner, Timothy E. Walsworth, and Paul D. Thompson. Benefits of stocking fewer but larger individuals with implications for native fish recovery. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80 (3):439–450, March 2023. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2022-0071.

Dunn:2023:IMA

[536] Muriel Dunn, Geir Pedersen, Sünnje L. Basedow, Malin Daase, Stig Falk-Petersen, Loïc Bachelot, Lionel Camus, and Maxime Geoffroy. Inverse method applied to autonomous broadband hydroacoustic survey detects higher densities of zooplankton in near-surface aggregations than vesselbased net survey. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80(3):451–467, March 2023. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/cjfas-
2022-0105.

Watson:2023:SCA

[537] Joe Watson, Andrew M. Edwards, and Marie Auger-Méthé. A statistical censoring approach accounts for hook competition in abundance indices from longline surveys. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80 (3):468–486, March 2023. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2022-0159.

Becu:2023:IST

[538] Mariella H. J. Becu, Tracy A. Michalski, and John S. Richardson. The influence of stream tributaries to lakes and upstream forest harvest on benthic, lake delta communities. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80 (3):487–502, March 2023. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2022-0134.

Swartz:2023:WWO

[539] Allison Swartz and Dana Warren. Wildfire in western Oregon increases stream temperatures, benthic biofilms, and juvenile coastal cutthroat trout size and densities with mixed effects on adult trout and coastal giant salamanders. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80(3):503–516, March 2023. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/cjfas-
2022-0053.

Goetz:2023:NNG

[540] Frederick Goetz, Shawn Sitar, Michael Seider, and Andrew Jasonowicz. Nature or nurture: a genetic basis for the behavioral selection of depth in siscowet and lean lake char (Salvelinus namaycush) ecomorphs. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80(3):517–532, March 2023. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2022-0104.

Charsley:2023:CPT

[541] Anthony R. Charsley, Nokuthaba Sibanda, Simon D. Hoyle, and Shannan K. Crow. Comparing the performance of three common species distribution modelling frameworks for freshwater environments through application to eel species in New Zealand. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80(3):533–548, March 2023. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/doi/

full/10.1139/cjfas-2022-0212.

Honsey:2023:GSB

[542] Andrew E. Honsey, Andrew L. Rypel, and Paul A. Venturelli. Guidance for selecting base temperatures when using degree-days in fish growth analyses. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80(3):549–562, March 2023. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2022-0197.

Marjadi:2023:EFR

[543] Meghna N. Marjadi, Allison H. Roy, Matthew T. Devine, Benjamin I. Gahagan, Adrian Jordaan, Julianne Rosset, and Andrew R. Whiteley. Effects of freshwater residence time on reproductive success in anadromous alewife (Alosa pseudoharengus): climate change implications. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80(3):563–576, March 2023. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2022-0086.

Persson:2023:ICL

[544] Lo Persson, Astrid Raunsgard, Eva B. Thorstad, Gunnel Østborg, Kurt Urdal, Harald Sægrov, Ola Ugedal, Kjetil Hindar, Sten Karlsson, Peder Fiske, and Geir H. Bolstad. Iteroparity and its contribution to life-history variation in Atlantic salmon. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80 (3):577–592, March 2023. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2022-0126.

Dean:2023:LTU

[545] Micah J. Dean, William S. Hoffman, Nicholas C. Buchan, Steven B. Scyphers, and Jonathan H. Grabowski. Lost in translation: understanding divergent perspectives on a depleted fish stock. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80(3):593–613, March 2023. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/cjfas-2022-0090.

Healy:2023:IAT

[546] Brian D. Healy, Charles B. Yackulic, and Robert C. Schelly. Impeding access to tributary spawning habitat and releasing experimental fall-timed floods increase brown trout immigration into a dam’s tailwater. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80(3):614–627, March 2023. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2022-0231.

Kluke:2023:SPE

[547] Calvin Kluke, Gretchen L. Lescord, Thomas A. Johnston, Brian W. Kielstra, Alan Lock, Satyendra Bhavsar, and John M. Gunn. Spatial patterns and environmental factors related to arsenic bioaccumulation in boreal freshwater fish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80(3):628–641, March 2023. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/cjfas-

2022-0106. See correction [565].

Chamberlin:2023:CAA

[548] Derek W. Chamberlin, Virginia R. Shervette, Darrell S. Kaufman, Jordon E. Bright, and William F. Patterson III. Can amino acid racemization be utilized for fish age validation? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80(3):642–647, March 2023. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/doi/

full/10.1139/cjfas-2022-0161.

Buckner:2023:IBL

[549] Jack H. Buckner, William H. Satterthwaite, Benjamin W. Nelson, and Eric J. Ward. Interactions between life history and the environment on changing growth rates of Chinook salmon. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80(4):648–662, April 2023. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/cjfas-2022-0116.

Calder:2023:DAH

[550] Ryan S. D. Calder, Jenni L. McDermid, and Stephanie A. Boudreau. Drivers of Atlantic herring decline and evidence basis for fisheries closures and rebuilding plans. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80(4):663–675, April 2023. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/cjfas-
2022-0169.

Powers:2023:DER

[551] Sean P. Powers, J. Marcus Drymon, Crystal L. Hightower, Liese M. Carleton, and John M. Hoenig. Direct estimates of reef fish abundance across an artificial reef network. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80 (4):676–689, April 2023. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2022-0094.

Nilsen:2023:ASW

[552] Cecilie I. Nilsen, Knut Wiik Vollset, Gaute Velle, Bjørn Torgeir Barlaup, Eirik Straume Normann, Elisabeth Stöger, and Robert James Lennox. Atlantic salmon of wild and hatchery origin have different migration patterns. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80(4):690–699, April 2023. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2022-0120.

James:2023:STH

[553] Samantha E. James, Beau Doherty, Sean P. Cox, Isobel A. Pearsall, and Brian Riddell. Size and timing of hatchery releases influence juvenileto-adult survival rates of British Columbia Chinook (Oncorhynchus tshawytscha) and Coho (Oncorhynchus kisutch) salmon. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80(4):700–718, April 2023. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2022-0121.

Masanovic:2023:EDL

[554] Marina Mašanović, Bent Herrmann, and Jure Brčić. Escape, discard, and landing probability of Nephrops norvegicus in the Mediterranean Sea creel fishery. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80(4):719–731, April 2023. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/cjfas-
2022-0190.

Pascoe:2023:EEB

[555] Sean Pascoe, André E. Punt, Trevor Hutton, Paul Burch, Pia Bessell-Browne, and L. Richard Little. Estimating economic-based target reference points for key species in multi-species multi-métier fisheries. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80(4):732–746, April 2023. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2022-0191.

Dayton:2023:EDM

[556] Alexa M. Dayton and Kanae Tokunaga. Economic diversity of Maine’s American lobster fishery. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80 (4):747–760, April 2023. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2022-0096.

Gimenez:2023:TPV

[557] Joan Giménez, Marta Albo-Puigserver, Raúl Laiz-Carrión, Elena Lloret-Lloret, José María Bellido, and Marta Coll. Trophic position variability of European sardine by compound-specific stable isotope analyses. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80(5):761–770, May 2023. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2022-0192.

Fontaine:2023:AIM

[558] Anthony Fontaine, Mackenzie Anne Clifford Martyniuk, Camille Garnier, and Patrice Couture. Assessing the impacts of mining activities on fish health in Northern Québec. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80 (5):771–794, May 2023. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2022-0200.

Dolan:2023:ASA

[559] Tara E. Dolan, Eric P. Palkovacs, Tanya L. Rogers, and Stephan B. Munch. Age structure augments the predictive power of time series for fisheries and conservation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80 (5):795–807, May 2023. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2022-0219.

Alglave:2023:IMF

[560] Baptiste Alglave, Youen Vermard, Etienne Rivot, Marie-Pierre Etienne, and Mathieu Woillez. Identifying mature fish aggregation areas during spawning season by combining catch declarations and scientific survey data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80(5):808–824, May 2023. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2022-0110.

Baldock:2023:RSM

[561] Jeffrey R. Baldock, Robert Al-Chokhachy, Timothy E. Walsworth, and Annika Walters. Redd superimposition mediates the accuracy, precision, and significance of redd counts for cutthroat trout. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80(5):825–839, May 2023. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2022-0267.

Togaki:2023:OSH

[562] Daisuke Togaki, Ayaka Sunohara, and Mikio Inoue. Overlap in spawning habitat characteristics between two salmonids in relation to stream size: redd superimposition hypothesis on longitudinal species replacement. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80(5):840–850, May 2023. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2022-0237.

Cortes:2023:DAR

[563] Federico Cortés, Jorge H. Colonello, Marina Sammarone, Anabela Zavatteri, and Natalia M. Hozbor. Demographic analyses reveal differential biological vulnerability in four Southwestern Atlantic skates. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80(5):851–865, May 2023. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2021-0340.

Melli:2023:EPT

[564] Valentina Melli, Bent Herrmann, Rikke P. Frandsen, Tiago Veiga-Malta, and Jordan P. Feekings. Escape panels in trawls: does placement matter when every individual contacting the panel can escape? Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80(5):866–891, May 2023. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2022-0205.

Kluke:2023:CSP

[565] Calvin Kluke, Gretchen L. Lescord, Thomas A. Johnston, Brian W. Kielstra, Alan Lock, Satyendra P. Bhavsar, and John M. Gunn. Correction: Spatial patterns and environmental factors related to arsenic bioaccumulation in boreal freshwater fish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80(5):892, May 2023. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2023-0062. See [547].

Poulton:2023:ODS

[566] Anna J. Poulton, Suresh A. Sethi, Stephen P. Ellner, and T. Scott Smeltz. Optimal dynamic spatial closures can improve fishery yield and reduce fishing-induced habitat damage. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80(6):893–912, June 2023. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/doi/

full/10.1139/cjfas-2022-0198.

Simmatis:2023:UMP

[567] Brigitte Simmatis, Courtney Lanigan, Breanna Lambert, Carsten Meyer-Jacob, Neal Michelutti, Andrew M. Paterson, and John P. Smol. Using a multi-proxy paleolimnological approach to track the effects of multiple stressors on a naturally acidic lake in Killarney Provincial Park (Ontario, Canada). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80(6):913–923, June 2023. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/cjfas-
2022-0225.

Carvalho:2023:RMM

[568] Paul G. Carvalho, William H. Satterthwaite, Michael R. O’Farrell, Cameron Speir, and Eric P. Palkovacs. Role of maturation and mortality in portfolio effects and climate resilience. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80(6):924–941, June 2023. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/doi/

full/10.1139/cjfas-2022-0171.

Kelly:2023:ENN

[569] Brett B. Kelly, Michael J. Siepker, and Michael J. Weber. Effects of non-native Salmo trutta and multiscale habitat factors on native fishes in the driftless area. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80(6):942–957, June 2023. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/cjfas-
2022-0213.

Crossman:2023:COE

[570] James A. Crossman, Josh Korman, Jason G. McLellan, Matthew D. Howell, and Andy L. Miller. Competition overwhelms environment and genetic effects on growth rates of endangered white sturgeon from a conservation aquaculture program. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80(6):958–977, June 2023. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/cjfas-
2022-0113.

Kindree:2023:CIB

[571] Meagan M. Kindree, Nicholas E. Jones, and Nicholas E. Mandrak. Competitive interactions between invasive round goby and native white sucker in streams. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80(6):978–991, June 2023. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/cjfas-
2022-0022.

Hallbert:2023:ICI

[572] Tyson B. Hallbert and Ernest R. Keeley. Instream complexity increases habitat quality and growth for cutthroat trout in headwater streams. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80(6):992–1005, June 2023. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2022-0189.

Tolimieri:2023:SLI

[573] N. Tolimieri and M. A. Haltuch. Sea-level index of recruitment variability improves assessment model performance for sablefish Anoplopoma fimbria. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80(6):1006–1016, June 2023. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2022-0238.

Nishijima:2023:MPD

[574] Shota Nishijima, Shigenori Suzuki, Ryo Fukuta, and Makoto Okada. Modeling pulse dynamics of juvenile fish enables the short-term forecasting of population dynamics in Japanese pufferfish: a latent variable approach and hindcasting. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80(6):1017–1031, June 2023. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/cjfas-
2022-0017.

Czaja:2023:ULR

[575] Raymond Czaja, Jr., Daniel Hennen, Robert Cerrato, Kamazima Lwiza, Emmanuelle Pales-Espinosa, Jennifer O’Dwyer, and Bassem Allam. Using LASSO regularization to project recruitment under CMIP6 climate scenarios in a coastal fishery with spatial oceanographic gradients. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80(6):1032–1046, June 2023. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2022-0091.

Patterson:2023:ABR

[576] William F. Patterson III and Derek W. Chamberlin. Application of the bomb radiocarbon chronometer with eye lens core Δ ^{14}C for age validation in deepwater reef fishes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80 (6):1047–1052, June 2023. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2023-0003.

Cunningham:2023:FIS

[577] Dylan S. Cunningham, Douglas C. Braun, Jonathan W. Moore, and Amanda M. Martens. Forestry influences on salmonid habitat in the North Thompson River watershed, British Columbia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80(7):1053–1070, July 2023. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/cjfas-2022-0255. See comment [672] and response

[673].

Forrest:2023:EBA

[578] Robyn E. Forrest, Allen R. Kronlund, Jaclyn S. Cleary, and Matthew H. Grinnell. An evidence-based approach for selecting a limit reference point for Pacific herring (Clupea pallasii) stocks in British Columbia, Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80(7):1071–1083, July 2023. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2022-0168.

Bourret:2023:QGD

[579] Audrey Bourret, Andrew Smith, Elisabeth Van Beveren, Stéphane Plourde, Kiersten L. Curti, Teunis Jansen, David E. Richardson, Martin Castonguay, Naiara Rodriguez-Ezpeleta, and Geneviève J. Parent. Quantifying genetic differentiation and population assignment between two contingents of Atlantic mackerel (Scomber scombrus) in the Northwest Atlantic. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80(7):1084–1097, July 2023. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/cjfas-
2022-0232.

Freshwater:2023:CLT

[580] Cameron Freshwater, Will Duguid, Francis Juanes, and Skip McKinnell. A century-long time series reveals large declines and greater synchrony in Nass River sockeye salmon size-at-age. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80(7):1098–1109, July 2023. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/doi/

full/10.1139/cjfas-2022-0259.

Davies:2023:TAF

[581] Peter Davies, J. Robert Britton, Theodore Castro-Santos, Charles Crundwell, Jamie R. Dodd, Andrew D. Nunn, Randolph Velterop, and Jonathan D. Bolland. Tracking anadromous fish over successive freshwater migrations reveals the influence of tagging effect, previous success, and abiotic factors on upstream passage over barriers. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80(7):1110–1125, July 2023. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2022-0196.

Fader:2023:NFL

[582] Joseph E. Fader, Jamie A. Marchetti, Robert S. Schick, and Andrew J. Read. No free lunch: estimating the biomass and ex-vessel value of target catch lost to depredation by odontocetes in the Hawai’i longline tuna fishery. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80(7):1126–1141, July 2023. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2022-0156.

Ozyurt:2023:CBB

[583] Caner Enver Ozyurt, Sefik Surhan Tabakoglu, Volkan Barıs Kiyaga, and Gulsun Ozyurt. Characterization of bluefish baits (keeled mullet, leaping mullet, and sardine) in regard to biochemical and physical properties: bait preference of bluefish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80 (7):1142–1155, July 2023. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2022-0135.

Cook:2023:PCI

[584] Jenna Cook, Charlie J. G. Loewen, Tamika L. Nagao, Mark D. Graham, and Rolf D. Vinebrooke. Phytoplankton communities as indicators of environmental change in the Canadian Rockies. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80(7):1156–1169, July 2023. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/cjfas-2022-0256.

Bernos:2023:EDE

[585] Thaïs A. Bernos, Matthew C. Yates, Margaret F. Docker, Amy Fitzgerald, Robert Hanner, Daniel Heath, Arshad Imrit, John Livernois, Erika Myler, Keta Patel, Sapna Sharma, Robert Young, and Nicholas E. Mandrak. Environmental DNA (eDNA) applications in freshwater fisheries management and conservation in Canada: overview of current challenges and opportunities. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80(7):1170–1186, July 2023. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/cjfas-
2022-0162.

Isaak:2023:CWH

[586] Daniel J. Isaak and Michael K. Young. Cold-water habitats, climate refugia, and their utility for conserving salmonid fishes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80(7):1187–1206, July 2023. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2022-0302.

Mullowney:2023:MIP

[587] Darrell R. J. Mullowney and Krista D. Baker. Multi-indicator precautionary approach frameworks for crustacean fisheries. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80(7):1207–1220, July 2023. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2022-0260.

Ou:2023:ACP

[588] Liguo Ou, Bilin Liu, Xinjun Chen, Qi He, Weiguo Qian, Wenlong Li, Leilei Zou, Yixi Shi, and Qinglian Hou. Automatic classification of the phenotype textures of three Thunnus species based on the machine learning SVM algorithm. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80(8):1221–1236, August 2023. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/cjfas-
2022-0270.

Shamaskin:2023:ERL

[589] Andrew C. Shamaskin, Michael E. Colvin, and L. E. Miranda. Evaluating regional length limits in freshwater fisheries. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80(8):1237–1247, August 2023. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/cjfas-2022-0179.

Klasios:2023:MSW

[590] Natasha Klasios and Michelle Tseng. Microplastics in subsurface water and zooplankton from eight lakes in British Columbia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80(8):1248–1267, August 2023. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2022-0293.

Su:2023:EMP

[591] Zhenming Su. Evaluation of management performance of a new statespace model for pink salmon (Oncorhynchus gorbuscha) stock–recruitment analysis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80(8):1268–1288, August 2023. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/cjfas-
2022-0262.

Kai:2023:MAD

[592] Mikihiko Kai, Hiroki Yokoi, and Yuki Fujinami. Modeling of agedependent natural mortality rates for long-lived fishes based on the Richards model family. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80(8): 1289–1300, August 2023. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2022-0257.

Bassi:2023:CNS

[593] Laélien Bassi, Réjean Tremblay, Anne-Laure Ferchaud, Louis Bernatchez, Dominique Robert, and Pascal Sirois. Connectivity and natal sources of Greenland halibut in the Gulf of St. Lawrence inferred from otolith chemistry. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80(8):1301–1312, August 2023. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/cjfas-
2022-0081.

Elmer:2023:CIA

[594] Laura K. Elmer, Arthur L. Bass, Stephen D. Johnston, Karia H. Kaukinen, Lisa A. Kelly, Shaorong Li, Amy K. Teffer, Kristi M. Miller, Steven J. Cooke, and Scott G. Hinch. Changes in infectious agent profiles and host gene expression during spawning migrations of adult sockeye salmon (Oncorhynchus nerka). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80(8):1313–1334, August 2023. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/cjfas-
2022-0132.

Peterson:2023:TCS

[595] Cheston T. Peterson and R. Dean Grubbs. Temporal community structure and seasonal climatic migration of coastal sharks and large teleost fishes in the northeast Gulf of Mexico. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80 (8):1335–1350, August 2023. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2022-0124.

Kallo:2023:FAM

[596] Kristi Källo, Kim Birnie-Gauvin, Henrik Baktoft, and Kim Aarestrup. On the factors affecting migration and straying in brown trout (Salmo trutta). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80(8):1351–1364, August 2023. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2022-0298.

Austin:2023:WSW

[597] Catherine S. Austin, Christian E. Torgersen, and Thomas P. Quinn. Who spawns where? Temperature, elevation, and discharge differentially affect the distribution of breeding by six Pacific salmonids within a large river basin. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80(8):1365–1384, August 2023. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/cjfas-
2022-0252.

Kornijow:2023:LAC

[598] Ryszard Kornijów. Lake aging concept revisited: alternative comprehensive approach to lake ontogeny. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80 (8):1385–1393, August 2023. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2023-0027.

Vainikka:2023:TGC

[599] Anssi Vainikka, Chris K. Elvidge, Jenni M. Prokkola, Alexandre Lemopoulos, Matti Vornanen, Laura S. Härkönen, Nico Alioravainen, and Pekka Hyvärinen. Two-generation common-garden experiment reveals a strong genetic contribution to migration tendency in brown trout (Salmo trutta). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80(9):1394– 1409, September 2023. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2022-0304.

Leahy:2023:MAA

[600] Susannah M. Leahy, Dean R. Jerry, Brett B. C. Wedding, Julie B. Robins, Carole L. Wright, Aleksey Sadekov, Stephen Boyle, David B. Jones, Samuel M. Williams, Steve Grauf, Luke Pavich, Mark McLennan, Michelle J. Sellin, Julie A. Goldsbury, and Richard J. Saunders. Multimethod approach to advance provenance determination of fish in stocked systems. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80(9):1410– 1424, September 2023. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2022-0292.

Munyandorero:2023:AES

[601] Joseph Munyandorero. Analytical expressions of sustainable benchmarks generic for all two-stage-structured models of exploited fish populations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80(9):1425–1435, September 2023. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2022-0283.

Howarth:2023:MCF

[602] Andrew Howarth, Vivian M. Nguyen, and Steven J. Cooke. Managing Canadian freshwater fisheries: persistent challenges and emerging opportunities. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80(9):1436– 1455, September 2023. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2023-0011.

Hansen:2023:LEP

[603] Lindsay E. Hansen, Charles B. Yackulic, Brett G. Dickson, Bridget R. Deemer, and Rebecca J. Best. Linking ecosystem processes to consumer growth rates: gross primary productivity as a driver of freshwater fish somatic growth in a resource-limited river. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80(9):1456–1469, September 2023. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2022-0229.

Sanborn:2023:DHR

[604] Delaney Sanborn, Alexis Base, Lysel Garavelli, Ranjoy Barua, Jiarong Hong, and Aditya R. Nayak. Digital holography for real-time noninvasive monitoring of larval fish at power plant intakes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80(9):1470–1481, September 2023. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2023-0058.

Benson:2023:FFA

[605] Irina M. Benson, Thomas E. Helser, Giovanni Marchetti, and Beverly K. Barnett. The future of fish age estimation: deep machine learning coupled with Fourier transform near-infrared spectroscopy of otoliths. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80(9):1482–1494, September 2023. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2023-0045.

Cresson:2023:FLD

[606] Pierre Cresson, Odile Bourgau, Remy Cordier, Clémence Couvreur, Manuel Rouquette, and Mélanie Gay. Fish length, diet, and depth drive Anisakis levels in a zooplankton-feeding fish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80(9):1495–1508, September 2023. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2022-0272.

Breivik:2023:DSR

[607] Olav Nikolai Breivik, Magne Aldrin, Edvin Fuglebakk, and Anders Nielsen. Detecting significant retrospective patterns in state space fish stock assessment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80(9):1509– 1518, September 2023. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2022-0250.

Yagi:2023:CFF

[608] Tatsunori Yagi and Takashi Yamakawa. A comprehensive functional form of the optimal harvest control rule for multiple fishery management objectives. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80(9):1519– 1532, September 2023. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2022-0195.

McLaren:2023:SDR

[609] John S. McLaren, Robert W. Van Kirk, Phaedra Budy, and Soren Brothers. The scale-dependent role of submerged macrophytes as drift-feeding lotic fish habitat. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80(9):1533– 1546, September 2023. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2022-0182.

Howell:2023:LTR

[610] Bradley E. Howell, Giulio Navarroli, Eric J. Mullen, Steven J. Cooke, and Caleb T. Hasler. Lake trout reflex impairment and physiological status following ice-angling. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80(10): 1547–1561, October 2023. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2023-0037.

Shuter:2023:IHT

[611] Brian J. Shuter, Scott W. Milne, Lee E. Hrenchuk, Derrick T. deKerckhove, and Michael D. Rennie. Integrating hydroacoustic and telemetric surveys to estimate fish abundance: a new approach to an old problem. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80(10):1562–1578, October 2023.

CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2022-0183.

Norstog:2023:LPS

[612] Jessica L. Norstog and Stephen D. McCormick. Landlocked populations have small but detectable differences in ionoregulatory physiology compared to anadromous sea lamprey, Petromyzon marinus. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80(10):1579–1594, October 2023. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2022-0242.

Gruss:2023:SVC

[613] Arnaud Grüss, James T. Thorson, Owen F. Anderson, Richard L. O’Driscoll, Madison Heller-Shipley, and Scott Goodman. Spatially varying catchability for integrating research survey data with other data sources: case studies involving observer samples, industry-cooperative surveys, and predators as samplers. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80(10): 1595–1615, October 2023. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2023-0051.

Burbank:2023:SFI

[614] Jacob Burbank, François Turcotte, François-Étienne Sylvain, and Nicolas Rolland. A statistical framework for identifying the relative importance of ecosystem processes and demographic factors on fish recruitment, with application to Atlantic herring (Clupea harengus) in the southern Gulf of St. Lawrence. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80(10):1616–1630, October 2023. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/cjfas-
2023-0081.

Kelley:2023:IYB

[615] Kayla Kelley, Eliza I. Gilbert, Casey A. Pennock, Mark C. McKinstry, Peter D. Mackinnon, Scott L. Durst, and Nathan R. Franssen. If you build it, will they pass? A systematic evaluation of fish passage efficiency for three large-bodied warm-water fishes. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80(10):1631–1643, October 2023. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/cjfas-2023-0030.

Wacker:2023:INN

[616] Sebastian Wacker, Geir H. Bolstad, Ola H. Diserud, Kjetil Hindar, and Sten Karlsson. Introgression of non-native mitochondrial haplotypes from farmed to wild Atlantic salmon. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80(10):1644–1652, October 2023. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/cjfas-2023-0044.

McIlwraith:2023:TMS

[617] Hayley K. McIlwraith, Eden K. Hataley, and Chelsea M. Rochman. Towards a management strategy for microplastic pollution in the Laurentian Great Lakes — monitoring (part 1). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80(10):1653–1668, October 2023. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/cjfas-2023-0022.

Hataley:2023:TMS

[618] Eden K. Hataley, Hayley K. McIlwraith, Dimple Roy, and Chelsea M. Rochman. Towards a management strategy for microplastic pollution in the Laurentian Great Lakes — ecological risk assessment and management (part 2). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80(10):1669–1678, October 2023. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/cjfas-

2023-0023. See correction [708].

Godwin:2023:SIE

[619] Sean C. Godwin, Andrew W. Bateman, Gideon Mordecai, Sean Jones, and Jeffrey A. Hutchings. Is scientific inquiry still incompatible with government information control? A quarter-century later. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80(10):1679–1695, October 2023. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2022-0286.

Falkegaard:2023:PAS

[620] Morten Falkegård, Robert J. Lennox, Eva B. Thorstad, Sigurd Einum, Peder Fiske, Øyvind A. Garmo, Åse H. Garseth, Helge Skoglund, Monica F. Solberg, Kjell R. Utne, Knut W. Vollset, Leif Asbjørn Vøllestad, Vidar Wennevik, and Torbjørn Forseth. Predation of Atlantic salmon across ontogenetic stages and impacts on populations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80(10):1696–1713, October 2023. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2023-0029.

He:2023:GAC

[621] Ji X. He, Charles P. Madenjian, and Todd C. Wills. A generalized application of the catch-curve regression with comparisons of adult mortality and year-class strength between hatchery-stocked and wild-reared lake trout in US waters of Lake Huron. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80(10): 1714–1722, October 2023. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2022-0275.

Freeman:2023:PNR

[622] Mary C. Freeman, Brett Albanese, Phillip M. Bumpers, Megan M. Hagler, Andrew J. Nagy, Byron J. Freeman, and Seth J. Wenger. Persistence of native riverine fishes downstream from two hydropower dams with contrasting operations. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80 (11):1723–1736, November 2023. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/doi/

full/10.1139/cjfas-2022-0297.

Goto:2023:AFA

[623] Akihiko Goto, Mari Kuroki, and Kentaro Morita. Active feeding of anadromous white-spotted char Salvelinus leucomaenis at the southern latitudinal rivers. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80(11):1737– 1747, November 2023. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2023-0028.

Benjamin:2023:CVD

[624] Joseph R. Benjamin, Jason B. Dunham, Kara J. Anlauf-Dunn, and Erika Eliason. Climate vulnerability for a desert fish: integrating hydrologic exposures, adaptive capacity, and growth potential. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80(11):1748–1758, November 2023. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2022-0280.

Karvonen:2023:ERE

[625] Anssi Karvonen, Ines Klemme, Ville Räihä, and Pekka Hyvärinen. Enriched rearing environment enhances fitness-related traits of salmonid fishes facing multiple biological interactions. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80(11):1759–1769, November 2023. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2023-0083.

Caballero:2023:FPL

[626] Smit Vasquez Caballero, Gil Sylvia, and Daniel S. Holland. Fishery participation and location choice model: the West Coast salmon troll commercial fishery. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80(11):1770– 1784, November 2023. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2023-0001.

Schacht:2023:ALU

[627] Joe R. Sánchez Schacht, Paul W. MacKeigan, Zofia E. Taranu, Yannick Huot, and Irene Gregory-Eaves. Agricultural land use and morphometry explain substantial variation in nutrient and ion concentrations in lakes across Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80(11): 1785–1797, November 2023. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2023-0109. See correction [632].

Hermann:2023:YRM

[628] Nathan T. Hermann, Lars J. Hammer, Nigel E. Hussey, Marianne Marcoux, Kevin J. Hedges, Ryan P. Walter, and Nathan B. Furey. Year-round monitoring of Arctic species of sculpin to identify residency and seasonality of movement behavior. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80 (11):1798–1812, November 2023. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/doi/

full/10.1139/cjfas-2023-0042.

Gallage:2023:IEL

[629] Kavishka S. Gallage, Alexander Van Nynatten, Nathan K. Lujan, Nathan R. Lovejoy, and Nicholas E. Mandrak. Identifying early life stages of Great Lakes fishes using a metabarcoding approach. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80(11):1813–1823, November 2023. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2023-0061.

McCleary:2023:CSS

[630] Richard J. McCleary and Ronald A. Ptolemy. Comment: Surprising salmonid response to water diversion at four run-of-river hydroelectric projects in British Columbia. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80 (11):1824–1830, November 2023. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/doi/

full/10.1139/cjfas-2022-0291. See [288] and reply [631].

Hocking:2023:RCS

[631] M. D. Hocking, S. G. Faulkner, E. Vogt, K. Akaoka, A. Harwood, T. Hatfield, and F. J. A. Lewis. Reply to “Comment: Surprising salmonid response to water diversion at four run-of-river hydroelectric projects in British Columbia”. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80(11): 1831–1840, November 2023. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2023-0153. See [288, 630].

SanchezSchacht:2023:CAL

[632] Joe R. Sánchez Schacht, Paul W. MacKeigan, Zofia E. Taranu, Yannick Huot, and Irene Gregory-Eaves. Correction: Agricultural land use and morphometry explain substantial variation in nutrient and ion concentrations in lakes across Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80 (11):1841–1843, November 2023. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/doi/

full/10.1139/cjfas-2023-0271. See [627].

Medinski:2023:CCC

[633] Nathan A. Medinski, Bryan M. Maitland, Timothy D. Jardine, D. Andrew R. Drake, and Mark S. Poesch. Correction: A catastrophic coal mine spill in the Athabasca River watershed induces isotopic niche shifts in stream biota including an endangered rainbow trout ecotype. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80(11):1844–1846, November 2023. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2023-0196. See [442].

Anonymous:2023:TTC

[634] Anonymous. Time to treat the climate and nature crisis as one indivisible global health emergency. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80(12), December 2023. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/doi/

full/10.1139/cjfas-2023-0324.

Paez:2023:CSL

[635] David J. Páez, Jacob L. Gregg, Ashley H. MacKenzie, Sophie A. Hall, and Paul K. Hershberger. Characteristics of a sea louse (Caligus clemensi) epizootic in wild Pacific herring (Clupea pallasii). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80(12):1847–1856, December 2023. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2023-0070.

Vaisvil:2023:NHS

[636] Alexander Vaisvil, Malte Willmes, Edward J. Enriquez, Zachary B. Klein, and Colleen A. Caldwell. A needle in a haystack: strontium isotopes (^{87}Sr/ 86Sr) in otoliths identify origin of largemouth bass from a large Southwest reservoir. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80(12):1857–1868, December 2023. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/cjfas-
2023-0148.

Nephin:2023:IDS

[637] Jessica Nephin, Patrick L. Thompson, Sean C. Anderson, Ashley E. Park, Christopher N. Rooper, Brendan Aulthouse, and Joe Watson. Integrating disparate survey data in species distribution models demonstrate the need for robust model evaluation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80 (12):1869–1889, December 2023. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/doi/

full/10.1139/cjfas-2022-0279.

Lucas:2023:EBS

[638] Jorrit Lucas, Albert Ros, Juergen Geist, and Alexander Brinker. Effects of behavioral strategies on catchability, bait selectivity, and hunting behavior in northern pike (Esox lucius). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80 (12):1890–1905, December 2023. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/doi/

full/10.1139/cjfas-2023-0004.

Wang:2023:SSA

[639] Zhen Wang, Hubert Keckeis, Fei Cheng, and Songguang Xie. Speciesspecific adaptations, individual condition, and life-history-induced dispersal strategies of fish larvae along a transverse gradient in a large river. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80(12):1906–1920, December 2023. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2023-0078.

Liao:2023:AWL

[640] Chuansong Liao, Jiacheng Wang, Sandra Bibiana Correa, Jixin Yu, Rui Yang, Jing Yuan, Wei Li, Shaowen Ye, Chuanbo Guo, and Jiashou Liu. Artificial water-level fluctuation modulates trophic niches of benthic fish assemblages in one of the world’s largest reservoirs. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80(12):1921–1934, December 2023. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2023-0114.

Askey:2023:CKS

[641] Paul J. Askey, Hillary G. M. Ward, Tyler Weir, and Kristen King. Comparison of known spawner abundance from fence counts to visual counts for simplified spawner estimation methods. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80(12):1935–1948, December 2023. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2023-0111.

Harnish:2023:FAP

[642] Ryan A. Harnish, Kenneth D. Ham, John R. Skalski, Richard L. Townsend, and Rebecca A. Buchanan. Factors affecting powerhouse passage of spring migrant smolts at federally operated hydroelectric dams of the Snake and Columbia rivers. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80(12):1949–1966, December 2023. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/cjfas-2022-0217.

McKelvey:2023:STP

[643] Veronica M. L. McKelvey and Nicholas E. Mandrak. Spatial and temporal patterns in the threats to at-risk freshwater fish species in Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80(12):1967–1983, December 2023. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2023-0170.

Routledge:2023:EGR

[644] Richard Routledge and Alexandra Morton. Effect of government removal of salmon farms on sea lice infection of juvenile wild salmon in the Discovery Islands. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 80(12): 1984–1989, December 2023. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2023-0039.

Anonymous:2024:NA

[645] Anonymous. Note of appreciation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 81(1), January 2024. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2023-0263.

Iacarella:2024:PFS

[646] Josephine C. Iacarella and J. Daniel Weller. Predicting favourable streams for anadromous salmon spawning and natal rearing under climate change. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 81(1):1–13, January 2024. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2023-0096. See cor-

rection [656].

OFarrell:2024:LRW

[647] Shay O’Farrell, Larry Perruso, James N. Sanchirico, and Iliana Chollett. Linking real world fisheries datasets for mapping of revenue from fishing grounds to dependent communities. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 81(1):14–27, January 2024. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/doi/

full/10.1139/cjfas-2023-0123.

Tong:2024:DCP

[648] Shanying Tong, Dong Xu, Hongjin Qiao, and Naihao Ye. Decreased calcification to photosynthesis ratio in coccolithophores under reduced O_{2} and elevated CO_{2} environment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 81 (1):28–37, January 2024. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2023-0073.

Mensinger:2024:DFP

[649] Matthew A. Mensinger, James P. Hawkes, Graham S. Goulette, Alessio Mortelliti, Erik J. Blomberg, and Joseph D. Zydlewski. Dams facilitate predation during Atlantic salmon (Salmo salar) smolt migration. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 81(1):38–51, January 2024. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2023-0175.

Sutela:2024:SER

[650] Tapio Sutela, Teppo Vehanen, Pekka Jounela, Jukka Aroviita, and Topi K. Lehtonen. Species–environment relationships, clusters, and thermal ranges of fish species inhabiting boreal rivers. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 81(1):52–62, January 2024. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/cjfas-2023-0016.

Hulson:2024:RSE

[651] Peter-John F. Hulson, Benjamin C. Williams, Meaghan D. Bryan, Jason Conner, and Matthew R. Siskey. Reductions in sampling effort for fishery-independent age and length composition: balancing sampling efficiency, data uncertainty, and workforce health. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 81(1):63–78, January 2024. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/cjfas-2023-0164.

Simonson:2024:HMP

[652] Martin A. Simonson, Michael J. Weber, Audrey L. McCombs, and Andrew R. Annear. Hierarchical model to predict common carp and bigmouth buffalo abundance from electrofishing data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 81(1):79–89, January 2024. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2023-0017.

vanDorst:2024:EHE

[653] Renee M. van Dorst, Anna Gårdmark, Kimmo K. Kahilainen, Leena Nurminen, Satu Estlander, Hannu Huuskonen, Mikko Olin, Martti Rask, and Magnus Huss. Ecosystem heating experiment reveals sex-specific growth responses in fish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 81 (1):90–96, January 2024. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2023-0034.

Murdoch:2024:MED

[654] A. Murdoch, B. M. Connors, N. W. R. Lapointe, J. Mills Flemming, S. J. Cooke, and C. Mantyka-Pringle. Multiple environmental drivers across life stages influence Yukon River Chinook salmon productivity. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 81(1):97–114, January 2024. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2022-0254.

Gerig:2024:TEJ

[655] Brandon S. Gerig, Shawn P. Sitar, Will F. Otte, Daniel L. Yule, Heidi K. Swanson, Charles R. Bronte, Dray Carl, and Joshua Blankenheim. Trophic ecology of juvenile lean and siscowet lake charr (Salvelinus namaycush) in Lake Superior: assessing for potential competition. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 81(1):115–128, January 2024. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2023-0102.

Iacarella:2024:CPF

[656] Josephine C. Iacarella and J. Daniel Weller. Correction: Predicting favourable streams for anadromous salmon spawning and natal rearing under climate change. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 81(1): 129–136, January 2024. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2023-0317. See [646].

DeSellas:2024:LWC

[657] Anna M. DeSellas, Andrew M. Paterson, Kathleen M. Rühland, and John P. Smol. Lake water chemistry and its relationship to shoreline residential development and natural landscape features in Algonquin Provincial Park, Ontario, Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 81 (2):137–153, February 2024. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2023-0103.

Malick:2024:SVE

[658] Michael J. Malick, Mary E. Hunsicker, Melissa A. Haltuch, Sandra L. Parker-Stetter, Kristin N. Marshall, John E. Pohl, Aaron M. Berger, Samantha A. Siedlecki, Stéphane Gauthier, and Albert J. Hermann. Spatially varying effects of the California Undercurrent on Pacific hake distribution. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 81(2):154–165, February 2024. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/cjfas-
2023-0202.

Vasiliadis:2024:ATE

[659] Michaelis Vasiliadis, Jennifer J. Freer, Martin A. Collins, and Alison C. Cleary. Assessing the trophic ecology of Southern Ocean Myctophidae: the added value of DNA metabarcoding. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 81(2):166–177, February 2024. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/doi/

full/10.1139/cjfas-2023-0079.

Brantschen:2024:CSE

[660] Jeanine Brantschen and Florian Altermatt. Contrasting strengths of eDNA and electrofishing compared to historic records for assessing fish community diversity and composition. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 81(2):178–189, February 2024. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/doi/

full/10.1139/cjfas-2023-0053.

Morissette:2024:EDR

[661] Olivier Morissette, Cristina Charette, Matthew J. S. Windle, Abraham Francis, Annick Drouin, Jesica Goldsmit, and Alison M. Derry. Environmental determinants of round goby invasion refuges at a river scale: implications for conservation of native biodiversity. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 81(2):190–201, February 2024. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/cjfas-2023-0171.

Papatheodoulou:2024:ESC

[662] Magdalene Papatheodoulou, Neil B. Metcalfe, and Shaun S. Killen. Effects of simulated catch-and-release angling of Atlantic salmon shortly before spawning on the viability and development of their offspring. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 81(2):202–211, February 2024. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2022-0306.

Madden:2024:BRU

[663] Jamie C. Madden, Luc LaRochelle, Declan Burton, Sascha C. Danylchuk, Andy J. Danylchuk, and Steven J. Cooke. Biologgers reveal unanticipated issues with descending angled walleye with barotrauma symptoms. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 81(2):212–222, February 2024. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2023-0141.

Bareille:2024:FFO

[664] G. Bareille, M. Vignon, A. Chappaz, A. Fontaine, H. Tabouret, F. Morat, J. Martin, J. C. Aymes, F. Daverat, C. Pécheyran, and O. Donard. Freshwater fish otoliths record signals from both water and physiological processes: new insights from Sr/Ca and Ba/Ca ratios. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 81(2):223–240, February 2024. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2022-0030.

McElroy:2024:AIF

[665] Katherine N. McElroy, Caitlin A. Stern, Thomas P. Quinn, and Ray Hilborn. Applying the ideal free distribution to the movement of a highly mobile gillnet fishery for Pacific salmon. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 81(3):241–252, March 2024. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/cjfas-2023-0134.

Arbeider:2024:EGR

[666] Michael Arbeider, Violaine Pemberton-Renaud, Emma E. Hodgson, and Jonathan W. Moore. The estuarine growth and residency of juvenile Pacific salmon in North America: a compilation of empirical data. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 81(3):253–267, March 2024. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2023-0225.

Kerametsidis:2024:CSE

[667] Georgios Kerametsidis, James Thorson, Vincent Rossi, Diego Álvarez-Berastegui, Cheryl Barnes, Gregoire Certain, Antonio Esteban, Encarnación García, Angélique Jadaud, Safo Piñeiro, Miguel Vivas, and Manuel Hidalgo. Cross-scale environmental impacts across persistent and dynamic aggregations within a complex population: implications for fisheries management. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 81(3):268–284, March 2024. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/cjfas-
2023-0120.

Thelen:2024:LCD

[668] Mary A. Thelen, David J. Gallagher, Payton E. Johnson, Sara M. Kangas, Dylan J. McNulty, Alexandra L. Morrison, Brian R. Herwig, David F. Staples, and Kyle D. Zimmer. Lake characteristics drive concordant trophic responses across ecosystems in three top predator fish species. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 81(3):285–296, March 2024. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2023-0240.

Jacobs:2024:LCM

[669] Gregory R. Jacobs, Russell F. Thurow, Charles E. Petrosky, Craig W. Osenberg, and Seth J. Wenger. Life-cycle modeling reveals high recovery potential of at-risk wild Chinook salmon via improved migrant survival. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 81(3):297–310, March 2024. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2023-0167.

Kikko:2024:AMN

[670] Takeshi Kikko, Takuya Sato, Minoru Kanaiwa, Daisuke Ishizaki, Kunihiko Kuwamura, Haruo Okamoto, Atsuhiko Ide, Morihito Nemoto, Shoichiro Takeoka, Yoshiaki Kai, Kouji Nakayama, and Yasuhiro Fujioka. Apparent migration and natal homing of a small minnow in a large ancient lake. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 81(3):311–325, March 2024. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2022-0207.

Lennox:2024:ETT

[671] Robert J. Lennox, Pedro Afonso, Kim Birnie-Gauvin, Lotte S. Dahlmo, Cecilie I. Nilsen, Robert Arlinghaus, Steven J. Cooke, Allan T. Souza, Ivan Jarić, Marie Prchalová, Milan Říha, Samuel Westrelin, William Twardek, Eneko Aspillaga, Sebastian Kraft, Marek Šmejkal, Henrik Baktoft, Tomas Brodin, Gustav Hellström, David Villegas-Ríos, Knut Wiik Vollset, Timo Adam, Lene K. Sortland, Michael G. Bertram, Marcelo Crossa, Emma F. Vogel, Natasha Gillies, and Jan Reubens. Electronic tagging and tracking aquatic animals to understand a world increasingly shaped by a changing climate and extreme weather events. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 81(3):326–339, March 2024. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/doi/

full/10.1139/cjfas-2023-0145.

Moore:2024:CFI

[672] R. D. (Dan) Moore. Comment: “Forestry influences on salmonid habitat in the North Thompson River watershed, British Columbia”. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 81(3):340–344, March 2024. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:

//cdnsciencepub.com/doi/full/10.1139/cjfas-2023-0259. See [577]

and response [673].

Cunningham:2024:FST

[673] Dylan S. Cunningham, Douglas C. Braun, Jonathan W. Moore, and Amanda M. Martens. Forestry and stream temperatures: a response to comment by R. D. Moore. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 81 (3):345–347, March 2024. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2023-0306. See [577, 672].

Gardner:2024:ENR

[674] Spencer T. Gardner and Tomas O. Höök. Efficacy of a novel reproductive tag to index spawn timing. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 81 (3):348–353, March 2024. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2023-0332.

Quennessen:2024:CAT

[675] V. I. Quennessen, E. A. Babcock, and J. W. White. Correction: Accounting for transient dynamics could improve the use of marine protected areas as a reference point for fisheries management. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 81(3):354, March 2024. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/doi/

full/10.1139/cjfas-2024-0024. See [509].

Huber:2024:SYD

[676] Eric R. Huber, Rachael E. Ryan, Rachel C. Johnson, Anna M. Sturrock, Robert A. Lusardi, and Stephanie M. Carlson. Seventy years of diminishing biocomplexity of California Central Valley hatchery steelhead, Oncorhynchus mykiss. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 81(4):355–367, April 2024. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/cjfas-
2023-0077.

Lerner:2024:SID

[677] Jacob E. Lerner and Brian P. V. Hunt. Stable isotopes delineate regional pelagic food web structure in British Columbia’s coastal ocean. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 81(4):368–386, April 2024. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2023-0057.

English:2024:DEM

[678] Greg English, Brent M. Wilson, Michael J. Lawrence, Matt Black, James P. Hawkes, David C. Hardie, Jason M. Daniels, Jonathan W. Carr, Claire Rycroft, Glenn T. Crossin, Fred G. Whoriskey, Cornelia E. den Heyer, Xavier Bordeleau, Christopher W. McKindsey, and Marc Trudel. Determining early marine survival and predation by endothermic predators on acoustically tagged Atlantic salmon (Salmo salar) post-smolts. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 81(4):387–402, April 2024. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2023-0206.

Haas:2024:ESE

[679] Taylor F. Haas, Travis O. Brenden, Zhiqun D. Deng, and C. Michael Wagner. Evaluation of survival estimates generated from tracking downstream migrating juvenile sea lamprey (Petromyzon marinus) with a miniature acoustic telemetry tag. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 81 (4):403–416, April 2024. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2023-0194.

Struthers:2024:SST

[680] Daniel P. Struthers, Lee F. G. Gutowsky, Tim C. D. Lucas, Neil J. Mochnacz, Christopher M. Carli, and Mark K. Taylor. Statistical stream temperature modelling with SSN and INLA: an introduction for conservation practitioners. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 81(4):417–432, April 2024. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/cjfas-
2023-0136.

Svantemann:2024:IEE

[681] Mette Munkholm Svantemann, Esther Savina, and Ludvig Ahm Krag. Improving escape efficiency in selective devices with the use of a dark tunnel. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 81(4):433–443, April 2024. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2023-0232.

Eighani:2024:SPS

[682] Morteza Eighani, Tiago Veiga Malta, Valentina Melli, and Finbarr G. O’Neill. Semi-pelagic self-adjusting otter boards: effects on the catching performance of a demersal trawl. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 81(4):444–458, April 2024. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/doi/

full/10.1139/cjfas-2023-0224.

Chen:2024:SMI

[683] Jiaying Chen, Jin Gao, and Fan Zhang. Spatiotemporal model improves survey indices for witch flounder stock assessment in the Grand Banks. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 81(4):459–487, April 2024. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2023-0101.

Warne:2024:TTH

[684] Connor P. K. Warne, Matthew M. Guzzo, Kevin Cazelles, Cindy Chu, Neil Rooney, and Kevin S. McCann. Thermal tolerance and habitat preferences mediate how freshwater fish body sizes respond to warming. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 81(4):488–496, April 2024. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2023-0233.

Cousido-Rocha:2024:NSA

[685] Marta Cousido-Rocha, Francisco Izquierdo, Joaquín Martínez-Minaya, M. Grazia Pennino, Hugo Mendes, Cristina Silva, Andreia V. Silva, María Saínza, and Santiago Cerviño. A novel statistical approach to deal with spatial bias in maturity ogive estimation. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 81(4):497–507, April 2024. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/cjfas-2023-0219.

Gilligan-Lunda:2024:HUA

[686] Erin K. Gilligan-Lunda, Adam Duarte, and James T. Peterson. Habitat use of anadromous and amphidromous sturgeons in North America: a systematic review. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 81(5):508–524, May 2024. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/cjfas-
2023-0222.

Davenport:2024:LBE

[687] Kaeli A. Davenport, Ronald G. Twibell, Matthew S. Piteo, Douglas P. Peterson, William P. Baker, Ryan K. Simmons, and Andrew R. Whiteley. Lab-based evaluation of the reproductive performance of Trojan (M_{Y Y}) brook trout (Salvelinus fontinalis). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 81(5):525–534, May 2024. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/doi/

full/10.1139/cjfas-2023-0245.

Sorel:2024:ILH

[688] Mark H. Sorel, Jeffrey C. Jorgensen, Richard W. Zabel, Mark D. Scheuerell, Andrew R. Murdoch, Cory M. Kamphaus, and Sarah J. Converse. Incorporating life history diversity in an integrated population model to inform viability analysis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 81(5):535–548, May 2024. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/doi/

full/10.1139/cjfas-2023-0118.

Burton:2024:USD

[689] Tim Burton, Ola Ugedal, Sten O. Karlsson, Eva B. Thorstad, and Geir H. Bolstad. Using scale-derived estimates of body size in analyses of Atlantic salmon life-history variation: a cautionary note. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 81(5):549–558, May 2024. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2023-0154.

Chang:2024:SDS

[690] Hsiao-Yun Chang, Richard Pendleton, Gregg Kenney, Kim McKown, William Eakin, John Maniscalco, and Yong Chen. Spatiotemporal dynamics of spawning habitat distribution of American shad (Alosa sapidissima) in the Hudson River Estuary under multi stressors. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 81(5):559–572, May 2024. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2023-0241.

Adelfio:2024:ECE

[691] Luca A. Adelfio, Steven M. Wondzell, Nathan J. Mantua, and Gordon H. Reeves. Expanded, compressed, or equal? Interactions between spawning window and stream thermal regime generate three responses in modeled juvenile emergence for Pacific salmon. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 81(5):573–588, May 2024. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/doi/

full/10.1139/cjfas-2023-0238.

Linner:2024:EST

[692] Robyn M. Linner and Yong Chen. Evaluating the spatial transferability of habitat suitability models: implications for conservation and management. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 81(5):589–599, May 2024. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2023-0293.

Bergman:2024:ECI

[693] Jordanna N. Bergman, Joseph R. Bennett, Valerie Minelga, Chantal Vis, Aaron T. Fisk, and Steven J. Cooke. Ecological connectivity of invasive and native fishes in a historic navigation waterway. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 81(5):600–619, May 2024. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2023-0207.

Weber:2024:EAE

[694] D. Nick Weber, Andrew T. Fields, Derek W. Chamberlin, William F. Patterson III, and David S. Portnoy. Epigenetic age estimation in a long-lived, deepwater scorpionfish: insights into epigenetic clock development. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 81(5):620–631, May 2024. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2023-0296.

Ridgway:2024:CWP

[695] Mark S. Ridgway, Darren A. Smith, and Allan H. Bell. Climate warming and projected loss of thermal habitat volume in lake populations of brook trout. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 81(5):632–645, May 2024. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2023-0192.

Kapur:2024:RWC

[696] Maia S. Kapur, M. A. Haltuch, B. Connors, A. Berger, K. Holt, K. Marshall, and A. E. Punt. Range-wide contrast in management outcomes for transboundary Northeast Pacific sablefish. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 81(7):810–827, July 2024. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/doi/

full/10.1139/cjfas-2024-0008.

Ransier:2024:ETR

[697] Krista T. Ransier, Paul Gatti, Arnault Le Bris, Cornelia E. den Heyer, Guy Claireaux, Brendan Wringe, and Jonathan A. D. Fisher. Electronic tags reveal high migratory diversity within the largest Atlantic halibut (Hippoglossus hippoglossus) stock. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 81(7):828–846, July 2024. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/doi/

full/10.1139/cjfas-2023-0282.

Tucker:2024:EFP

[698] Caroline M. Tucker, Henrique C. Giacomini, Nicholas E. Mandrak, Lifei Wang, and Derrick T. de Kerckhove. Estimating fish production in wetlands. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 81(7):847–861, July 2024. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2023-0220.

Quinn:2024:DLH

[699] Thomas P. Quinn, Lisa A. Wetzel, Daniel J. Hasselman, and Kim Larsen. Differences in life history patterns of American shad (Alosa sapi dissima) populations between ancestral, Atlantic coast, and non-native Pacific coast rivers of North America. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 81(7):862–878, July 2024. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/doi/

full/10.1139/cjfas-2023-0286.

Bihun:2024:ETB

[700] Christian J. Bihun, Megan K. Murphy, Timothy B. Johnson, Aaron T. Fisk, Matthew M. Guzzo, Charles P. Madenjian, and Graham D. Raby. Effects of temperature, body size, and sex on the standard metabolic rates of a sexually dimorphic freshwater piscivore, walleye (Sander vitreus). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 81(7):879–886, July 2024. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2023-0323.

Fortin:2024:MEF

[701] William Fortin, Olivier Morissette, Serge Lavoie, Marc Mingelbier, and Pascal Sirois. Matching the elemental fingerprints of fish otoliths with water masses of the St. Lawrence River and its tributaries, Canada. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 81(7):887–903, July 2024. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2023-0097.

Dawson:2024:SIA

[702] Jade C. Dawson, Matthew M. Guzzo, John M. Gunn, Erik J. S. Emilson, Kevin S. McCann, and Brie A. Edwards. Stable isotope analysis provides novel insights for measuring lake ecosystem recovery following acidification. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 81(7):904–918, July 2024. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2023-0305.

Howard:2024:EWG

[703] Tanner J. Howard, Anthony J. Orr, Kimberly Narel, and Carolyn M. Kurle. Estimating whisker growth rates and foraging profiles for northern fur seal pups and adult females using stable isotope analysis. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 81(7):919–934, July 2024. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2023-0156.

Battaile:2024:ESG

[704] Brian C. Battaile, Ethan A. Mora, Brendan M. Lehman, and Peter N. Dudley. Estimating spawning green sturgeon (Acipenser medirostris Ayres, 1854) abundance in the Sacramento River using side-scan sonar and N-mixture models. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 81(7):935–951, July 2024. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/cjfas-
2023-0228.

Pepin:2024:PFS

[705] Pierre Pepin. A probabilistic foundation for the study of larval fish feeding, growth, and mortality rates. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 81(7):952–967, July 2024. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/doi/

full/10.1139/cjfas-2023-0339.

Koelmans:2024:NAA

[706] Albert A. Koelmans, Todd Gouin, Alvine C. Mehinto, and Scott Coffin. On the need to avoid apple-to-orange comparisons in microplastic research. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 81(7):968–971, July 2024. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/cjfas-

2023-0325. See response [707].

Hataley:2024:RNA

[707] Eden K. Hataley, Hayley K. McIlwraith, Dimple Roy, and Chelsea M. Rochman. Response to: On the need to avoid apple-to-orange comparisons in microplastic research. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 81 (7):972–977, July 2024. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2024-0074. See [706].

Hataley:2024:CTM

[708] Eden K. Hataley, Hayley K. McIlwraith, Dimple Roy, and Chelsea M. Rochman. Correction: Towards a management strategy for microplastic pollution in the Laurentian Great Lakes — ecological risk assessment and management (part 2). Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 81(7):978–983, July 2024. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/cjfas-

2023-0346. See [708].

Rooper:2024:SPF

[709] Christopher N. Rooper, Jennifer L. Boldt, Andres Uriarte, Cecilie Hansen, Tim Ward, and Sarah Gaichas. Small pelagic fish: new frontiers in science and sustainable management. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 81 (8):984–989, August 2024. CODEN CJFSDX. ISSN 0706-652X (print),

1205-7533 (electronic). URL https://cdnsciencepub.com/doi/full/
10.1139/cjfas-2024-0104.

Gaichas:2024:ASP

[710] Sarah K. Gaichas, James Gartland, Brian E. Smith, Anthony D. Wood, Elizabeth L. Ng, Michael Celestino, Katie Drew, Abigail S. Tyrell, and James T. Thorson. Assessing small pelagic fish trends in space and time using piscivore stomach contents. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 81(8):990–1012, August 2024. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/doi/

full/10.1139/cjfas-2023-0093.

Citores:2024:BSE

[711] Leire Citores, Leire Ibaibarriaga, Maria Santos, and Andres Uriarte. A Bayesian spatially explicit estimation of daily egg production: application to anchovy in the Bay of Biscay. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 81(8):1013–1028, August 2024. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/doi/

full/10.1139/cjfas-2023-0126.

Wildermuth:2024:ERH

[712] Robert P. Wildermuth, Desiree Tommasi, Peter Kuriyama, James Smith, and Isaac Kaplan. Evaluating robustness of harvest control rules to climate-driven variability in Pacific sardine recruitment. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 81(8):1029–1051, August 2024. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https:/

/cdnsciencepub.com/doi/full/10.1139/cjfas-2023-0169.

Makinen:2024:EIT

[713] K. Mäkinen, M. Rajasilta, S. Ruuskanen, T. Karpela, A. Lauerma, and J. Sahlstén. Effects of incubation temperature and maternal phenotype on Baltic herring (Clupea harengus membras) eggs and larvae: an experimental study. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 81(8):1052–1065, August 2024. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/cjfas-
2023-0032.

Moosa:2024:IIM

[714] Naseera Moosa and Doug S. Butterworth. Investigating the influence of minor krill-predators on the krill-predator dynamics of the Antarctic ecosystem in the International Whaling Commission’s Management Area II. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 81(8):1066–1080, August 2024. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/cjfas-
2023-0086.

Schiano:2024:EAH

[715] Samantha Schiano, Geneviève M. Nesslage, Katie Drew, Amy M. Schueller, Ryan J. Woodland, and Michael J. Wilberg. Evaluation of alternative harvest policies for striped bass and their prey, Atlantic menhaden. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 81(8):1081–1103, August 2024. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/cjfas-
2023-0089.

Kell:2024:DMP

[716] Laurence T. Kell, Jacob W. Bentley, David A. Feary, Afra Egan, and Cormac Nolan. Developing management plans for sprat (Sprattus sprattus) in the Celtic Sea to advance the ecosystem approach to fisheries. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 81(8):1104–1121, August 2024. CO-DEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2023-0090.

deMoor:2024:EIE

[717] Carryn L. de Moor. Explicitly incorporating ecosystem-based fisheries management into management strategy evaluation, with a focus on small pelagics. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 81(8):1122–1134, August 2024. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (elec-

tronic). URL https://cdnsciencepub.com/doi/full/10.1139/cjfas-
2023-0092.

Quezada:2024:CCC

[718] Felipe J. Quezada, Desiree Tommasi, Timothy H. Frawley, Barbara Muhling, Isaac Kaplan, and Stephen Stohs. Catch as catch can: markets, availability, and fishery closures drive distinct responses among the U.S. West Coast coastal pelagic species fleet segments. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 81(8):1135–1153, August 2024. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL https://cdnsciencepub.com/

doi/full/10.1139/cjfas-2023-0094.

Beckensteiner:2024:SAA

[719] Jennifer Beckensteiner, Sebastian Villasante, Anthony Charles, Pierre Petitgas, Christelle Le Grand, and Olivier Thébaud. A systemic approach to analyzing post-collapse adaptations in the Bay of Biscay anchovy fishery. Canadian Journal of Fisheries and Aquatic Sciences = Journal canadien des sciences halieutiques et aquatiques, 81(8):1154–1173, August 2024. CODEN CJFSDX. ISSN 0706-652X (print), 1205-7533 (electronic). URL

https://cdnsciencepub.com/doi/full/10.1139/cjfas-2023-0087.